### v2

# Discontinuous integro-differential equations and sliding mode control

Abstract
The paper deals with analysis and design of sliding mode control systems modeled by finite-dimensional integro-differential equations. Filippov method and equivalent control approach are extended to a class of nonlinear discontinuous integro-differential equations. Sliding mode control algorithms are designed for distributed input delay systems and for systems modeled by differential equations in Banach spaces. The obtained results are illustrated by numerical examples.
1 Introduction
1.1 State of the art
Sliding Mode Control (SMC) is a robust control methodology, which uses discontinuous feedback laws for rejection of matched disturbances. The design and analysis of SMC is well-developed for finite-dimensional systems [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book]. Mathematical models of finite-dimensional SMC systems are ill-posed discontinuous ODEs (Ordinary Differential Equations). Filippov method [Filippov1988:Book], [Cortes2008:CSM] is aimed at regularization of discontinuous ODEs (Ordinary Differential Equations). The method consists in a minimal (in a set-theoretic sense) extension of an ill-posed discontinuous ODE to a well-posed ODI (Ordinary Differential Inclusion). The Filippov regularization has a rather simple mathematical formulation [Filippov1988:Book] based on a notion of the set of Lebesgue measure zero. Filippov method is the main tool for regularization of a SMC system [Filippov1988:Book], [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book], [PolyakovFridman2014:JFI], [Utkin_etal2020:Book] while an analysis of the closed-loop dynamics of the SMC system is usually based on the so-called equivalent control method [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book].
Formally, the finite-dimensional Filippov method is inapplicable to discontinuous infinite-dimensional systems in abstract spaces [DaleckyKrein1974:Book] (e.g., due to the absence of a notion of the set of Lebesgue measure zero in an abstract space). Traditionally (see, e.g., [OrlovUtkin1985:DAN], [Orlov_elal2004:IJC], [Orlov2008:Book] and references therein), SMC design for infinite-dimensional systems is based on an equivalent control method. Well-posedness analysis of infinite-dimensional SMC systems requires rather complicated constructions [Levaggi2002:DIE], [Pisano_etal2011:SIAM_JCO], [Levaggi2013:DCDS], [Orlov2020:Book]. Infinite-dimensional differential inclusion obtained from discontinuous infinite-dimensional differential equation in an abstract space (like in [Mohet_etal2023:SCL]) is a non-trivial object for research [Tolstonogov2000:Book]. Some generalized solutions to discontinuous functional differential equations of retarded type are introduced on [KolmanovskiiMyshkis1992:Book, page 58]. The corresponding regularization procedure consists in a construction of a convex hull of limits of all possible convergent sequences in a function space. Applicability of such non-constructive procedure to analysis and design of SMC systems seems questionable too.
Integro-Differential Equations (IDEs) [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book] model various finite and infinite dimensional systems, such as epidemics [KermackMcKendrick1991:BMB], dispersive waves [Debnath2005:Book], systems with delays [Hale1977:Book], [Fridman2014:Book], [Orlov2020:Book], and others. Input-output dynamics of some systems modeled by PDEs (Partial Differential Equations) can also be described by IDEs (see Section 1.2). Being a model of (possibly) infinite-dimensional system, an IDE is usually defined by functions in finite-dimensional spaces (see, e.g., [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book]). To the best of author’s knowledge, finite-dimensional Filippov method and SMC theory has never been developed for IDEs.
1.2 Motivation example: SMC design for a heat system
Let us consider a control system modeled by the PDE
$\tfrac{\partial x}{\partial t}\!=\!\nu\tfrac{\partial^{2}x}{\partial z^{2}}+\!\beta(z)(u(t)\!+\!\gamma(t)),\quad t\geq t_{0}\vskip-2.84526pt$ (1)
with the homogeneous Dirichlet boundary conditions
$x(t,0)=0,\quad x(t,1)=0,\vskip-2.84526pt$ (2)
and the initial condition
$x(t_{0},z)=x^{0}(z),\quad z\in[0,1],\vskip-2.84526pt$ (3)
where $x(t,\cdot)$ is a distributed state of the system, $u(t)\in\mathbb{R}$ is the control input, $\beta:\mathbb{R}\to\mathbb{R}$ is a function that models an effect of the input to the system, $\gamma\in L^{\infty}(\mathbb{R};\mathbb{R})$ is an unknown function (perturbation) with a known upper bound $\|\gamma\|_{L^{\infty}(\mathbb{R};\mathbb{R})}\leq\overline{\gamma}$, $\nu>0$ is a conductivity coefficient and $x^{0}$ is the initial state satisfying the compatibility condition, e.g.,
$x^{0}\in\tilde{\mathcal{D}}=\{x\in C^{2}([0,1];\mathbb{R}):x(0)=x(1)=0\}.$ (4)
A control system similar to (1) also appears in the case of the boundary control design using the Dirichlet lifting [PrieurTrelat2019:TAC], [KatzFridman2021:EJC].
Let a measured output of the system be modeled as
[…]
The control aim is to steer the output $y$ to zero in a finite time rejecting the effect of the unknown perturbation $\gamma$.
Let a discontinuous output control law be defined as
$u(t)\!=\!q(t)\mathrm{sign}(y(t)),\quad q\in L^{\infty}([t_{0},+\infty);\mathbb{R}).\vskip-2.84526pt$ (6)
This is a SMC provided that $q$ is properly selected [Orlov_elal2004:IJC], [Orlov2020:Book].
The infinite-dimensional closed-loop control system (1) – (6) is ill-posed in the SMC case. A regularization of this discontinuous PDEs with SMC is a nontrivial problem (see, e.g., [Orlov2020:Book, pages 24-28], [Mohet_etal2023:SCL]). However, the considered infinite-dimensional control system has the scalar input $u$ and the scalar output $y$. The discontinuous feedback $u$ is a function of output variable only. For such infinite-dimensional system, the well-posedness analysis can be reduced to a study of a finite-dimensional IDE that represents the input-output dynamics:
$\dot{y}(t)\!=\!p(t-t^{0})+b(u(t)+\!\gamma(t))+\!\!\int^{t}_{t_{0}}\!\!\!\!\Phi(t-\tau)(u(\tau)\!+\!\gamma(\tau))d\tau,\vskip-5.69054pt$ (7)
where $y(0)=\int^{1}_{0}\xi(z)x^{0}(z)dz$, $b=\int^{1}_{0}\xi(z)\beta(z)dz$, and $p,\Phi$ are some decreasing functions (see Section 4). If the control law $u$ is defined by the formula (6), then the input-output dynamics of the system (1) - (6) should satisfy the discontinuous IDE (7), (6). Due to discontinuity of $u$ on $y$, this IDE is ill-posed [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book]. However, the IDE is defined in terms of finite-dimensional functions, which admit Filippov regularization [Filippov1988:Book]. This paper extends the Filippov method to a class of finite-dimensional IDEs, which covers the IDE (7), (6) as a particular case. A regularization of the IDE (7), (6) allows the original discontinuous PDE (1) – (6) to be easily regularized (see, e.g., [Polyakov2025:CDC]). Such an IDE-based approach to SMC design and analysis is applicable to a rather large class of systems modeled by differential equations in Banach spaces
1.3 Contributions of the paper
  • The Filippov method is extended to the IDE
    $\begin{split}\dot{x}(t)=&f(t,x(t),u(t,x(t)))\\ &+\!\int^{t}_{t^{0}}\!\!\Phi(t,\tau)\tilde{f}(\tau,x(\tau),u(\tau,x(\tau)))d\tau,\vskip-2.84526pt\end{split}$ (8)
    where $t>t^{0}\in\mathbb{R},x(t)\in\mathbb{R}^{n}$, $f,\tilde{f}:\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$, $\Phi:\mathbb{R}\times\mathbb{R}\to\mathbb{R}^{n\times n}$, $u:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a (possibly) discontinuous state feedback, which may be a SMC.
  • The well-posedness analysis (existence, uniqueness and continuous dependence of Filippov solutions on initial data) is presented for the discontinuous IDE (8) .
  • The equivalent control method is developed for affine-in-control IDEs (i.e., for $f=a+bu$ and $\tilde{f}=\tilde{a}+\tilde{b}u$).
  • A SMC design for some control systems with distributed input delay is presented.
  • An IDE-based method of SMC design and analysis for systems modeled by differential equations in Banach spaces is developed and illustrated on the system (1)-(6).
1.4 Organization of the paper
The paper is organized as follows. First, Filippov method for discontinuous ODEs is briefly surveyed. The surveyed results are utilized in the next section, where an extension of Filippov method to IDEs is proposed. After that, the developed theoretical results are applied for SMC design and analysis of finite-dimensional systems with distributed input delay and of infinite-dimensional systems modeled by differential equations in Banach spaces. Finally, concluding remarks are given.
2 Preliminaries: Discontinuous ODEs
2.1 Some classes of functions
For $1\!\leq\!p\!\leq\!\infty$, $L^{p},$ denotes the well-known Lebesgue space. A function $(t,x)\in\mathbb{R}\times\mathbb{R}^{n}\mapsto f(t,x)\in\mathbb{R}^{n}$ is said to satisfy
  • Carathéodory Condition if it is measurable in $t$ for any $x$, continuous in $x$ for almost any $t$ and for any $\alpha,r>0$ there exists $m\in L^{1}((-\alpha,\alpha);\mathbb{R})$ such that $\left\|f(t,x)\right\|\leq m(t)$ for almost all $t\!\in\!(-\alpha,\alpha)$, for all $x\!\in\!\mathcal{B}(r)$, where $\mathcal{B}(r)$ is an open ball of the radius $r>0$.
  • One-Sided Lipschitz Condition on $\mathcal{X}$ if $\exists\ell\in L^{\infty}_{\rm loc}(\mathbb{R};\mathbb{R})$:
    $(x-y)^{\top}(f(t,x)-f(t,y))\!\leq\!\ell(t)\|x-y\|^{2},\vskip-2.84526pt$ (9)
    $\forall t\in\mathbb{R},\forall x,y\!\in\!\mathcal{X}$, where $\ell$ may depend on $\mathcal{X}\subset\mathbb{R}^{n}$;
  • Lipschitz Condition on $\mathcal{X}$ if $\exists\ell\in L^{\infty}_{\rm loc}(\mathbb{R};\mathbb{R})$:
    $\|f(t,x)-f(t,y)\|\leq\ell(t)\|x-y\|,\quad t\!\in\!\mathbb{R},x,y\!\in\!\mathcal{X};\vskip-2.84526pt$ (10)
  • Filippov Condition if $f$ is measurable and for any $\alpha,r\!>\!0$ there exists $m\!\in\!L^{1}((-\alpha,\alpha);\mathbb{R})$ : $\left\|f(t,x)\right\|\!\leq\!m(t)$ for almost all $t\in(-\alpha,\alpha)$ and for almost all $x\in\mathcal{B}(r)$.
A measurable function $f:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is said to be locally (one-sided) Lipschitz on $\mathcal{X}$ if it satisfies (One-Sided) Lipschitz Condition on any compact from $\mathcal{X}\subset\mathbb{R}^{n}$. If $\mathcal{X}=\mathbb{R}^{n}$ we write shortly locally (one-sided) Lipschitz.
A measurable function $(t,x)\in\mathbb{R}\times\mathbb{R}^{n}\mapsto f(t,x)\in\mathbb{R}^{n}$ is said to be piece-wise (one-sided) Lipschitz continuous in $x$ if
  • the space $\mathbb{R}^{n}$ can be split into a finite number of open connected sets $G^{j}\subset\mathbb{R}^{n},j=1,2,...,N$ with non-empty interiors and continuous boundaries $\partial G^{j}$;
  • $\bigcup_{i=1}^{n}\overline{G^{j}}=\mathbb{R}^{n},G^{i}\bigcap G^{j}=\emptyset\text{ for }i\neq j$ and the boundary set $\mathcal{S}=\bigcup_{i=1}^{N}\partial G^{j}$ is of measure zero
  • for each j=1,,Nj=1,...,N, there exists a locally (one-sided) Lipschitz continuous function $(t,x)\in(\mathbb{R}\times\mathbb{R}^{n}\mapsto f^{j}(t,x)\in\mathbb{R}^{n})$ such that $f^{j}(t,x)\!=\!f(t,x)$ for all $x\!\in\!G^{j}$ and almost all $t\!\in\!\mathbb{R}$.
2.2 Well-posed ODEs
A problem is said to be well-posed in Hadamard sense if 1) it has a solution; 2) the solution is unique; 3) the solution changes continuously with the initial data. The ODE
$\dot{x}(t)=f(t,x(t)),\quad f:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}\vskip-2.84526pt$ (11)
defines a rule of an evolution of the state $x(t)\in\mathbb{R}^{n}$ of the system in time $t\in\mathbb{R}$. Let $\mathcal{I}$ be one of the following intervals: $[a,b]$, $(a,b)$, $[a,b)$ or $(a,b]$, where $a,b\in[-\infty,+\infty]$.
Definition 1 (Classical solution of ODE).
A function $x\in C^{1}(\mathcal{I};\mathbb{R}^{n})$ is said to be a classical solution of ODE (11) if it satisfies the equation (11) everywhere on $\mathcal{I}$.
The Initial Value Problem (IVP) is to find a solution (11) for a given initial state $x_{0}\in\mathbb{R}^{n}$ at a given initial time $t_{0}\in\mathbb{R}$:
$x(t^{0})=x^{0}\in\mathbb{R}^{n}.\vskip-2.84526pt$ (12)
For the proofs of the classical theorem surveyed in this subsection we refer the reader, for example, to [CoddingtonLevinson1955:Book], [Filippov1988:Book].
Theorem 1 (Peano Existence Theorem, 1890).
If $f$ is continuous then the IVP (11), (12) has a classical solution defined, at least, locally in time.
The identity (11) may hold almost everywhere if $x:\mathcal{I}\mapsto\mathbb{R}^{n}$ is an absolutely continuous function [^1] A function $x:\mathcal{I}\mapsto\mathbb{R}^{n}$ is said to be absolutely continuous if $\forall\varepsilon>0$, $\forall\delta>0$ such that whenever a finite sequence of pairwise disjoint intervals $(\alpha_{k},\beta_{k})\subset\mathcal{I}$ satisfies $\sum_{k}(\beta_{k}-\alpha_{k})<\delta$ then $\sum_{k}|x(\beta_{k})-x(\alpha_{k})|<\varepsilon$..
Definition 2 (Carathéodory solution of ODE).
An absolutely continuous function $x:\mathcal{I}\subset\mathbb{R}^{n}\to\mathbb{R}^{n}$ is said to be a Carathéodory (or strong) solution of ODE (11) if it satisfies the equation (11) almost everywhere on the time interval $\mathcal{I}$.
The existence of Carathéodory solutions can be proven for ODEs with discontinuous in $t$ right-hand sides.
Theorem 2 (Carathéodory Existence Theorem, 1918).
f $f$ satisfies Carathéodory Condition then the IVP (11), (12) has a Carathéodory solution defined, at least, locally in time.
The strong solutions allow some systems with deterministic noise to be modeled by ODEs, e.g., the system $\dot{x}(t)=-x(t)+\gamma(t)$ does not have a classical solution if $\gamma\in L^{\infty}(\mathbb{R};\mathbb{R})$ is a discontinuous signal, but it has the strong solution.
Let $x(t,t^{0},x^{0})$ denote a solution of the IVP (11), (12).
Theorem 3 (Continuity of solutions on initial data [CoddingtonLevinson1955:Book], page 58).
If $f$ satisfies Carathéodory Condition, then $x(t,t^{1},x^{1})\to x(t,t^{0},x^{0})$ as $(t^{1},x^{1})\to(t^{0},x^{0})$ uniformly on $\mathcal{I}$ provided that the solution $x(t,t^{0},x^{0})$ is unique on $\mathcal{I}$.
Uniqueness of solutions can be guaranteed under additional restrictions to the right-hand side of (11).
Theorem 4 (Uniqueness of solutions).
If a locally (one-sided) Lipschitz continuous function $f$ satisfies Carathéodory Condition then the IVP (11), (12) has a unique Carathéodory solution defined, at least, locally in the (resp., forward) time.
2.3 Regularization of ill-posed ODE by Filippov method
Systems with discontinuous in $x$ right-hand sides may have Carathéodory solutions in some cases [Liberzon2003:Book]. However, SMC theory deals with control systems modeled by ill-posed ODEs with discontinuous right-hand sides [Utkin_etal2009:Book, Liberzon2003:Book, GeligLeonovYakubovich2004:Book].
Example 1.
Let us consider the following SMC system
$\dot{x}(t)\!=\!u(t)\!+\!0.5\sin(3t),\quad u(t)\!=\!-\mathrm{sign}(x(t)),\vskip-2.84526pt$ (13)
where $t\in\mathbb{R}$ is the time variable, $x(t)\in\mathbb{R}$ is the state variable, $u(t)\in\mathbb{R}$ is the control variable, $\mathrm{sign}(x(t))=1$ for $x(t)>0$, $\mathrm{sign}(x(t))=-1$ for $x(t)<0$ and $\mathrm{sign}(0)=0$. The ODE (13) admits classical solutions on $\{x\!\in\!\mathbb{R}:x\!>\!0\}$ and on $\{x\!\in\!\mathbb{R}:x\!<\!0\}$. Each solution with $x(t_{0})\!=\!x_{0}\!\neq\!0$ reaches zero in a finite time $t=t^{0}+T(t^{0},x^{0})$, since $\frac{d|x(t)|}{dt}\leq-0.5$ for $x(t)\neq 0$. Hence, the natural continuation of a solution for $t>t^{0}+T(t^{0},x^{0})$ is $x(t)=0$, but neither Peano nor Carathéodory’s existence theorem is applicable in this case. In addition, formally, if $x(t)=0$ for all $t>t^{0}+T(t^{0},x^{0})$ then $\dot{x}(t)=0=u(t)+0.5\sin(3t)$ for $t>t^{0}+T(t^{0},x^{0})$, but, by definition, $\mathrm{sign}(x(t))=0$ for $x(t)=0$.
To define a solution of the system (11) with discontinuous (on $x$) right-hand side, a special regularization procedure known as the Filippov method [Filippov1988:Book] can be utilized. It extends an ill-posed ODE to a well-posed ODI
$\dot{x}\in K[f](t,x),\vskip-2.84526pt$ (14)
where $K[f]:\mathbb{R}\times\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$ is defined as [Filippov1988:Book, page 85]:
[…]
where $\overline{\mathrm{co}}$ denotes the convex closure, and the intersections are taken over all sets $M\subset\mathbb{R}^{n}$ of measure zero and all $\delta>0$. The set-valued mapping $(t,x)\to K[f](t,x)$ is nonempty-valued, compact-valued, convex-valued and upper semi-continuous in $x$ for almost all $t$ provided that $f$ is a measurable function (see, e.g., [Filippov1988:Book]). Filippov regularization (15) does not modify continuous in $x$ function $f$.
Definition 3 (Filippov solution).
An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ is said to be a Filippov solution of (11) if it satisfies the ODI (14), (15) almost everywhere on $\mathcal{I}$.
Filippov method regularizes various discontinuous ODEs.
Example 1 (continued).
Any solution of the discontinuous differential equation (13) can be prolonged by Filippov method for $t>t^{0}+T(t^{0},x^{0})$. The corresponding differential inclusion
[…]
obviously, admits the forward unique zero solution.
Theorem 5 (Filippov Existence Theorem, 1960).
If $f$ satisfies Filippov Condition, then the IVP (11), (12) has a Filippov solution defined, at least, locally in time.
Continuity of Filippov solutions on initial data [Filippov1988:Book, page 98] repeats the classical result if solutions of (14) are unique.
Theorem 6 (Continuity of Filippov solutions on initial data).
Let $f$ satisfy Filippov Conditions and let a Filippov solution $x(t,t^{0},x^{0})$ of the IVP (11), (12) be defined and unique on some time interval $\mathcal{I}$. Then $x(t,t^{1},x^{1})\to x(t,t^{0},x^{0})$ as $(t^{1},x^{1})\to(t^{0},x^{0})$ uniformly on compacts from $\mathcal{I}$.
2.4 Uniqueness of Filippov solutions of ODEs
Lipschitz condition does not hold for discontinuous (on state variable) ODEs. Filippov solution of a discontinuous ODE is understood as a solution of an ODI. The uniqueness analysis for solutions of the ODI is a non-trivial problem in the general case [Filippov1988:Book], [Tolstonogov2000:Book]. However, One-Sided Lipschitz Condition is still applicable for uniqueness analysis of Filippov solutions.
Theorem 7 (Uniqueness of Filippov solutions, [Filippov1988:Book], §10).
If a locally one-sided Lipschitz function $f$ satisfies Filippov Condition then the IVP (11), (12) has a unique Filippov solution defined, at least, locally in the forward time.
Notice that One-Sided Lipschitz Condition is still rather conservative. For example, the right-hand side of the system
$\dot{x}_{1}(t)=-\mathrm{sign}(x_{1}(t),\quad\dot{x}_{2}(t)=\mathrm{sign}(x_{1}(t))\vskip-2.84526pt$ (16)
does not satisfy One-Sided Lipschitz Condition, but its Filippov solution is unique. Less conservative sufficient conditions for uniqueness of Filippov solution can be obtained for particular classes of discontinuous ODEs. The Filippov regularization (15) of a piece-wise continuous system gives
[…]
where the set-valued function $\mathcal{N}:\mathcal{S}\to 2^{\{1,2,...,N\}}$ indicates the domains, which have a common boundary point $x\in\mathcal{S}$, i.e., $\mathcal{N}(x)=\left\{j\in\{1,2,...,N\}:x\in\partial G^{j}\right\}.$
Let $\mathcal{C}_{\Omega}(x)\subset\mathbb{R}^{n}$ denote the (Bouligand) tangent cone to the set $\Omega\subset\mathbb{R}^{n}$ at $x\in\mathbb{R}^{n}$ defined as (see., e.g. [Nagumno1942:PPMSJ])
$\mathcal{C}_{\Omega}(x)=\left\{z\in\mathbb{R}^{n}:\liminf_{h\to 0^{+}}\tfrac{\inf_{y\in\Omega}\|x+hz-y\|}{h}=0\right\}.\vskip-2.84526pt$ (18)
Tangent cones characterize positively invariant sets [^2] A set $M\!\subset\!\mathbb{R}^{n}$ is positively invariant if $x(t^{0})\!\in\!M\Rightarrow x(t)\!\in\!M,t\!\geq\!t_{0}.$ of dynamical systems (see, e.g., [Clarke_etal1995:JDCS], [Blanchini1999:Aut], [AubinEkeland1984:Book] and references therein).
Assumption 1.
Let a piece-wise one-sided Lipschitz continuous function $f$ be such that $f^{j}(t,x)\!\notin\!\mathcal{C}_{\bar{G}^{j}}(x)$ for all $x\!\in\!\partial G^{j}$, for almost all $t\!\in\!\mathbb{R}$ and for all j=1,2,,Nj\!=\!1,2,...,N.
Let us denote $F_{\mathcal{S}}(t,x)=K[f](t,x)\cap\mathcal{C}_{\mathcal{S}}(x),t\in\mathbb{R},x\in\mathcal{S}$. The following result is partially inspired by [Filippov1988:Book, §10].
Theorem 8 (Uniqueness of solutions in sliding mode case).
If the set $F_{\mathcal{S}}$ is a singleton $F_{\mathcal{S}}(t,x)=\{f^{0}(t,x)\}$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$ and for some locally one-sided Lipschitz continuous function $f^{0}$, then, under Assumption 1, the IVP (11), (12) has a Filippov solution being unique, at least, locally in the forward time.
Proof.
[…]
Assumption 1 guarantees that the discontinuity set $\mathcal{S}$ is the sliding set of the system. The dynamics on $\mathcal{S}$ is governed by
$\dot{x}(t)\in F_{\mathcal{S}}(t,x(t)).\vskip-2.84526pt$ (19)
The system (16) satisfies the above theorem with $F_{\mathcal{S}}=\{0\}$. The relay SMC system is the conventional example of discontinuous ODE satisfying the above theorem [Utkin1992:Book], [Filippov1988:Book].
Example 2.
The simplest example of discontinuous ODE with unique solutions can be discovered in [Utkin1992:Book] for the system
$\dot{x}(t)=Ax(t)+Bu(t),\vskip-2.84526pt$ (20)
with $u(t)=-(CB)^{-1}CAx(t)+L\,\mathrm{sign}(Cx(t))$, where $A\in\mathbb{R}^{n\times n},B\in\mathbb{R}^{n\times m},L\in\mathbb{R}^{m\times m}$, $C\in\mathbb{R}^{m\times n}$, $\det(CB)\neq 0$ and $\mathrm{sign}$ of a vector is understood component-wisely. The dynamics of the sliding variable $\sigma=Cx$ is given by
$\dot{\sigma}(t)=CB{L}\mathrm{sign}(\sigma(t)).\vskip-2.84526pt$
If ${L}\!=\!\rho(CB)^{-1}$, $\rho\!>\!0$ then conditions of Theorem 8 are fulfilled and $f^{0}(t,x)\!=\!(A\!-\!B(CB)^{-1}CA)x$ for $Cx\!=\!\mathbf{0}$.
The uniqueness of Filippov solutions of piece-wise continuous ODE without sliding sets can be proven under the assumption.
Assumption 2.
Let $f$ be piece-wise one-sided Lipschitz continuous and $F_{\mathcal{S}}(t,x)$$\emptyset$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$.
Theorem 9 (Uniqueness of solutions in switching case).
Under Assumption 2, the IVP (11), (12) has a Carathéodory solution being unique, at least, locally in the forward time.
Proof.
The condition ${F_{\mathcal{S}}(t,x)=}K[f](t,x)\cap\mathcal{C}(x)=\emptyset$ means that any Filippov solution initiated on the discontinuity set $\mathcal{S}$ immediately leaves the set $\mathcal{S}$ by entering into some $G^{j}$. Since $f^{j}$ satisfies Theorem 4 then any Filippov solution of (11) is a unique Carathéodory solution. ∎
Assumption 3.
Let $f$ be piece-wise one-sided Lipschitz continuous function and there exists a closed set $\mathcal{M}\subset\mathcal{S}\subset\mathbb{R}^{n}$:
  • $F_{\mathcal{S}}(t,x)$$\emptyset$, $\forall x\!\in\!\mathcal{S}\backslash\mathcal{M}$, for almost all $t\!\in\!\mathbb{R}$;
  • $f^{j}(t,x)\!\notin\!\mathcal{C}_{\bar{G}^{j}}(x)$ for all $x\!\in\!\partial G^{j}\cap\mathcal{M}$, for almost all $t\!\in\!\mathbb{R}$ and for all $j\!=\!1,2,...,N:\partial G^{j}\cap\mathcal{M}\neq\emptyset$.
The following corollary combines both above theorems.
Corollary 1 (Uniqueness of solution of switching sliding mode).
Under Assumption 3, if $F_{\mathcal{S}}(t,x)\!=\!\{f^{0}(t,x)\}$} is a singleton for all $x\!\in\!\mathcal{M}$ and almost all $t\!\in\!\mathbb{R}$ and $f^{0}$ is locally one-sided Lipschitz continuous, then the IVP (11), (12) has a Filippov solution being unique, at least, locally in the forward time.
Assumption 3 ensures that $\mathcal{M}$ is the sliding set and the (switching) set $\mathcal{S}\backslash\mathcal{M}$ is crossed by trajectories without sliding.
Example 3.
The system (20) with $m=2$, $\sigma=(\sigma_{1},\sigma_{2})^{\top}$, ${L}=\rho(CB)^{-1}\left(\begin{smallmatrix}-1&2\\ -2&{-1}\end{smallmatrix}\right)$ and $\rho>0$ satisfy all conditions of Corollary 1 with $\mathcal{M}=\{0\}$.
2.5 Equivalent control method for systems modeled by ODEs
Let us consider a control system modeled by ODE
$\dot{x}(t)=f(t,x(t),u(t,x(t))),\vskip-2.84526pt$ (21)
where $t$ are $x$ are as before, $f\in\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ and $u:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a measurable function (e.g., SMC). In this case, some alternative methods of regularization of discontinuous ODE can be proposed (e.g., [Utkin1992:Book], [Filippov1988:Book], [PolyakovFridman2014:JFI]).
Definition 4 (Utkin solution).
An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ is said to be an Utkin solution of the first kind (resp., of the second kind) to (21) if there exists a measurable function $u_{\rm eq}:\mathcal{I}\to\mathbb{R}^{m}$ such that
$\dot{x}(t)=f(t,x(t),u_{\rm eq}(t)),\quad u_{\rm eq}(t)\in U(t,x(t))\vskip-2.84526pt$ (22)
almost everywhere on $\mathcal{I}$, where $U=K[u]$ is defined by (15) (resp., $U=(K[u_{1}],...,K[u_{m}])^{\top}$ for $u=(u_{1},u_{2},...,u_{m})^{\top}$).
Since $K[u](t,x(t))\subset(K[u_{1}](t,x(t)),...,K[u_{m}(t,x(t))])^{\top}$ then the existence of the Utkin solutions of the first kind always implies the existnce of the Utkin solutions of the second kind. They always coincide for single input systems.
Example 4 (Neimark 1960).
A simple analysis (see, e.g., [Neimark1961:IFAC_Congress] or [PolyakovFridman2014:JFI, Example 2]) shows that the set of Utkin solutions of the first kind do not coincide with the set of Utkin solutions of the second kinds for the linear time-invariant control system (20) with $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, $A=\mathbf{0}$, $B=I_{2}$ and the static relay feedback law $u=(u_{1},u_{2})^{\top},u_{1}=u_{2}=\mathrm{sign}(x_{1})$.
The function $u_{eq}$ is the so-called equivalent control [Utkin1992:Book], [Filippov1988:Book], which, in SMC theory, is defined as a solution of a special algebraic equation (see below). The equivalent control method is the main tool for analysis and design of SMC systems [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book], [Utkin_etal2020:Book], [Orlov2020:Book]. The existence of $u_{\rm eq}$ follows from [Filippov1962:SIAM].
Lemma 1 (Filippov 1959).
Let a function $(z,u)\in\mathbb{R}^{p}\times\mathbb{R}^{m}\mapsto w(z,u)\in\mathbb{R}^{n}$ be measurable in $z$ and continuous in $u$. Let a compact-valued mapping $U:\mathbb{R}^{p}\to 2^{\mathbb{R}^{m}}$ and a function $v:\mathbb{R}^{p}\to\mathbb{R}^{m}$ be measurable. If $v(z)\in\{w(z,u):u\in U(z)\}$ almost everywhere then $\exists u_{eq}:\mathbb{R}^{p}\to\mathbb{R}^{m}$ being measurable, $u_{\rm eq}(z)\in U(z)$ and $v(z)=w(z,u_{eq}(z))$ almost everywhere.
The set-valued mapping $t\mapsto U(t,x(t))$ from Definition 4 satisfies Lemma 1, so an Utkin solution exists if the ODI
$\dot{x}(t)\in f(t,x(t),U(t,x(t))),\quad t\in\mathbb{R},\vskip-2.84526pt$ (23)
has a solution (see, e.g., [Filippov1988:Book] or [PolyakovFridman2014:JFI]). The mapping $(t,x)\mapsto f(t,x,U(t,x))$ may be nonconvex-valued, so the conventional existence theorems for ODIs [Filippov1988:Book] are not applicable [BartoliniZolezzi1985:IEEE_TAC].
The equivalence of Filippov and Utkin solutions for affine-in-control systems was studied in the literature [Utkin1992:Book].
Theorem 10.
Let the system (21) be affine on the input
$f(t,x)=a(t,x)+b(t,x)u(t,x),\vskip-2.84526pt$
and the locally essentially bounded [^3] A function $\xi:\mathcal{X}\subset\mathbb{R}^{p}\to\mathbb{R}^{q}$ is said to be locally essentially bounded on $\mathcal{X}$ if $\operatorname*{ess\,sup}_{z\in\mathcal{Z}}\|\xi(z)\|<+\infty$ for any compact $\mathcal{Z}\subset\mathcal{X}$. functions $a:\mathbb{R}\times\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ and $b:\mathbb{R}\times\mathbb{R}^{n}\mapsto\mathbb{R}^{n\times m}$ be continuous in $x$ for almost all $t$ and measurable in $t$ for all $x$. If a measurable function $u:\mathbb{R}\times\mathbb{R}^{n}\mapsto\mathbb{R}^{m}$ is locally essentially bounded then, $\forall t^{0}\in\mathbb{R}$, $\forall x^{0}\in\mathbb{R}^{n}$, the IVP (11), (12) has an Utkin solution of the first kind defined, at least, locally in the time; any Utkin solution of the first kind is a Filippov solution and vice versa.
This follows from Filippov existence theorem and the identity
$K[f](t,x,u(t,x))=a(t,x)+b(t,x)K[u](t,x),\vskip-2.84526pt$
being fulfilled for almost all $t$, due to continuity of $a$, $b$ in $x$.
A sufficient condition of equivalence of Utkin solutions of the first and of the second kinds is presented, for example, in [Zolezzi2002:inBook, Theorem 14, page 44] for affine-in-control systems.
If $\mathcal{S}=\{x\in\mathbb{R}^{n}:s(x)=\mathbf{0}\}$, $s\in C^{1}(\mathbb{R}^{n};\mathbb{R}^{k})$ is a sliding set of the system (21) then the equivalent control on $\mathcal{S}$ is defined as a solution of the algebraic equation [Utkin1992:Book]:
$\tfrac{\partial s(x)}{\partial x}f(t,x,u_{\rm eq})=\mathbf{0}.\vskip-2.84526pt$ (24)
Let this algebraic equation have a unique solution $u_{\rm eq}(t,x)$. If $u_{\rm eq}(t,x)\in U(t,x)$ for all $x\in\mathcal{S}$ and for almost all $t\in\mathbb{R}$ and $f_{0}(t,x):=f(t,x,u_{eq}(t,x))$ satisfies Carathéodory existence theorem then the system (11) has an Utkin solution sliding on $\mathcal{S}$ and governed by the ODE $\dot{x}=f_{0}(t,x)$. If
$\det\left(\tfrac{\partial s(x)}{\partial x}b(t,x)\right)\neq 0,\quad\forall x\in\mathcal{S},\quad\forall t\in\mathbb{R},\vskip-2.84526pt$ (25)
for the affine-in-control system from Theorem 10, then [Utkin1992:Book]:
$u_{\rm eq}(t,x)=-\left(\tfrac{\partial s(x)}{\partial x}b(t,x)\right)^{-1}\tfrac{\partial s(x)}{\partial x}a(t,x),\quad x\in\mathcal{S}.\vskip-5.69054pt$ (26)
If $(t,x)\!\mapsto\!\!f^{0}(t,x)\!=\!a(t,x)\!-\!b(t,x)\!\left(\!\frac{\partial s(x)}{\partial x}b(t,x)\!\right)^{\!\!-\!1}\!\!\frac{\partial s(x)}{\partial x}a(t,x)$ is locally one-sided Lipschitz on $\mathcal{S}$ continuous function then any Filippov solution belonging to the sliding set is unique. This does not mean all solutions with $x_{0}\in\mathcal{S}$ are unique, since $\mathcal{S}$ may not be a strongly positively invariant set of the system.
3 Discontinuous Integro-differential equations
Let us consider the integro-differential equation
$\dot{x}(t)\!=\!f(t,x(t))\!+\!\int^{t}_{t^{0}}\!\Phi(t,\tau)\tilde{f}(\tau,x(\tau))d\tau,\quad t>t^{0},\vskip-2.84526pt$ (27)
where the functions $\Phi:\Delta\to\mathbb{R}^{n\times n}$ with $\Delta=\{(t,\tau)\in\mathbb{R}\times\mathbb{R}:t\geq\tau\}$ and $f,\tilde{f}:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ are measurable.
Assumption 4.
Let $\Phi$ be locally essentially bounded on $\Delta$ and $f,\tilde{f}$ satisfy Filippov Condition.
3.1 Well-posed IDEs
Continuous integro-differential equations of the form (27) are well studied in the literature (see, e.g., [Lakshmikantham_etal1995:Book]). Below we use the notation $\mathcal{I}=[t^{0},t^{0}+\alpha)$ with $\alpha>0$ or $\mathcal{I}=[t^{0},+\infty)$.
Definition 5 (Classical solution of IDE).
A function $x\in C^{1}(\mathcal{I};\mathbb{R}^{n})$ with $x(t^{0})=x^{0}$ is said to be a classical solution of the IVP (27), (12) if it satisfies (27) everywhere on $\mathcal{I}$.
The existence theorem of classical solutions of the integro-differential equation (27) repeats the Peano existence theorem.
Theorem 11 (see, e.g., [Lakshmikantham_etal1995:Book]).
If $f$, $\tilde{f}$, $\Phi$ are continuous functions then the IVP (27), (12) has a classical solution defined, at least, locally in the forward time.
Carathéodory solutions of integro-differential equations are defined similarly to solutions of ODEs [OReganMeehan1998:Book].
Definition 6 (Carathéodory solution of IDE).
An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ with $x(t^{0})=x^{0}$ is said to be a Carathéodory (or strong) solution of the IVP (27), (12) if it satisfies (27) almost everywhere on $\mathcal{I}$.
Carathéodory solutions of IDEs are well-studied too [OReganMeehan1998:Book].
Theorem 12.
Let Assumption 4 hold. If $f~\tilde{f}$ satisfy Carathéodory Condition then the IVP (27), (12) has a Carathéodory solution defined, at least, locally in the forward time.
Similarly to ODEs, a Lipschitz condition is sufficient for uniqueness of solutions of IDEs (see, e.g., [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book]).
Theorem 13.
Let Assumption 4 be fulfilled and $f$, $\tilde{f}$ satisfy Carathéodory Condition. If $f$ is locally one-sided Lipschitz and $\tilde{f}$ is locally Lipschitz, then the IVP (27), (12) has a unique solution defined, at least, locally in the forward time.
3.2 Regularization of ill-posed IDEs by Filippov method
If $f,\tilde{f}$ are discontinuous in $x$ functions then the IVP (27), (12) may be ill-posed. Let us regularize it by Filippov method.
Definition 7 (Filippov solution of IDE).
An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ is said to be a Filippov solution of the IVP (27), (12) if $x(t^{0})=x^{0}$ and, almost everywhere on $\mathcal{I}$, it satisfies the integral differential inclusion [^4] The integral of the set-valued function is understood in the sense of Aumann [Aumann1965:JMAA].
$\begin{split}\dot{x}(t)\in&K[f](t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)K[\tilde{f}](\tau,x(\tau))d\tau.\vskip-2.84526pt\end{split}$ (28)
If $\Phi=\mathbf{0}$ or $\tilde{f}$ is independent of $x$, the definition is equivalent to the definition of Filippov solution of discontinuous ODE.
Theorem 14.
If Assumption 4 is fulfilled, then for any $x^{0}\in\mathbb{R}^{n}$ and any $t^{0}\in\mathbb{R}$, the IVP (27), (12) has a Filippov solution defined, at least, locally in the forward time.
Proof.
1) The set-valued mappings $K[f]$ and $K[\tilde{f}$] are nonempty-valued, convex-valued, compact-valued and upper semi-continuous in $x$ for almost all $t$ (see, e.g., [Filippov1988:Book, page 85]).
Let us define the functions $f^{\eta},\tilde{f}^{\eta}:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}$ as follows
$f^{\eta}(t,x)\!=\!\tfrac{1}{\eta}\!\!\int\limits_{\mathcal{B}(\eta)}\!\!\!f(t,x+y)dy,\;\;\;\tilde{f}^{\eta}(t,x)\!=\!\tfrac{1}{\eta}\!\!\int\limits_{\mathcal{B}(\eta)}\!\!\tilde{f}(t,x+y)dy,\vskip-2.84526pt$
where $\eta>0$. In [Filippov1988:Book, Theroem 8, page 85] it is shown that these functions satisfy Carathéodory Condition and
[…]
for all $x\in\mathbb{R}^{n}$ and almost all $t\in\mathbb{R}$. By Theorem 12, the IDE
$\dot{x}^{\eta}(t)=f^{\eta}(t,x)+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{f}^{\eta}(\tau,x^{\eta}(\tau))d\tau\vskip-2.84526pt$ (29)
has a Carathéodory solution. By definition of Aumann’s Integral [Aumann1965:JMAA], for almost all $t\in\mathcal{I}$, this solution satisfies
$\dot{x}^{\eta}(t)\in F^{\eta}(t,x^{\eta}(t))+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{F}^{\eta}(\tau,x^{\eta}(\tau))d\tau.\vskip-2.84526pt$ (30)
2) Since $\Phi$ is locally essentially bounded on $\Delta$ then, for any $\alpha>0$ there exists $\overline{\Phi}>0$ such that
$\operatorname*{ess\,sup}_{t_{0}\leq\tau\leq t\leq t_{0}+\alpha}\|\Phi(t,\tau)\|\leq\overline{\Phi}.\vskip-2.84526pt$
Let $\bar{\eta}=\min\{1/4,1/(4\overline{\Phi}\alpha)\}$. Since, by Filippov Condition, for any $\alpha>0$ and any $r>0$ there exist a Lebesgue integrable function $m:[t^{0},t^{0}+\alpha]\to\mathbb{R}_{+}$ such that $\|f(t,x)\|\leq m(t)$ and $\|{\tilde{f}}(t,x)\|\leq m(t)$ for almost all $t\in[t^{0},t^{0}+\alpha]$ and almost all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$ then (see, e.g., [Filippov1988:Book, page 76])
$K[f](t,x)\subset\mathcal{B}(m(t)),\quad K[\tilde{f}](t,x)\subset\mathcal{B}(m(t))\vskip-2.84526pt$
for almost all $t\in\mathcal{I}=[t^{0},t^{0}+\alpha]$ and for all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$, and (see, e.g., [Filippov1988:Book, page 82])
$F^{\eta}(t,x)\subset\mathcal{B}(m(t)+\eta),\quad\tilde{F}^{\eta}(t,x)\subset\mathcal{B}(m(t)+\eta)\vskip-2.84526pt$
for almost all $t\!\in\!\mathcal{I}$, all $x\!\in\!x^{0}+\mathcal{B}(r+\bar{\eta})$, all $\eta\!\in\!(0,\bar{\eta})$. Hence,
$\|\dot{x}^{\eta}(t)\|\leq m(t)+\eta+\overline{\Phi}\int^{t}_{t^{0}}m(\tau)+\eta\,d\tau\vskip-2.84526pt$
for almost all $t\in\mathcal{I}$ and all $\eta\in(0,\bar{\eta}]$. Since $m$ is Lebesgue integrable on $\mathcal{I}$ then $M:=\int^{t^{0}+\alpha}_{t^{0}}m(\tau)d\tau<+\infty$ and for any $\epsilon>0$ there exist $\delta\in(0,\epsilon\min\{1,1/(4\overline{\Phi}M)\})$ such that for all $a,b]\subset\mathcal{I}:|b-a|\leq\delta$ and all $\eta\in\left(0,\bar{\eta}\right)$ it holds
$\begin{split}\|x^{\eta}(b)-x^{\eta}(a)\|\leq&\left\|\int^{b}_{a}{\dot{x}}^{\eta}(s)ds\right\|\leq\int^{b}_{a}\|{\dot{x}}^{\eta}(s)\|ds\\ \leq&\int^{b}_{a}m(s)ds+(b-a)(\eta+\overline{\Phi}M+\overline{\Phi}\alpha\eta)\\ \leq&\frac{\epsilon}{4}+\delta\eta+\delta\overline{\Phi}M+\delta\overline{\Phi}\alpha\eta\leq\epsilon.\vskip-2.84526pt\end{split}$
This means that the family of function $\{x^{\eta}\},\eta\in(0,\bar{\eta}]$ is equicontinuous and uniformly bounded for all $t\in\mathcal{I}$,
$\|x^{\eta}(t)\|\leq\|x_{0}\|+\int^{t}_{t_{0}}\|\dot{x}^{\eta}(\tau)\|d\tau\leq M+\alpha\eta+\overline{\Phi}(M\alpha+\alpha^{2}\eta),\vskip-2.84526pt$
so, by Arzelà–Ascoli theorem, it contains a sequence $x^{\eta_{k}}$ uniformly convergent to a continuous function $x$ as $\eta_{k}\to\mathbf{0}$. The above inequalities imply that $x$ is absolutely continuous.
3) The set-valued mappings $F^{\eta}$ and $\tilde{F}^{\eta}$ are nonempty-valued, compact-valued, convex-valued and upper semi-continuous continuous in $x$ for almost all $t$ (see, e.g., [Filippov1988:Book, page 86]). For any solution $x^{\eta}$ of the IDE (29), the set-valued mappings $t\to F^{\eta}(t,x^{\eta}(t))$ and $\tau\to\Phi(t,\tau)\tilde{F}^{\eta}(t,x^{\eta}(t))$ are compact-valued, convex-valued, measurable and bounded by the integrable function $t\mapsto\bar{\Phi}(m(t)+\eta)$. By [Aumann1965:JMAA, Theorems 1, 2 and 4] [^5] The requirement of the paper [Aumann1965:JMAA] about Borel graph of the integrated function has been relaxed (see, e.g., [CastaingValadier1977:Book, Chapter 3]) using Kuratowski–Ryll-Nardzewski measurable selection theorem published in the same year [KuratowskiRyll-Nardzewski1965:BAPC]., the right-hand side of (30)
$Q^{\eta}(t):=F^{\eta}(t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{F}^{\eta}(\tau,x(\tau))d\tau\subset\mathbb{R}^{n}\vskip-2.84526pt$
is convex-valued and compact-valued for almost all $t$. Moreover, by construction, for almost all $t$ and for all $x$, we have $K[f](t,x)\subset F^{\eta}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{\eta}(t,x)$, $\forall\eta\in[0,\overline{\eta}]$, and $F^{\eta}(t,x)\to K[f](t,x)$, $\tilde{F}^{\eta}(t,x)\to K[\tilde{f}](t,x)$ as $\eta\to 0$, where the limit is understood in the Hausdorff topology (see, e.g., [Filippov1988:Book, §5]). By [Aumann1965:JMAA, Theprem 5], for almost all $t\in\mathcal{I}$, we conclude that $Q^{\eta_{k}}(t)\to Q(t)$ as $\eta_{k}\to 0$, where
$Q(t)=K[f](t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)K[\tilde{f}](\tau,x(\tau))d\tau\subset\mathbb{R}^{n}\vskip-2.84526pt$
is a convex and compact set. Hence, for any $v\in\mathbb{R}^{n}$ and for almost all $t\in\mathcal{I}$ we have
$v^{\top}\dot{x}^{\eta_{k}}(t)\leq\sup_{\ell\in Q^{\eta_{k}}(t)}v^{\top}\ell$ and
$\phi(t):=\limsup_{\eta_{k}\to 0}v^{\top}\dot{x}_{\eta_{k}}(t)\leq\sup_{\ell\in Q(t)}v^{\top}\ell,\vskip-5.69054pt$ (31)
Moreover, for any $a,b]\subset\mathcal{I}$, it holds
$v^{\top}(x^{\eta_{k}}(b)-x^{\eta_{k}}(a))\!=\!\int^{b}_{a}v^{\top}\dot{x}^{\eta_{k}}(s)ds\!\leq\!\int^{b}_{a}\sup\limits_{j\geq k}v^{\top}x^{\eta_{j}}(s)ds.\vskip-2.84526pt$
Since $v^{\top}(x^{\eta_{k}}(b)-x^{\eta_{k}}(a))=\int^{b}_{a}v^{\top}\dot{x}^{\eta_{k}}(s)ds\to v^{\top}(x(b)-x(a))=\int^{b}_{a}v^{\top}\dot{x}(s)ds$ as $\eta^{k}\to 0$ and $\sup_{j\geq k}v^{\top}x^{\eta_{j}}(s)$ does not increase as $k$ grows, then we derive the inequality
$\int^{b}_{a}v^{\top}\dot{x}(s)ds=v^{\top}(x(b)-x(a))\leq\int^{b}_{a}\phi(s)ds\vskip-2.84526pt$
by tending $\eta_{k}\!\to\!0$. The segment $[a,b]$ is arbitrary selected, so
$v^{\top}\dot{x}(t)\leq\sup_{\ell\in Q(t)}v^{\top}\ell,\quad\forall v\in\mathbb{R}^{n},\vskip-2.84526pt$
almost everywhere on $t^{0},t^{0}+\alpha]$. This means $\dot{x}(t)\in Q(t)$ for almost all $t\in\mathcal{I}$. The proof is complete. ∎
Under some assumptions, the integro-differential equation (27) can be interpreted as a functional differential equation of retarded type [Hale1977:Book], [Fridman2014:Book]. Indeed, if, there exist $d>0$ such that $\Phi(t,\tau)=0$ for $t-\tau>d$, then $\dot{x}(t)$ depends only on $x(\tau)$ with $\tau\in[\max\{t^{0},t-d\},t]$, $t\geq t^{0}$. Generalized solutions of discontinuous retarded functional differential equations are introduced in [KolmanovskiiMyshkis1992:Book, page 58] by means of a limiting procedure in a function space, which does not have such a simple formalization by the Filippov regularization (15). A study of links between the generalized solutions and the Filippov solutions is an interesting problem for the future research.
Theorem 15 (On continuous dependence on parameters).
[…]
$E(t,x)\!=\!\left\{\!z\!\in\!\mathbb{R}^{n}:\!\begin{smallmatrix}\|x-z\|\leq m_{\mu}(t),\,\|f_{\mu}(t,x)-f(t,z)\|\leq m_{\mu}(t)\\ \text{ and }\|\tilde{f}_{\mu}(t,x)-\tilde{f}(t,z)\|\leq m_{\mu}(t)\end{smallmatrix}\!\right\}\!\!\!\vskip-2.84526pt$ (32)
is of positive measure for almost all $t\in(-\alpha,\alpha)$ and almost all $x\!\in\!\!B(r)$. Let $\sup_{\mu\in\mathcal{M}}\|m_{\mu}\|_{L^{\infty}}<+\infty$ and $\|m_{\mu}\|_{L^{1}}\to 0$ as $\mu\to\mu_{0}$. If, for a given $t^{0}\in\mathbb{R}$ and a given $x^{0}\in\mathbb{R}^{n}$, all Filippov solutions of the IVP (27), (12) exist on $t^{0},t^{0}+T]$ then, for any $\varepsilon>0$, there exists $\delta\in(0,\varepsilon)$ such that any Filippov solution $x_{\mu}$ of the IVP
$\dot{x}_{\mu}(t)\!=\!f_{\mu}(t,x_{\mu}(t))\!+\!\!\int^{t}_{t^{0}_{\mu}}\!\!\Phi(t,\tau)\tilde{f}_{\mu}(\tau,x_{\mu}(\tau))d\tau,\;\;x_{\mu}(t^{0}_{\mu})\!=\!x^{0}_{\mu}\vskip-2.84526pt$ (33)
with $t>t^{0}_{\mu}\in[t^{0},t^{0}+\delta]$, $x_{\mu}(t^{0}_{\mu})=x^{0}_{\mu}\in x_{0}+\mathcal{B}(\delta)$ and $\|m_{\mu}\|_{L^{1}(t^{0},t^{0}+T)}\!\leq\!\delta$ differs from some solution $x$ of the IVP (27), (12) for less than $\varepsilon$, i.e., $\sup\limits_{t\in[t^{0}+\delta,t^{0}+T]}\|x(t)-x_{\mu}(t)\|\leq\epsilon$.
Proof.
Let $\bar{r}>0$ be such that any Filippov solution of (27), (12) belongs to the ball $x^{0}+\mathcal{B}(\bar{r})$ for all $t\in[t^{0},t^{0}+T]$. Let us study the case $t^{0}_{\mu}=t^{0}$ and define
$F^{*}_{\mu}(t,x)=\overline{\mathrm{co}}\bigcup_{y\in\mathcal{B}(m_{\mu}(t)),z\in\mathcal{B}(m_{\mu}(t))}K[f](t,x+y)+z$
$\tilde{F}^{*}_{\mu}(t,x)=\overline{\mathrm{co}}\bigcup_{y\in\mathcal{B}(m_{\mu}(t)),z\in\mathcal{B}(m_{\mu}(t))}K[\tilde{f}](t,x+y)+z.\vskip-2.84526pt$
For almost all $t\in(t^{0},t^{0}+T)$ and almost all $x\in x^{0}+\mathcal{B}(\bar{r})$, by construction $K[f](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$, and $K[f_{\mu}](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$ due to positive measure of $E(t,x)$ (see, [Filippov1988:Book, page 98] for more details). Since $r_{\mathcal{M}}=\sup_{\mu\in\mathcal{M}}\|m_{\mu}\|_{L^{\infty}(t^{0},t^{0}+T)}<+\infty$ then, due to Filippov Condition, there exists $m\in L^{1}((t^{0},t^{0}+T),\mathbb{R}_{+})$ such that $\sup_{y\in F^{*}_{\mu}(t,x)}\|y\|\leq m(t)+m_{\mu}(t)$ for almost all $t\in(t^{0},t^{0}+T)$, for all $x\in x^{0}+\mathcal{B}(\bar{r}+r_{\mathcal{M}})$ and for all $\mu\in\mathcal{M}$. Hence, any Filippov solution of (33) satisfies
$\dot{x}_{\mu}(t)\in F^{*}_{\mu}(t,x_{\mu}(t))+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{F}^{*}_{\mu}(\tau,x_{\mu}(\tau))d\tau.\vskip-2.84526pt$ (34)
and $\|\dot{x}_{\mu}(t)\|\in m(t)+m_{\mu}(t)+\overline{\Phi}\int^{t}_{t^{0}}m(\tau)+m_{\mu}(\tau)\,d\tau$ almost everywhere, where $\overline{\Phi}\geq 0$ is defined in the proof of Theorem 14. Repeating the step 2) of the proof of Theorem 14, we show that the family of solutions of (33) with $\mu\in\mathcal{M}$ is uniformly bounded and equicontinuous. Since, by assumption, $\|m_{\mu}\|_{L^{1}}\to 0$ as $\mu\to\mu_{0}$, then, repeating the step 3) of the proof of Theorem 14, we conclude that, for any uniformly convergent sequence of solutions $x_{\mu_{i}}\to x_{\mu_{0}}$ as $\mu_{i}\to\mu_{0}$, the limit $x_{\mu_{0}}$ is a Filippov solution of the IVP (27), (12). Applying [Filippov1988:Book, Lemma 6, page 10] we complete the proof. ∎
Continuous dependence of solutions of IDEs on initial data (as in Theorem 3) follows from the above result.
3.3 Uniqueness of Filippov solutions of IDEs
For continuous $\tilde{f}$, the uniqueness analysis of Filippov solutions repeats the analysis of Charathéodory solutions.
Theorem 16.
Under Assumption 4, if $f$ is locally one-sided Lipschitz and $\tilde{f}$ is locally Lipschitz then the IVP (27), (12) has
a unique solution defined, at least, locally in the forward time.
Proof.
If $x_{1},x_{2}$ are two solutions of the IVP (27), (12) then, for $e(t)=x_{1}(t)-x_{2}(t)$ and for almost all $t$, it holds
$\begin{split}\tfrac{d\|e(t)\|^{2}}{dt}=&2e^{\top}(t)(f_{1}(t)-f_{2}(t))\\ &+\!2\!\!\int^{t}_{t^{0}}\!\!\!e^{\top}\!(t)\Phi(t,\tau)\tilde{f}(\tau,x_{1}(\tau)\!-\!\tilde{f}(\tau,x_{2}(\tau))d\tau\\ \leq&2\ell(t)\|e(t)\|^{2}+2\bar{\Phi}\|e(t)\|\int^{t}_{t^{0}}\tilde{\ell}(\tau)\|e(\tau)\|d\tau\\ \leq&(2\ell(t)+\bar{\Phi}^{2})\|e(t)\|^{2}+\|\tilde{\ell}\|_{L^{2}}^{2}\int^{t}_{t^{0}}\|e(\tau)\|^{2}d\tau,\end{split}\vskip-2.84526pt$ (35)
where $f_{1}(t)\in K[f](t,x_{1}(t)),f_{2}(t)\in K[f](t,x_{2}(t))$. One-Sided Lipschitz Condition for $K[f]$ [Filippov1988:Book, page 106], Lipschitz Condition for $\tilde{f}$ and Cauchy–Schwartz inequality in $L^{2}$ have been utilized on the last steps. Since, $e(t^{0})=0$ then the obtained inequality implies e(t)=0e(t)\!=\!0 for $t\!>\!t^{0}$ (see [Pachpatte1975:JMAA]). ∎
For discontinuous $\tilde{f}$, the uniqueness analysis is challenging.
Let us denote $F_{\mathcal{S}}(t,x,\gamma)\!=\!\left(K[f](t,x)+\gamma\right)\,\cap\,\mathcal{C}_{\mathcal{S}}(x)$, where $\gamma\in\mathbb{R}^{n}$ and $\mathcal{S}\subset\mathbb{R}^{n}$ is the discontinuity set of $f$.
Theorem 17 (Uniqueness of solution in sliding mode case).
Let Assumption 4 be fulfilled. Let $\tilde{\mathcal{S}}\subset\mathbb{R}^{n}$ (resp., $\mathcal{S}\subset\mathbb{R}^{n}$) be a discontinuity set of the piece-wise (resp., one-sided) Lipschitz continuous function $\tilde{f}$ (resp., $f$). Let a measurable function $(t,x,\gamma)\in\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\mapsto f^{0}(t,x,\gamma)\in\mathbb{R}^{n}$ be locally one-sided Lipschitz in $x$ and locally Lipschitz in $\gamma$. If
  • 1)
    $f$ satisfies Assumption 1 and $\mathcal{S}\cap\tilde{\mathcal{S}}$$\emptyset$;
  • 2)
[…]
  • 3)
    $K[f](t,x)\cap\mathcal{C}_{\tilde{\mathcal{S}}}(x)\!=\!\emptyset$ for all $x\!\in\!\tilde{\mathcal{S}}$ and almost all $t\!\in\!\mathbb{R}$;
then then the IVP (27), (12) has a Filippov solution being unique, at least, locally in the forward time.
Proof.
Let us study the case $x^{0}\in\mathcal{S}$. This means that $x^{0}\in\partial G^{j}$ for all $j\in\mathcal{N}(x^{0})$, where, as before, $\mathcal{N}:\mathbb{R}^{n}\to 2$ is the indicator function (i.e., $j\in\mathcal{N}(x_{0})$ if $x_{0}\in\overline{G}^{j}$). Since the function $f$ is piece-wise continuous then there exists $\bar{\epsilon}>0$ such that $\mathcal{N}(x^{0}+\delta)\subset\mathcal{N}(x^{0})$ for $x_{0}+\delta\in S:\|\delta\|\leq\bar{\epsilon}$.
By Assumption 1, the vector $f^{j}(t,x^{0})$ belongs to the set $\mathbb{R}^{n}\backslash\mathcal{C}_{\bar{G}^{j}}(x^{0})$ for all $t\in\mathbb{R}$. Since $\mathcal{C}_{\bar{G}^{j}}(x^{0})$ is a positive closed cone then its complement $\mathbb{R}^{n}\backslash\mathcal{C}_{\bar{G}^{j}}(x^{0})$ is a positive open cone. Hence, taking into account, $f^{j}\in C(\mathbb{R}\times\mathbb{R}^{n};\mathbb{R}^{n})$ we conclude that there exists $\epsilon^{j}>0$ such that
$f^{j}(t^{0}+s,x^{0}+B(\epsilon^{j}))+B(\epsilon^{j})\subset\left(\mathbb{R}^{n}\backslash{\mathcal{C}}_{\bar{G}^{j}}(x^{0})\right)\vskip-2.84526pt$
for all $s\in(0,\epsilon^{j})$. Let $\epsilon=\min\left\{\bar{\epsilon},{\hat{\epsilon}},\min_{j\in\mathcal{N}(x^{0})}\{\epsilon^{j}\}\right\}$. Repeating considerations of the proof of Theorem 8, we conclude that any solution $x^{\epsilon}$ of the IVP
$\dot{x}^{\epsilon}(t)\in K[f](t,x^{\epsilon}(t))+B(\epsilon),\quad x^{\epsilon}(t^{0})=x^{0}\in\mathcal{S}\vskip-2.84526pt$ (36)
belongs to $\mathcal{S}$ as long as $x^{\epsilon}(t)\in x^{0}+B(\epsilon)$. For any continuous function $t\in[t^{0},t^{0}+\epsilon]\to x(t)\in\mathbb{R}^{n}$, the set-valued function
$t\in[t^{0},t^{0}+\epsilon]\mapsto\tilde{Q}(t):=\int^{t}_{t^{0}}\Phi(t,\tau)K[\tilde{f}](\tau,x(\tau))d\tau\subset\mathbb{R}^{n}\vskip-2.84526pt$
is bounded as follows (see, the proof of Theorem 14)
$\sup_{y\in\tilde{Q}(t)}\|y\|\leq\overline{\Phi}\int^{t}_{t^{0}}m(\tau)+\eta\,d\tau\vskip-2.84526pt$
for some $\eta>0,\overline{\Phi}>0$ and some integrable function $m:[t^{0},t^{0}+\epsilon]\to\mathbb{R}$. This means that $\exists t^{\prime}\in(0,\epsilon)$ such that
$\sup_{y\in\tilde{Q}(t)}\|y\|\leq\epsilon,\quad\forall t\in[t^{0},t^{0}+t^{\prime}].\vskip-2.84526pt$
Any solution the system (27) with $x(t^{0})=x^{0}\in\mathcal{S}$ is a solution of (36) for $t\in[t^{0},t^{0}+t^{\prime}]$. Since $\mathcal{S}\cap\tilde{\mathcal{S}}$$\emptyset$ and $\tilde{f}$ satisfies Lipschitz condition on any compact from $\mathcal{S}$ then the IDE
$\dot{x}(t)\!=\!f^{0}\left(t,x(t),\gamma(t)\right)+\gamma(t),\quad\gamma(t)\!=\!\int^{t}_{t^{0}}\!\!\!\Phi(t,\tau)\tilde{f}(\tau,x(\tau))d\tau\vskip-2.84526pt$
[…]
Theorem 17 asks $\tilde{f}$ to be continuous on the discontinuity set of $f$. For IDEs with sliding modes, this condition cannot be violated to ensure the uniqueness of Filippov solutions.
Example 5.
A Filippov solution of the IVP
$\left\{\begin{split}\dot{x}_{1}(t)=&\;u(x(t)),\quad\quad\quad\quad\quad\quad\quad\quad\quad x_{1}(0)=0,\\ \dot{x}_{2}(t)=&\int^{t}_{0}{\sin(t+\tau)}u(x(\tau))d\tau,\quad\quad x_{2}(0)=0,\\ u(x)=&-\mathrm{sign}(x_{1}),\quad\quad\quad\;\;x=(x_{1},x_{2})^{\top}\in\mathbb{R}^{2}\end{split}\right$. (37)
is $x_{1}(t)=0$ and $x_{2}(t)=\int^{t}_{0}\int^{s}_{0}{\sin(s+\tau)}q(\tau)d\tau ds$, where $q:\mathbb{R}\mapsto[-1,1]$ is any measurable function.
The IDE (27) without sliding mode may have unique solutions if an intersection of discontinuity sets of $f$ and $\tilde{f}$ is non-empty.
Theorem 18 (Uniqueness of solution in switching case).
Let Assumption 4 be fulfilled. Let $\tilde{\mathcal{S}}\subset\mathbb{R}^{n}$ (resp., $\mathcal{S}\subset\mathbb{R}^{n}$) be a discontinuity set of the piece-wise (resp., one-sided) Lipschitz continuous function $\tilde{f}$ (resp., $f$). If $\tilde{\mathcal{S}}\subset\mathcal{S}$ and $f$ satisfies Assumption 2 then the IVP (27), (12) has a Filippov solution being unique, at least, locally in the forward time.
Proof.
For $x_{0}\notin\mathcal{S}$, Filippov solution of the IVP (27), (12) coincides Carathéodory solution as long as $x(t)\notin\mathcal{S}$. This solution is unique due to Theorems 12 and 13. Since $K[f](t,x)$ and $\mathcal{C}_{\mathcal{S}}(x)$ are compact sets for all $x$ and for almost all $t$, then Assumption 2 implies that for any $x\in\mathcal{S}$ there exists $\hat{\epsilon}>0$ such that $K[f](t,x)+\gamma\cap\mathcal{C}_{\mathcal{S}}(x)=\emptyset$ for all $\gamma\in\mathbb{R}^{n}:\|\gamma\|\leq\hat{\epsilon}$. A Filippov solution $x$ of the IVP (27), (12) initiated at $x^{0}\in\mathcal{S}$ immediately enters in a domain $G^{j}$, where the solution is unique due to Theorems 12 and 13. ∎
As a straightforward consequence of the two above theorems we deduce the following corollary.
Corollary 2.
Let Assumption 4 hold. Let $\tilde{\mathcal{S}}\subset\mathbb{R}^{n}$ (resp., $\mathcal{S}\subset\mathbb{R}^{n}$) be a discontinuity set of the piece-wise (resp., one-sided) Lipschitz continuous function $\tilde{f}$ (resp., $f$). Let a measurable function $(t,x,\gamma)\in\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\mapsto f^{0}(t,x,\gamma)\in\mathbb{R}^{n}$ be locally one-sided Lipschitz in $x$, locally Lipschitz in $\gamma$. If
  • 1)
    $f$ satisfies Assumption 3 with $\mathcal{M}\subset\mathcal{S}:\mathcal{M}\cap\tilde{\mathcal{S}}$$\emptyset$;
  • 2)
[…]
  • 3)
    $K[f](t,x)\cap\mathcal{C}_{\tilde{\mathcal{S}}}(x)\!=\!\emptyset$ for all $x\!\in\!\tilde{\mathcal{S}}$ and almost all $t\!\in\!\mathbb{R}$;
then the IVP (27), (12) has a Filippov solution being unique, at least, locally in the forward time.
Remark 1 (On explicit Euler method for IDEs).
If $f,\tilde{f}$ are piece-wise locally Lipschitz continuous functions in $x$ then approximate (numerical solution) of the IDE (27) can be found using the explicit Euler method
$\begin{split}x_{k+1}=&x_{k}+hf(t^{0}+kh,x_{k})+hI_{k},\\ I_{k}=&h\sum^{k}_{i=0}\Phi(t^{0}+kh,t^{0}+ih)\tilde{f}(t^{0}+ih,x_{i}),\vskip-2.84526pt\end{split}$ (38)
where $h>0$ is the sampling period, $k=0,1,2,$..., $x_{k}$ approximates $x(t^{0}+kh)$ and $I_{k}$ approximates the integral $\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{f}(\tau,x(\tau))d\tau$ by means of the method of rectangles. Indeed, since, on continuity domains, both functions $f,\tilde{f}$ are locally Lipschitz continuous in $x$, any solution $x:\mathcal{I}\to\mathbb{R}^{n}$ of IDE (27) belonging to this domain is unique and the sequence of the Euler approximates
$x_{h}(t)\!=\!\tfrac{t^{0}\!+(k+1)h-t}{h}x_{k}+\tfrac{t-t^{0}\!-kh}{h}x_{k+1},t\!\in\![t^{0}\!+kh,t^{0}\!+(k+1)h],$
(being uniformly bounded and equicontinuous on $\mathcal{I}$) converges to $x$ uniformly on $\mathcal{I}$ as $h\to 0$. This conclusion remains valid if the discontinuous IDE (27) has no sliding mode. In the sliding mode case, similarly to ODEs, the explicit Euler method correctly approximates the Filippov solution of the IDE away an $\epsilon$-neighborhood of a sliding set $\mathcal{S}$, where $\epsilon\to 0$ as $h\to 0$. A consistent discrete-time approximation of the motion on the sliding surface requires development of advanced numerical schemes [AcaryBrogliato2010:SCL], [Polyakov_etal2019:SIAM_JCO]. This goes out of the scope of the paper.
3.4 Equivalent control method for systems modeled by IDEs
Let us consider a control system modeled by the IDE (8) with measurable functions $f$ and $\tilde{f}$ being continuous in $(x,u)\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ for almost all $t\in\mathbb{R}$.
Definition 8 (Utkin solution of IDE).
An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ is said to be an Utkin solution of the first kind (resp., of the second kind) of the IVP (8), (12) if $x(t^{0})=x^{0}$ and there exists a measurable function $u_{\rm eq}:\mathcal{I}\to\mathbb{R}^{m}$ such that $u_{\rm eq}(t)\in U(t,x(t))$ and
$\dot{x}(t)\!=\!f(t,x(t),u_{\rm eq}(t))\!+\!\!\int^{t}_{t^{0}}\!\!\!\Phi(t,\tau)\tilde{f}(\tau,x(\tau),u_{\rm eq}(\tau))d\tau,\!\vskip-2.84526pt$ (39)
almost everywhere on $\mathcal{I}$, where $U=K[u]$ is defined by (15) (resp., $U=(K[u_{1}],...,K[u_{m}])^{\top}$ for $u=(u_{1},u_{2},...,u_{m})^{\top}$).
Similarly to ODEs, the set of Utkin solutions of the first kind is a subset of the set of the Utkin solutions of the second kind, since $K[u](t,x(t))\subset(K[u_{1}](t,x(t)),...,K[u_{m}(t,x(t))])^{\top}$. They always coincide for single input systems (i.e., if $m=1$). Utkin solutions always exist for affine-in-control IDEs.
Theorem 19.
Let $f$ and $\tilde{f}$ in (8) have the form
$f(t,x,u)\!=\!a(t,x)\!+\!b(t,x)u,\quad\tilde{f}(t,x,u)\!=\!\tilde{a}(t,x)\!+\!\tilde{b}(t,x)u,\vskip-2.84526pt$
where $t\in\mathbb{R},x\in\mathbb{R}^{n},u\in\mathbb{R}^{m}$ and measurable functions $a,\tilde{a}:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, $b,\tilde{b}:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ are locally essentially bounded and continuous in $x$. If a measurable function $u:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is locally essentially bounded then, for any $t^{0}\in\mathbb{R}$ and any $x^{0}\in\mathbb{R}^{n}$, the IVP (8),(12) has an Utkin solution defined, at least, locally in the forward time. For the affine-in-control system, any Utkin solution of the first kind is a Filippov solution.
Proof.
Let us denote $f^{\eta}\!=\!a+bu^{\eta},\tilde{f}^{\eta}\!=\!\tilde{a}+\tilde{b}u^{\eta}$,
$u^{\eta}(t,x)\!=\!\tfrac{1}{\eta}\int_{\|y\|\leq\eta}\!\!\!\!\!\!u(t,x+y)dy,\vskip-2.84526pt$
$U^{\eta}(t,x)=\overline{\mathrm{co}}\bigcup_{y,z\in\mathcal{B}(\eta)}U(t,x+y)+z,\vskip-2.84526pt$
[…]
where $U=K[u]$ or $U=(K[u_{1}],....,K[u_{m}])$. By construction, $u^{\eta}(t,x)\in U^{\eta}(t,x)$, $f^{\eta}(t,x)\in F^{\eta}(t,x)$, $\tilde{f}^{\eta}(t,x)\in\tilde{F}^{\eta}(t,x)$ for almost all $t$ and all $x$. Since $a,\tilde{a},b,\tilde{b}$ are continuous in $x$ then, for any $\eta>0$, the functions $f^{\eta}$ and $\tilde{f}^{\eta}$ satisfy Caratheodory Condition and the set-valued mappings $F^{\eta}$ and $\tilde{F}^{\eta}$ are compact-valued, convex-valued and upper-semicontinuous in $x$. Let us denote $\xi^{\eta}(t)=\int^{t}_{t^{0}}u^{\eta}(\tau,x^{\eta}(\tau))d\tau$. Repeating the step 2 of the proof of Theorem 14 we conclude that the family of pairs $(x^{\eta},\xi^{\eta})$ is uniformly bounded and equicontinuous on some time interval $\mathcal{I}$ and there exists a sequence $\eta_{i}:\lim_{i\to\infty}\eta_{i}=0$ and a pair of absolutely continuous functions $(x,\xi)$ such that $(x^{\eta_{i}},\xi^{\eta_{i}})\to(x,\xi)$ as $i\to+\infty$ uniformly on $\mathcal{I}$. Since $a,\tilde{a},b,\tilde{b}$ are continuous in $x$, then, repeating the step 3) of the proof of Theorem 14 we derive $\dot{\xi}(t)\in U(t,x(t))$ and $\dot{x}(t)\in a(t,x(t))+b(t,x(t))U(t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{a}(\tau,x(\tau))d\tau+\int^{t}_{t^{0}}\Phi(t,\tau)b(\tau,x(\tau))\dot{\xi}(\tau)d\tau$ for almost all $t\!\in\!\mathcal{I}$. So, $x$ is an Utkin solution with $u_{\rm eq}\!=\!\dot{\xi}$. Due to continuity of $f$ and $\tilde{f}$ in $x$, it holds $K[f]=a+bK[u]$ and $K[\tilde{f}]=\tilde{a}+\tilde{b}K[u]$. So, any Utkin solution of the first kind is a Filippov solution. ∎
A set of Filippov solutions of the affine-in-control discontinuous IDE may be larger than a set of Utkin solutions of the first kind. This is a specific feature of IDEs. For affine-in-control discontinuous ODEs ($\Phi=\mathbf{0}$), these sets are identical.
Example 5 (Continued).
The IVP (37) has the unique Utkin solution of the first kind: $x_{1}(t)=x_{2}(t)=0,\forall t\geq 0$ with the equivalent control $u_{\rm eq}(t)=0$.
At the first sight, in view of the above example, the Utkin solutions of the first kind seems to be more appropriate for control systems design. Indeed, the integral part in IDE (8) contains the control signal defined in past, i.e., for $\tau\leq t$. At time $t$, this control is already determined, so a non-uniqueness of a control signal in past looks artificial. However, Utkin solutions of the second kind (as well-as Filippov solutions) may correspond to a more realistic dynamics of a MIMO system with inputs generated by independent actuators.
Example 6.
The Utking solution of the first kind for
[…]
is unique and given by $x_{1}(t)=x_{2}(t)=0,\forall t\geq 0$. The set of Filippov solutions of the IVP (40) coincides with the sets of Utkin solutions of the second kind of the IVP (40) and coincides with the set of Filippov solutions of the IVP (37) (see, Example 5). If control inputs $u_{1}$ and $u_{2}$ are assumed to be independent (e.g., realized by different actuators), then, in practice, the input signals cannot be identical. In this particular case, the Utkin solution of the second kind (40) provides a more realistic description of the system motion than the Utkin solution of the first kind.
Corollary 3 (On equivalent control on smooth surface).
Let $\mathcal{S}=\{x\in\mathbb{R}^{n}:s(x)=0\},s\in C^{1}(\mathbb{R}^{n};\mathbb{R}^{m})$ be the discontinuity set of the piece-wise Lipschitz continuous function $u:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{m}$. Let conditions of Theorem 19 be fulfilled, the matrix $G(t,x)=\frac{\partial s(x)}{\partial x}b(t,x)$ be non-singular for almost all $t\in\mathbb{R}$ and for all $x\in\mathbb{R}^{n}$ and $(t,x)\to G^{-1}(t,x)$ be locally essentially bounded on $\mathbb{R}\times\mathbb{R}^{n}$.
  • i)
    If $t\in[t^{0},t^{1}]\mapsto x(t)\in\mathbb{R}^{n}$ with $\mathcal{I}=[t^{0},t^{1}]$ is an Utkin solution of (8), (12) such that $x(t)\notin\mathcal{S}$ for $t\in[t^{0},t^{*}]$ and $x(t)\in\mathcal{S}$ for $t\in[t^{*},t^{1}]$ then the equivalent control $u_{\rm eq}:[t^{0},t^{1}]\to\mathbb{R}^{m}$ is given by
[…]
    where $\mathcal{K}:L^{\infty}((t^{*},t^{1});\mathbb{R}^{m})\to L^{\infty}((t^{*},t^{1});\mathbb{R}^{m})$ is a linear bounded operator
    $(\mathcal{K}\tilde{u})(t)\!=\!\!\int^{t}_{t^{*}}\!\!\!G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\tilde{u}(\tau)d\tau,\vskip-2.84526pt$ (42)
    and $g\!\in\!L^{\infty}((t^{*},t^{1});\mathbb{R}^{n})$ is as follows
    $g(t)\!=\!-G^{-1}(t,x(t))\left.\tfrac{\partial s(y)}{\partial y}\right|_{y=x(t)}\!\left(a(t,x(t))\!+\!\tilde{g}(t)\right),\vskip-5.69054pt$
    $\tilde{g}(t)\!=\!\int^{t^{*}}_{t^{0}}\!\!\!\!\!\Phi(t,\tau)\left(a(\tau,x(\tau))\!+\!b(\tau,x(\tau))u(t,x(\tau)\right)d\tau.\vskip-2.84526pt$
  • ii)
    If $a,b,\tilde{a},\tilde{b},\frac{\partial s}{\partial x}$ are locally Lipschitz continuous and $G(t,y_{1})=G(t,y_{2}),\tilde{b}(t,y_{1})=\tilde{b}(t,y_{2})$, $\forall y_{1},y_{2}\in\mathcal{S}$ then any Utkin solution of (8), (12) belonging to $\mathcal{S}$ is unique.
Proof.
i) If $x(t)\!\in\!\mathcal{S}$ then $\frac{ds(x(t))}{dt}\!=\!\left.\frac{\partial s(y)}{\partial y}\right|_{y=x(t)}\dot{x}(t)\!=\!0$ and, taking into account,
$\begin{split}\dot{x}(t)\!=&a(t,x(t))+b(t,x(t))u_{\rm eq}(t)\\ &+\!\int^{t}_{t^{0}}\!\!\!\Phi(t,\tau)\left(\tilde{a}(\tau,x(\tau))\!+\!\tilde{b}(\tau,x(\tau))u_{\rm eq}(\tau)\right)d\tau,\vskip-5.69054pt\end{split}$ (43)
we conclude that, in the sliding mode, $u_{\rm eq}$ is the unique solution of the integral Volterra equation of the second kind
$u_{\rm eq}(t)\!=\!g(t)\!+\!(\mathcal{K}u_{\rm eq})(t),\quad t\in[t^{*},t^{1}].\vskip-2.84526pt$ (44)
This integral equation has a unique solution since
$M\!=\!\!\operatorname*{ess\,sup}_{t^{*}\leq\tau\leq t\leq t^{1}}\|G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\|d\tau\!<\!+\infty.\vskip-5.69054pt$
Indeed, since $\|\mathcal{K}^{n}\|_{\mathcal{L}(L^{\infty};L^{\infty})}\leq\frac{M^{n}(t^{1}-t^{*})^{n}}{n!}$, then the Neumann series $\sum^{\infty}_{i=0}\mathcal{K}^{i}$ is convergent in the Banach space $\mathcal{L}(L^{\infty};L^{\infty})$ of bounded linear operators $L^{\infty}\to L^{\infty}$, and $u_{\rm eq}=\sum^{\infty}_{i=0}\mathcal{K}^{i}g$ is the solution of the integral equation (44).
ii) For $t\in[t^{*},t^{1}]$, the dynamics is given by (43). Let $x_{1}(t),x_{2}(t)$ satisfy the above identity for $t\geq t^{0}$, $x_{1}(t)=x_{2}(t)=x(t)\notin\mathcal{S}$ for $t^{0}\leq t<t^{*}$ and $x_{1}(t)\in\mathcal{S},x_{2}(t)\in\mathcal{S}$ for $t\in[t^{*},t^{1}]$. By assumption, the operator $\mathcal{K}$ is independent of $x(t)$. For $j=1,2$, let $g_{i}$ be defined by replacing $x$ with $x_{j}$ in the formula for $g$, respectively. Notice than $\tilde{g}$ is identical for both $x_{1}$ and $x_{2}$. Then, using Lipschitz continuity in $x$ and local essential boundedness in $t$ of the functions $a,b,\tilde{a},\tilde{b},\frac{\partial s}{\partial x}$, for $e(t)=x_{1}(t)-x_{2}(t)$, we derive $\|g_{1}(t)-g_{2}(t)\|\leq\ell_{1}\|e(t)\|$ and $\|Kg_{1}(t)-Kg_{2}(t)\|\leq\ell_{2}\int^{t}_{t^{*}}\|e(\tau)\|d\tau\leq\ell_{2}\sqrt{t-t^{*}}\|e\|_{L^{2}}$ for some $\ell_{1},\ell_{2}\!>\!0$, where $\|e\|^{2}_{L^{2}(t^{*},t)}=\int^{t}_{t^{*}}e(\tau)d\tau$. Moreover, for $i\geq 1$, it holds $\|(K^{i}g_{1})(t)-(K^{i}g_{2})(t)\|\leq\tfrac{(2\ell_{2})^{i}(t-t^{*})^{i}}{(2i-1)!\sqrt{t-t^{*}}}\|e\|_{L^{2}}.$ Therefore, for some $\ell_{3},\ell_{4},\ell_{5}\!>\!0$ and for all $t\!\in\![t^{*},t^{1}]$ we have
[…]
where $s\mapsto\ell(s)\!=\!\ell_{5}\sqrt{s}\!+\!(\ell_{4}\!+\!\ell_{5})\sum_{i=1}^{\infty}\tfrac{(2\ell_{2})^{i}s^{i}}{(2i-1)!\sqrt{s}}$. Since $\sum_{i=1}^{\infty}\tfrac{(2\ell_{2})^{i}s^{i}}{(2i-1)!\sqrt{s}}=\sqrt{2\ell_{2}}(\exp(\sqrt{2\ell_{2}s})-1)$ and $e(t^{*})=0$ then the obtained estimate yields $e(t)\!=\!0,\forall t\in[t^{*},t^{1}]$. ∎
The table 1 summarizes classes of solutions of the IDE (8) and some conditions to its right-hand side, which are required to ensure existence/uniqueness.
Table 1: Types of solutions of the IDE (8) and conditions to $f,\tilde{f},u$
Solution Class of $f,\tilde{f}$ Class of $u$ Uniqueness
Classical continuous continuous Lipschitz
[Lakshmikantham_etal1995:Book] in $(t,x,u)$ in $(t,x)$
Carathéodory discontin. in $t$; discontin. in $t$; Lipschitz
[OReganMeehan1998:Book] contin. in $(x,u)$ continuous in $x$
Filippov discontinuous discontinuous piece-wise Lipschitz;
in $(t,x,u)$ in $(t,x)$ $\tilde{f}$ is independent of $u$
 and has no common
discontinuity with $f$
Utkin discontin. in $t$; discontinuous piece-wise Lipschitz;
continuous in $(t,x)$ affine-in-control system;
in $(x,u)$ smooth sliding surface
4 Application of discontinuous IDEs for SMC design and analysis
4.1 SMC for a distributed input delay system
Let us consider a control system modeled by the IDE
$\begin{split}\dot{x}(t)=&Ax(t)+B(u(t)+\gamma(t))+p(t)\\ &+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{B}(u(\tau)+\gamma(\tau))d\tau,\end{split}\vskip-2.84526pt$ (45)
where, $t>t_{0}\in\mathbb{R},x(t)\in\mathbb{R}^{n}$, $u(t)\in\mathbb{R}^{m}$, $\Phi:\Delta\to\mathbb{R}^{n\times n}$ be locally essentially bounded on $\Delta$, $\gamma\in L^{\infty}(\mathbb{R};\mathbb{R}^{m})$ is a matched perturbation, $p\in L^{\infty}(\mathbb{R};\mathbb{R}^{n})$ is a possibly mismatched perturbation and $A\in\mathbb{R}^{n\times n},B,\tilde{B}\in\mathbb{R}^{n\times m}$ are known matrices. This IDE can be considered as a linear system with a distributed input delay. SMC design for a linear system with a distributed state delay can be found in [Gouaisbaut_etal2001:TDS]. The aim is to stabilize the output
$y(t)=Cx(t),\quad C\in\mathbb{R}^{m\times n},\quad\det(CB)\neq 0\vskip-2.84526pt$ (46)
to zero in a finite time. For ODEs (i.e. for $\Phi=\mathbf{0}$), the SMC theory (see, [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book]) suggests the feedback
$u(t)=-\rho(t)(CB)^{-1}\frac{y(t)}{\|y(t)\|_{P}},\quad\rho(t)>0,\vskip-2.84526pt$ (47)
where $\|y\|_{P}=\sqrt{y^{\top}Py}$ is the weighted Euclidean norm with a positive definite matrix $P\succ 0$. Let $\|Q\|_{P}=\sup_{y\neq 0}\frac{\|Qy\|_{P}}{\|y\|_{P}}$, $Q\in\mathbb{R}^{m\times m}$ be the matrix norm is induced by the weighted Euclidean norm $\|\cdot\|_{P}$. The condition $\det(CB)\neq 0$ means that the output has the relative degree 1 with respect to the input. This is the standard assumption for the first order sliding mode control design [Utkin1992:Book], [Utkin_etal2020:Book]. If the output has a higher relative degree with respect to the input, then, to solve the considered regulation problem, the high order sliding mode control design algorithms [Levant2003:IJC], [Shtessel_etal2014:Book] must be extended to the IDE (45). This extension goes out of the scope of the paper.
The single-input closed-loop system (45), (47), obviously, satisfies Theorem 14. So, for any initial condition (12), it has a local-in-time Filippov solution. So, the closed-loop system is well-posed provided that solutions are unique. Let us show that, under some restrictions, the feedback (47) is a global finite-time stabilizer of $y$ to zero.
Theorem 20.
If $CA\!=\!{\Lambda C}$ with $\Lambda\!\in\!\mathbb{R}^{m\times m}$ such that
[…]
$M=\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|C\Phi(t,\tau)\tilde{B}(CB)^{-1}\|_{P}d\tau<1,\vskip-2.84526pt$ (49)
$\rho(t)\!\geq\!\tfrac{\left\|C(B\gamma(t)+p(t))\right\|_{P}+\int^{t}_{t^{0}}\!\|C\Phi(t,\tau)\tilde{B}\gamma(\tau)\|_{P}d\tau+\delta}{1-M},\;\delta\!>\!0,\vskip-2.84526pt$ (50)
then, any Filippov solution of the closed-loop system (45), (46), (47) exists and defined for all $t\geq t^{0}$. Moreover, ${\frac{d}{dt}}\|y(t)\|_{P}\leq-\delta$ for $\|y(t)\|_{P}\neq 0,t\geq t^{0}$ and there exists $T\in\left[t^{0},t^{0}+|Cx(t^{0})|/\delta\right]$ such that
$y(t)=Cx(t)=\mathbf{0},\quad\forall t\geq T.$ (51)
Proof.
For any Filippov solution of the closed-loop system (45), (46), (47) it holds
$\begin{split}\tfrac{d\|x(t)\|}{dt}\!\leq&\|Ax(t)\|\!+\!\rho\|B(CB)^{-1}y\|/\|y\|_{P}\!+\!\|B\gamma\|_{L^{\infty}}\!+\!\|p\|_{L^{\infty}}\!\\ &\!\!+\!\left(\!\rho\|\tilde{B}(CB)^{-1}y\|/\|y\|_{P}\!+\!\|\tilde{B}\gamma\|_{L^{\infty}}\!\right)\!\!\int^{t}_{t^{0}}\!\!\|\Phi(t,\tau)\|d\tau.\end{split}\vskip-2.84526pt$
Since $\Phi$ is measurable and locally bounded on $\Delta$ then the function $t\mapsto\int^{t}_{t^{0}}\|\Phi(t,\tau)\|d\tau$ is measurable and locally bounded on $t_{0},+\infty)$. So, using Grönwall–Bellman inequality, we conclude that $\|x(t)\|<+\infty$ for any finite $t\geq t^{0}$.
Assumption $CA={\Lambda C}$ implies that the output $y$ satisfies
$\begin{split}\dot{y}(t)\in&\,{\Lambda y(t)+}CB\gamma(t)-\tfrac{\rho y(t)}{\|y(t)\|_{P}}+Cp(t)\\ &+\!\int^{t}_{t^{0}}C\Phi(t,\tau)\tilde{B}\left(\gamma(\tau)-\tfrac{\rho(CB)^{-1}y(\tau)}{\|y(\tau\|)_{P}}\right)d\tau\end{split}\vskip-2.84526pt$ (52)
for almost all $t\geq t^{0}$. Using (49), (50) and (48), we derive
[…]
as long as $|y(t)|\neq 0$, so $y(t)=0$ for all $t\geq T$. ∎
The above theorem presents a sufficient condition under which $Cx=0$ is a sliding surface of the closed-loop system (45), (46), (47). The SMC theory [Utkin1992:Book] also studies a motion of the system on this surface.
Similarly to ODEs, the sliding motion can be studied using equivalent control method.
Corollary 4.
Under conditions of Theorem 20, the closed-loop system (45), (46), (47) has the unique Utkin solution with
$u_{\rm eq}(t)\!=\!\left\{\begin{array}[]{cl}\!-\rho\frac{(CB)^{-1}y(t)}{\|y(t)\|_{P}},&\!\!t^{0}\!\leq\!t\!<\!T,\\ \!\!\tilde{u}_{\rm eq}(t)-\gamma(t),&\!\!t\!\geq\!T,\end{array}\right.\;\tilde{u}_{\rm eq}\!=\!\sum_{i=0}^{\infty}\!\mathcal{K}^{i}g,\,\vskip-2.84526pt$ (53)
where $T\in[t^{0},t^{0}+T^{\max}]$ is, as before, the reaching time, and the linear bounded operator $\mathcal{K}:L^{\infty}((T,t);\mathbb{R})\to L^{\infty}((T,t);\mathbb{R})$ is given for $s\in[T,t]$ by
$(\mathcal{K}g)(s)\!=\!-(CB)^{-1}\!\int^{s}_{T}\!\!\!C\Phi(t,\tau)\tilde{B}g(\tau)d\tau,\vskip-2.84526pt$ (54)
$g(t)\!=-(CB)^{-1}\left(Cp(t)\!+\tilde{g}(t)\right),\vskip-2.84526pt$ (55)
$\tilde{g}(t)=\int^{T}_{t^{0}}\!\!C\Phi(t,\tau)\tilde{B}\left(\gamma(\tau)-\tfrac{\rho(CB)^{-1}y(\tau)}{\|y(\tau)\|_{P}}\right)\,d\tau.\vskip-2.84526pt$ (56)
For $t\!\geq\!t^{0}\!+\!T$, the dynamics of (45), (46), (47) is governed by
$\begin{split}\!\!\!\!\!\dot{x}(t)\!=&\,Ax(t)\!+\Pi p(t)\!+\!\!\int^{t}_{t^{0}}\!\!\!\!\Pi\Phi(t,\tau)\tilde{B}(u_{\rm eq}(\tau)\!+\!\gamma(\tau))d\tau,\\ 0=&\,Cx(t),\end{split}\vskip-2.84526pt$ (57)
where $\Pi=\left(I_{n}\!-\!B(CB)^{-1}C\right)$ is a projector on $Cx=0$.
Proof.
By Theorem 19, the system has an Utkin solution coinciding with some Filippov solution. Since, by Theorem 20, any Filippov solution belongs a sliding set for all $t\geq T$, then, by Corollary 3, any Utkin solution is unique and
$\begin{split}0=&CB(u_{\rm eq}(t)+\gamma(t))+Cp(t)\\ &+\int^{t}_{t^{0}}C\Phi(t,\tau)\tilde{B}\left(u_{\rm eq}(\tau)+\gamma(\tau)\right)d\tau,\end{split}\quad t\geq T.\vskip-2.84526pt$ (58)
Hence, we derive (57). ∎
Under some additional conditions the dynamics (57) of the system in the sliding mode can be defined by an ODE, e.g., if $\Phi(t,\tau)\tilde{B}\in\mathrm{range}(B)$ then $\Pi\Phi(t,\tau)\tilde{B}=\mathbf{0}.$
Let $\mathrm{spec}(\cdot)$ denotes the spectrum of a matrix. The case of a bounded input delay is treated by the following corollary.
Corollary 5.
Let there exist $d>0$ such that
$\Phi(t,\tau)=\mathbf{0}\quad\text{for all}\quad(t,\tau)\in\Delta:t-\tau\geq d.\vskip-2.84526pt$ (59)
[…]
$\operatorname*{ess\,sup}_{s\geq t}\|p(s)\|\to 0\text{ as }t\to+\infty,\vskip-2.84526pt$ (60)
then for any Utkin solution of the closed-loop system (45), (46), (47) it holds $x(t)\to 0\text{ as }t\to+\infty.$
Proof.
Since $\Phi(t,\tau)=0$ for all $t\geq T+d$ and all $\tau\in[t^{0},T]$ then, for all $t\geq T+d$, it holds $Cx(t)=0$ and
$\begin{split}\!\!\!\!\!\dot{x}(t)\!=&\,Ax(t)\!+\Pi p(t)\!+\!\Pi\!\int^{t}_{t-d}\!\!\!\!\Phi(t,\tau)\tilde{B}\tilde{u}_{\rm eq}(\tau)d\tau,\\ 0=&\,CB\tilde{u}_{\rm eq}(t)+Cp(t)+\int^{t}_{t-d}C\Phi(t,\tau)\tilde{B}\tilde{u}_{\rm eq}(\tau)d\tau,\vskip-2.84526pt\end{split}$ (61)
[…]
$\dot{z}(t)=\tilde{\Lambda}z(t)+C^{\bot}\Pi p(t)\!+\!C^{\bot}\Pi\!\int^{t}_{t-d}\!\!\!\!\Phi(t,\tau)\tilde{B}\tilde{u}_{\rm eq}(\tau)d\tau,\vskip-2.84526pt$
for all $t\geq T+d$. Since, by assumption, the matrix $\tilde{\Lambda}$ is Hurwitz, then $z(t)=C^{\bot}x(t)\to 0$ as $t\to+\infty$. ∎
Example 7.
Let us consider the system (45), (46), (47) with
$A=\left(\begin{smallmatrix}-2&4&2\\ 0&-3&1\\ -1&2&1\end{smallmatrix}\right),\quad B=\tilde{B}=\left(\begin{smallmatrix}0\\ 0\\ 1\end{smallmatrix}\right),\quad C=\left(\begin{smallmatrix}1&0&-2\end{smallmatrix}\right),$
$p=\mathbf{0}$ and $\Phi(t,\tau)=\xi(t-\tau)I_{n}$, where $\xi(t)=1$ for $0\leq t-\tau<1$ and $\xi(t)=0$ for $t-\tau\geq 1$. In this case, the condition (49) is fulfilled with $M=1$ and $|CB|=2$. Therefore, by Theorem 20, the SMC (47) with $\rho>6\|\gamma\|_{L^{\infty}}$ stabilizes the output $y=Cx$ of the system (45) to zero in a finite time. The results of numerical simulations for $\rho=4$, $\gamma(t)=0.5\cos(2t)$ and $t^{0}=0$ are depicted on Figures 1, 2. The simulations have been done by means of explicit Euler method (38) with the sampling period $h=0.001$. The integral is computed using the method of rectangles using the same sampling period.
The control signal depicted on Figure 3 is the typical signal generated by SMC, which has a bounded chattering during the sliding mode (after the time instant $t=T\approx 0.55$). Recall that (see, e.g., [Utkin1992:Book]), in practice, the equivalent control can be estimated by filtering the discontinuous input signal $u(t)$:
$\epsilon\dot{u}_{\epsilon}(t)=-u_{\epsilon}(t)+u(t),\quad t>t^{0},\quad u_{\epsilon}(t^{0})=0,\vskip-2.84526pt$ (62)
where $\epsilon>0$ is a small enough constant parameter. For $t\geq t^{0}+T$, the SMC engineering considers $u_{\epsilon}(t)\approx u_{\rm eq}(t)$. In the view of (58), the function $\delta:[t^{0},+\infty)\to\mathbb{R}$ given by
$\delta(t)=CB(u_{\epsilon}(t)+\gamma(t))+\int^{t}_{t^{0}}C\Phi(t,\tau)\tilde{B}(u(\tau)+\gamma(\tau))d\tau\vskip-2.84526pt$
can be considered as a sliding mode indicator. It must be close to zero for $t\geq T$, provided that Utkin solution is an adequate description of a motion of the considered discontinuous control system. The evolution of the signal $\delta$ computed for $\epsilon=0.01$ is depicted on Figure 4. The numerical computation of $\delta$ has been done using the sampling period $h=0.0001$. The obtained $\delta$ is, indeed, close to zero for $t\geq 0.6$. This indirectly confirms a validity of the Utkin solutions for description of a motion of discontinuous SMC control systems modeled by IDEs. A small delay in detection of SMC is caused by the inertia of the linear filter (62) and by numerical errors.
Figure 1: The states of the closed-loop system (45)-(47)
Figure 2: The output of the closed-loop system (45)-(47)
Figure 3: The control input of the closed-loop system (45)-(47)
Figure 4: The sliding mode indicator $\delta$ of the closed-loop system (45)- (47)
To compare the SMC for IDE with the SMC for ODE , we also present the simulation results of the system (45)-(47) with $\Phi=\mathbf{0}$ (please see Figures 5, 6 and 7). The simulations show a faster convergence of the delay-free ($\Phi=\mathbf{0}$) system to zero.
Figure 5: The states of the closed-loop system (45)-(47) with $\Phi=\mathbf{0}$
Figure 6: The output of the closed-loop system (45)-(47) with $\Phi=\mathbf{0}$
Figure 7: The control input of the closed-loop system (45)-(47) with $\Phi=\mathbf{0}$
4.2 SMC for dynamical systems in Banach spaces
Let us consider a control system modeled by a differential equation in a Banach space [DaleckyKrein1974:Book], [Yosida1980:Book], [Pazy1983:Book]
$\dot{x}(t)\!=\!\!Ax(t)\!+\!B(u(t)+\gamma(t)),\,t\!>\!t_{0},\,x(t_{0})\!=\!x^{0}\!\!\in\!\mathcal{D}(A),\vskip-2.84526pt$ (63)
where $x(t)\in\mathbb{B}$ is the system state at time $t$, $A:\mathcal{D}(A)\subset\mathbb{B}\to\mathbb{B}$ is a linear closed densely defined (possibly) unbounded operator, which generates a strongly continuous semigroup $\Psi(s):\mathbb{B}\to\mathbb{B}$, $s\geq 0$ of linear bounded operators on a real Banach space $\mathbb{B}$, $B:\mathbb{R}^{m}\to\mathbb{B}$ is a linear bounded operator, $u:\mathbb{R}\to\mathbb{R}^{m}$ is the finite-dimensional control input, $\gamma\in L^{\infty}(\mathbb{R};\mathbb{R}^{m})$ is the matched perturbation and $\mathcal{D}(A)$ is the domain of $A$. Let the output of the system be defined as
$y(t)=Cx(t)\in\mathbb{R}^{k},\vskip-2.84526pt$ (64)
where $C:\mathbb{B}\to\mathbb{R}^{k}$ is a linear bounded operator. For any $u\in L^{1}_{\rm loc}([t^{0},+\infty);\mathbb{R}^{m})$, the open-loop systems (63) is well-posed and has the so-called mild solution [Pazy1983:Book, page 106]
$x(t)=\Psi(t-t^{0})x^{0}+\int^{t}_{t^{0}}\Psi(t-\tau)B(u(\tau)+\gamma(\tau))d\tau,\vskip-2.84526pt$ (65)
where the integral is understood in the sense of Bochner [Yosida1980:Book]. The mild solution $x:[t^{0},t^{0}+T)\to\mathbb{B}$ with $0<T\leq+\infty$ is the classical (resp., strong [Pazy1983:Book, page 109]) solution of (63) if $x(t)\!\in\!\mathcal{D}(A)$ satisfies (63) for (almost) all $t\in[t^{0},t^{0}+T)$.
Let a feedback law be defined as follows
$u(t)=\tilde{u}(t,y(t)),\quad\tilde{u}:[t^{0},+\infty)\times\mathbb{R}^{k}\to\mathbb{R}^{m},\vskip-2.84526pt$ (66)
where $\tilde{u}$ is a measurable function. Even if a function $\tilde{u}$ is continuous in $y$, the non-linearity in the right-hand side of the closed-loop system may not be locally Lipschitz continuous, so, the conventional theory of nonlinear abstract differential equations (see, e.g., [Pazy1983:Book], [EngelNagel2000:Book]) is not applicable for its analysis. In the case of SMC design [Orlov_elal2004:IJC], the closed-loop system (63), (64), (66) is ill-posed. Let us regularize the system (63), (64) following the concept of Utkin solutions [Orlov2008:Book], [PolyakovFridman2014:JFI].
Definition 9.
A continuous function $x:[t_{0},t^{0}+T)\to\mathbb{B}$ with $x(t^{0})\!=\!x^{0}$ and $0<T\leq+\infty$ is said to be an Utkin solution of the first (resp., second) kind for the system (63), (64), (66) if $x$ satisfies (65) with $u=u_{\rm eq}\in L^{1}((t_{0},t^{0}+T);\mathbb{R}^{m})$ such that $u_{\rm eq}(t)\in U(t,Cx(t))$ for almost all $t\!\in\!(t_{0},t^{0}+T)$, where $U\!=\!K[\tilde{u}$] (resp., $U=(K[\tilde{u}_{1}],...,K[\tilde{u}_{m}])^{\top}$ for $\tilde{u}=(\tilde{u}_{1},...,\tilde{u}_{m})^{\top}$).
In other words, an Utkin solution $x$ of (63), (64), (66) is a mild solution with a properly selected $u(t)\!\in\!U(t,Cx(t))$. If $B$ maps $\mathbb{R}^{m}$ into the domain of $A$ (i.e., $\mathrm{range}(B)\!\subset\!\mathcal{D}(A)$), then, by [Pazy1983:Book, Theorem 2.9, Chapter 4], any mild solution is a strong solution. By [Pazy1983:Book, Theorem 2.4, Chapter 1], the function
$s\in[0,+\infty)\mapsto\Phi(s):=C\Psi(s)AB\in\mathbb{R}^{k\times m}\vskip-2.84526pt$
is continuous. Moreover, taking into account $\frac{d\Psi(s)B}{ds}=\Psi(s)AB$ and $\frac{d\Psi(s)x^{0}}{ds}=\Psi(s)Ax^{0},\forall s>0,\forall x^{0}\in\mathcal{D}(A)$ (see, [Pazy1983:Book, Theorem 2.4, Chapter 1]), we derive
$\begin{split}\tfrac{d}{dt}y(t)=\tfrac{d}{dt}Cx(t)=&\frac{d}{dt}C\Psi(t-t)x^{0}\\ &+\frac{d}{dt}\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)B}_{\in\mathbb{R}^{k\times m}}(u(\tau)+\gamma(\tau))d\tau\\ =&C\Psi(t-t)Ax^{0}+CB(u(t)+\gamma(t))\\ &+\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)AB}_{\Phi(t-\tau)}(u(\tau)+\gamma(\tau))d\tau\end{split}$
almost everywhere. Therefore, the input-output dynamics of the system (63), (64) is given by the IDE
$\dot{y}(t)=p(t-t^{0})+CB(u(t)+\gamma(t))+\int^{t}_{\sigma}\!\!\Phi(t-\tau)(u(\tau)+\gamma(\tau))d\tau,\vskip-2.84526pt$ (67)
where $t>t^{0}$, $CB\in\mathbb{R}^{k\times m}$ and
$y(t^{0})\!=\!Cx^{0},\quad p(t-t^{0})\!=\!C\Psi(t-t^{0})Ax^{0}.\vskip-2.84526pt$ (68)
Notice that the function $s\mapsto p(s)$ is continuous if $x_{0}\in\mathcal{D}(A)$.
Let $y_{\rm eq}$ be an Utkin solution of the IDE (67), (68) with the feedback law (66), and let $u_{\rm eq}$ be the corresponding equivalent control. By Theorem 19, the pair $(y,u_{\rm eq})$ always exists if $\tilde{u}$ is a locally essentially bounded measurable function. If $x_{\rm eq}$ is a mild solution of (63) given by (65) with $u\!=\!u^{\rm eq}$, then
$\begin{split}\tfrac{d}{dt}Cx_{\rm eq}(t)=&C\Psi(t-t)Ax^{0}+CB(u_{\rm eq}(t)+\gamma(t))\\ &+\int^{t}_{\sigma}\!\!\Phi(t-\tau)(u_{\rm eq}(\tau)+\gamma(\tau))d\tau,\end{split}\vskip-2.84526pt$
i.e., $y_{\rm eq}(t)=Cx_{\rm eq}(t)$ and $u_{\rm eq}(t)\in U(t,Cx(t))$ almost everywhere. By Definition 9, this means, that the function $x_{\rm eq}$, constructed by solving the IDE (67), is an Utkin solution of the nonlinear (possibly discontinuous) differential equation (63), (66) in a Banach space. As a trivial corollary of Theorem 19 (resp., Theorem 12), we derive the following result.
Corollary 6.
If, $\mathrm{range}(B)\subset\mathcal{D}(A)$ and $\tilde{u}$ satisfies Filippov Condition, then, for any $x_{0}\in\mathcal{D}(A)$, the system (63), (64), (66) has an Utkin solution solution $x:[t^{0},t^{0}+T)\to\mathcal{D}(A)$, which is defined, at least, locally in time and satisfies the differential equation (63) with $u=u_{\rm eq}$ almost everywhere, where $0<T\leq+\infty$. If, additionally, $\Phi\in L^{\infty}((0,+\infty);\mathbb{R}^{k\times m})$ and $\operatorname*{ess\,sup}_{t\in\mathbb{R},y\in\mathbb{R}^{k}}\|\tilde{u}(t,y)\|<+\infty$, then $T=+\infty$.
The uniqueness of Utkin solution of the IDE (67), (66) implies the uniqueness of Utkin solution of the system (63), (64), (66).
Corollary 7.
Let $\mathrm{range}(B)\subset\mathcal{D}(A^{2})$ and $\tilde{u}$ satisfy Filippov (resp., Carathéodory) Condition. If Utkin solution $x(t,t^{0},x^{0})$ of (63), (64), (66) is unique on $\mathcal{I}$ then it depends continuously on initial data: $x(t,t^{1},x^{1})\to x(t,t^{0},x^{0})$ as $|t^{1}-t^{0}|+\|x^{1}-x^{0}\|_{\mathbb{B}}+\|Ax^{1}-Ax^{0}\|_{\mathbb{B}}\to 0$ unformly on compacts from $\mathcal{I}$.
Proof.
Let $x(t,t^{0},x^{0})$ be an Utkin solution with the corresponding equivalent control $u_{\rm eq}$. Let us denote $\xi(t,t^{0},x^{0})=\int^{t}_{t^{0}}u_{\rm eq}(\tau,t^{0},x^{0})d\tau$. Since $\mathrm{range}(B)\subset\mathcal{D}(A^{2})$ then $\frac{d\Psi(s)AB}{ds}=\Psi(s)A^{2}B,\forall s>0$ and
$\begin{split}-\xi(t,t^{0},x^{0})=&\,\Phi(t-t^{0})\xi(t^{0},t^{0},x^{0})-\Phi(0)\xi(t,t^{0},x^{0})\\ =&\int^{t-t^{0}}_{0}\frac{d}{d\sigma}\left(\Phi(\sigma)\xi(t-\sigma,t^{0},x^{0})\right)d\sigma\\ =&\int^{t-t^{0}}_{0}C\Psi(\sigma)A^{2}B\xi(t-\sigma,t^{0},x^{0})d\sigma\\ &-\int^{t-t^{0}}_{0}\Phi(\sigma)\dot{\xi}(t-\sigma,t^{0},x^{0})d\sigma\end{split}\vskip-2.84526pt$
Hence, $\int^{t}_{t^{0}}\Phi(t-\tau)\dot{\xi}(\tau,t^{0},x^{0})d\tau=\xi(t,t^{0},x^{0})+\int^{t}_{t^{0}}C\Psi(t-\tau)A^{2}B\xi(\tau,t^{0},x^{0})d\tau$ and the pair $y(t,t^{0},x^{0})=Cx(t,t^{0},x^{0})$, $\xi(t,t^{0},x^{0})$ is an Utkin solution of the IDE
$\begin{split}\dot{y}(t,t^{0},x^{0})\in&\,p(t,t^{0},x^{0})+CB(U(t,y(t,t^{0},x^{0}))+\gamma(t))\\ &+\int^{t}_{t^{0}}\Phi(t-\tau)\gamma(\tau)d\tau+\xi(t,t^{0},x^{0})\\ &+\int^{t}_{t^{0}}C\Psi(t-\tau)A^{2}B\xi(\tau,t^{0},x^{0})d\tau,\\ \dot{\xi}(t,t^{0},x^{0})\in&\,U(t,y(t,t^{0},x^{0})),\quad\xi(t^{0})=0,\quad y(t^{0})=Cx^{0},\end{split}\vskip-2.84526pt$
where $p(t,t^{0},x^{0})=C\Psi(t-t^{0})Ax^{0}$. Since $p(t,t^{0},x^{0})=C\Psi(t-t^{0})Ax^{0}$ depends continuously on $Ax^{0}$, then, using Theorem 15, we conclude that $\xi(t,t^{0},x^{0})$ depends continuously on $(t^{0},x^{0},Ax^{0})$. Since $\mathrm{range}(B)\!\subset\!\mathcal{D}(A^{2})\!\subset\!\mathcal{D}(A)$ then
$\begin{split}x(t,t^{0}\!,x^{0})\!=\,&\Psi(t\!-\!t^{0})x^{0}\!+\!\!\int^{t}_{t^{0}}\!\!\Psi(t\!-\!\tau)B(\dot{\xi}(\tau,t^{0}\!,x^{0})\!+\!\gamma(\tau))d\tau\\ =&\Psi(t-t^{0})x^{0}\!+\!B\xi(t,t^{0}\!,x^{0})\!+\!\!\int^{t}_{t^{0}}\!\!\!\Psi(t\!-\!\tau)B\gamma(\tau)d\tau\\ &+\int^{t}_{t^{0}}\Psi(t-\tau)AB\xi(\tau,t^{0},x^{0})d\tau,\end{split}\vskip-2.84526pt$
where the identity $\frac{d\Psi(s)B}{ds}=\Psi(s)AB,\forall s>0$ and the integration by parts has been utilized. Continuous dependence of $\xi(t,t^{0},x^{0})$ on $(t^{0},x^{0},Ax^{0})$ implies continuous dependence of $x(t,t^{0},x^{0})$ on $(t^{0},x^{0},Ax^{0})$. ∎
The discontinuous IDE (67) is utilized above for well-posedness analysis of the differential equation in the Banach space. This IDE, can also be used for SMC design by means of Theorem 20. However, since, by Corollary 6, any Utkin solution satisfies (63) with $u(t)=u_{\rm eq}(t)\in U(t,Cx(t))$ almost everywhere, then a control design as well as stability/completeness/robustness analysis can also be based on the original differential equation (63) in the Banach space.
Example 8.
1) Regularization of PDE using IDE. The example (1)–(6) admits the representation (63) with $A\!=\!\nu\frac{\partial^{2}}{\partial z^{2}}$, $D(A)\!=\!H^{2}((0,1);\mathbb{R})\cap H^{1}_{0}((0,1);\mathbb{R})$, $(Bx)(z)\!=\!\beta(z)x(z)$, where $H^{2}$, $H^{1}_{0}$ are Sobolev spaces. The closed densely defined operator $A:\mathcal{D}(A)\subset L^{2}((0,1);\mathbb{R})\to L^{2}((0,1);\mathbb{R})$ is a generator of the strongly continuous semigroup of linear bounded operators on $L^{2}((0,1);\mathbb{R})$ (see, e.g., [Pazy1983:Book, page 212]). The output (64) is given by $y(t)=Cx(t)=\int^{1}_{0}\xi(z)x(t,z)dz$. The control law (66) is defined by the formula (6). Since $\beta\in\mathcal{D}(A)$, then, by Corollary 6, the system (1)–(6) has an Utkin solution $x(t)\!\in\!\mathcal{D}(A)$, $\forall x^{0}\in\mathcal{D}(A)$, $\forall q\!\in\!L^{\infty}((t^{0},+\infty);\mathbb{R})$ in (6).
2) SMC design using PDE-based approach. The closed-loop heat control system (1)-(6) is forward complete, i.e., all Utkin solutions exist globally in time. Indeed, considering $V(t)\!=\!\int^{1}_{0}\!x^{2}(t,z)dz$ and using integration by parts, we derive
$\dot{V}(t)=2\int^{1}_{0}\nu x(t,z)\tfrac{\partial^{2}x(t,z)}{\partial z^{2}}+x(t,z)\beta(z)u_{\rm eq}(t)dz\vskip-5.69054pt$
$2\int^{1}_{0}\nu\left(\tfrac{\partial x(t,z)}{\partial z}\right)^{2}+x(t,z)\beta(z)u_{\rm eq}(t)dz,\vskip-2.84526pt$
where $u_{\rm eq}(t)\in q(t)\,\overline{\mathrm{sign}}(Cx(t))$. Since $x(t,\cdot)\in\mathcal{D}(A)$ almost everywhere then the function $z\mapsto x(t,z)$ is continuously differentiable in $z$ for almost all $t$. Using Wirtinger and Cauchy-Schwartz inequalities, we come up with the estimate
$\dot{V}(t)\leq-\nu 2\pi^{2}V(t)+2\|q\|_{L^{\infty}}\|\beta\|_{L^{2}}\sqrt{V(t)},$
that is fulfilled almost everywhere. Therefore, any Utkin solution of the closed-loop system is bounded and
$\limsup_{t\to+\infty}\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{\|q\|_{L^{\infty}}\|\beta\|_{L^{2}}}{\nu\pi^{2}}.\vskip-2.84526pt$ (69)
Since $CB=\int^{1}_{0}\xi(z)\beta(z)dz$ and $Cx(t)=\int^{1}_{0}\xi(z)x(t,z)dz$ then, due to $\xi(0)=\xi(1)=x(t,0)=x(t,1)=0$ and $\xi\in C^{2}([0,1];\mathbb{R})$, the integration by parts yields
$\begin{split}\tfrac{dCx(t)}{dt}=&\int^{1}_{0}\nu\xi(z)\tfrac{\partial^{2}x(t,z)}{\partial z^{2}}dz+CB(u_{\rm eq}(t)+\gamma(t))\\ =&\int^{1}_{0}\nu\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz+CB(u_{\rm eq}(t)+\gamma(t))\\ =&-\nu\lambda Cx(t)+\nu\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)x(t,z)dz\\ &+CB(u_{\rm eq}(t)+\gamma(t)),\quad\quad\forall\lambda\geq 0.\end{split}\vskip-2.84526pt$ (70)
Taking into account that $u_{\rm eq}(t)\in-q(t)\overline{\mathrm{sign}}(Cx(t))$, we conclude that, for $Cx(t)\!\neq\!0$, it holds $u_{\rm eq}(t)\!=\!-q(t)\mathrm{sign}(Cx(t))$,
$\begin{split}\tfrac{d|Cx(t)|}{dt}\leq&\,-\nu\lambda|Cx(t)|+CBq(t)\\ &+\nu\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\!\|x(t,\cdot)\|_{L^{2}}+\|CB\gamma\|_{L^{\infty}}\end{split}\vskip-2.84526pt$
almost everywhere. If $CB=\int^{1}_{0}\xi(z)\beta(z)dz\neq 0$ then, taking $\delta>0$ and $q(t)=-\frac{\rho}{CB}$ with $\rho>0$, we have due to (69)):
$\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{\rho\|\beta\|_{L^{2}}}{\pi^{2}|CB|}+\delta\vskip-2.84526pt$
for almost all $t\geq t^{0}+\tilde{T}(x_{0})$ provided that $\tilde{T}(x_{0})\geq 0$ is large enough. If $\exists\lambda\geq 0$ such that
$\sqrt{\!\int^{1}_{0}\!\!\!\left(\!\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}\!+\!\lambda\xi(z)\!\right)^{\!2}\!\!\!dz\!\int^{1}_{0}\!\!\!\!\beta^{2}(z)dz}\!<\!\pi^{2}\!\left|\int^{1}_{0}\!\!\!\!\xi(z)\beta(z)dz\right|\vskip-2.84526pt$ (71)
then, for
$\rho>\tfrac{\pi^{2}|CB|\left(\bar{\gamma}|CB|+\nu\delta\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\right)}{\pi^{2}\left|CB\right|-\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}},\quad\bar{\gamma}\geq\|\gamma\|_{L^{\infty}},\vskip-2.84526pt$
we derive the following estimate
$\tfrac{d|Cx(t)|}{dt}\!\leq\!\nu\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}\!+\!\lambda\xi\right\|_{L^{2}}\!\left(\tfrac{\rho\|\beta\|_{L^{2}}}{\nu\pi^{2}|CB|}\!+\!\delta\right)\!+|CB|\cdot\|\gamma\|_{\infty}-\rho\!<\!0\;$
for almost all $t\geq t^{0}+\tilde{T}(x^{0})$ such that $Cx(t)\neq 0$. Therefore, there exists a finite number $T\geq 0$ such that $Cx(t)\to 0$ as $t\to T$. Moreover, since $t\mapsto Cx(t)$ is an absolutely continuous function defined for all $t\geq t^{0}$ then the implication $|Cx(t)|\neq 0\Rightarrow\frac{d|Cx(t)|}{dt}<0$ implies that $Cx(t)=0$ for all $t\geq T$. Therefore, the system (1), (2), (4), (5), (6) has a sliding mode and the identity $\frac{dCx(t)}{dt}=0$ implies
$u_{\rm eq}(t)=-\gamma(t)-\tfrac{\nu\int^{1}_{0}\frac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz}{CB}\vskip-2.84526pt$
for almost all $t\geq T$. The dynamics of the original system in the sliding mode is defined by the equation
$\tfrac{\partial x(t,z)}{\partial t}\!=\!\nu\tfrac{\partial^{2}x(t,z)}{\partial z^{2}}\!-\!\tfrac{\nu\int^{1}_{0}\frac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz}{\int^{1}_{0}\xi(z)\beta(z)dz}\beta(z),\vskip-2.84526pt$ (72)
with the homogeneous boundary conditions (2) and
$\int^{1}_{0}\xi(z)x(t,z)=0\vskip-2.84526pt$ (73)
for almost all $t\geq T$.
3) SMC design using IDE-based approach. The same result can be obtained considering the input-output dynamics of the heat control system. Using the modal decomposition
$x(t,z)=\sum_{i=1}^{\infty}x_{i}(t)\phi_{i}(z),\quad x_{i}(t)\in\mathbb{R},\vskip-2.84526pt$ (74)
where $\phi_{i}(z)=\sqrt{2}\sin(\sqrt{\lambda_{i}}z),\lambda_{i}=\pi^{2}i^{2}$, the system (1) - (3) can be rewritten (see, e.g. [Balas1988:JMAA], [PrieurTrelat2019:TAC], [KatzFridman2021:EJC]) in the form
$\dot{x}_{i}(t)=-\nu\lambda_{i}x_{i}(t)+b_{i}(u(t)+\gamma(t)),\quad t>t_{0},\vskip-2.84526pt$ (75)
where $x_{i}(0)=x^{0}_{i}=\int^{1}_{0}\!\phi_{i}(z)x^{0}(z)dz$, $b_{i}=\int^{1}_{0}\!\phi_{i}(z)\beta(z)dz$. By Cauchy formula, we have
$x_{i}(t)\!=\!e^{-\nu\lambda_{i}(t-t_{0})}x^{0}_{i}+\int^{t}_{t_{0}}\!\!e^{-\nu\lambda_{i}(t-\tau)}b_{i}(u(\tau)+\gamma(\tau))d\tau.\vskip-2.84526pt$
Therefore, the input-output dynamics (70) is given by
$\begin{split}\dot{y}(t)=&-\nu\lambda y(t)+p(t,t^{0})+CB(u(t)+\gamma(t))\\ &+\int^{t}_{t^{0}}\Phi(t,\tau)(u(\tau)+\gamma(\tau))d\tau,\end{split}\vskip-5.69054pt$ (76)
$p(t,t^{0})=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-t^{0})}x_{i}^{0}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz,\vskip-5.69054pt$
$\Phi(t,\tau)=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-\tau)}b_{i}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz.\vskip-2.84526pt$
Since $|\Phi(t,\tau)|\leq\nu e^{-\nu\pi^{2}(t-\tau)}\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}$ then the inequality (71) yields $M=\frac{\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|\Phi(t,\tau)|d\tau}{|CB|}\!<\!1$. Let $q(t)=-\frac{|CB|\cdot\|\gamma\|_{L_{\infty}}+|p(t,t^{0})|+M\|\gamma\|_{L^{\infty}}+\delta}{(1-M)CB}$ with $\delta>0$. By Theorem 20 and Corollary 4, the system (76), (6) has the unique Utkin solution and $y(t)=0$ for all $t\geq T$, where $T\geq t^{0}$ is some instant of time. This means that the Utkin solution of the system (1)–(6) satisfies (73) for all $t\geq T$.
4) Stability of the heat system in the sliding mode. Let us show that the system (72), (73), (3) is globally exponentially stable. Considering again the functional $t\to V(t)=\int^{1}_{0}x^{2}(t,z)dz$ , we derive
$\dot{V}(t)\leq-2\nu\pi^{2}V(t)-2\nu\tfrac{\int^{1}_{0}\frac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz\int^{1}_{0}\beta(z)x(t,z)dz}{\int^{1}_{0}\xi(z)\beta(z)dz}.\vskip-2.84526pt$
The identity (73) gives
$\begin{split}\left|\int^{1}_{0}\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz\right|=&\left|\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)x(t,z)dz\right|\\ \leq&\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\sqrt{V(t)}.\vskip-2.84526pt\end{split}$
Using the inequality $\left|\int^{1}_{0}\beta(z)x(t,z)dz\right|\leq\|\beta\|_{L^{2}}\sqrt{V(t)}$ and the condition (71) we come up with the estimate
$\dot{V}(t)\leq 2V(t)\left(\tfrac{\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\|\beta\|_{L^{2}}}{|CB|}-\pi^{2}\right)<0\text{ for }V(t)\neq 0,\vskip-2.84526ptx$
i.e., the system (72), (73), (3) is globally exponentially stable.
The condition (71) is fulfilled, for example, if
$\beta(z)=\left\{\begin{array}[]{ll}\exp\left(\frac{1}{(0.3-x)(0.6-x)}\right)&\text{ if }x\in(0.3,0.6),\\ 0&\text{ otherwise},\end{array}\right$.
$\xi(z)=\sin(\pi z)+z(1-z)\quad\text{ and }\quad\lambda=\pi^{2}.$
In this case, we have $CB\approx 0.33$, $\|\beta\|_{L^{2}}\approx 0.55$ and $\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\approx 0.8$ for $\lambda=\pi^{2}$. The figures 8, 9, 10 show simulation results for the heat system (1)-(5) with the sliding mode control (6), $q(t)=-\frac{\rho}{CB},\rho=1$ and the matched perturbation $\gamma(t)=0.5$. The initial state $x^{0}\in\mathcal{D}(A)$ is defined as follows $x_{0}(z)=10z(1-z),z\in[0,1]$. The method of modal decomposition [Balas1988:JMAA], [PrieurTrelat2019:TAC], [KatzFridman2021:EJC], [Ayamou_etal2025:IMA] (with 60 modes) is utilized for numerical simulation of the closed-loop SMC system. The time sampling is $0.001$.
Figure 8: The states of the heat system (1)-(5) with the SMC (6)
Figure 9: The output of the heat system (1)-(5) with the SMC (6)
Figure 10: The control input of the heat system (1)-(5) with the SMC (6)
The finite-dimensional IDE-based procedure of regularization of infinite-dimensional systems with sliding mode control can be extended to other systems using, for example, the methodology suggested in [Polyakov2025:CDC]. Further study of the SMC design for such systems goes out of the scope of this paper.
5 Conclusions
In this paper, Filippov method is extended to systems modeled by integro-differential equations. This method allows an ill-posed discontinuous system to be regularized by means of its transformation to a well-posed differential inclusion. Such ill-posed discontinuous systems appear in sliding mode control theory, which uses discontinuous feedback laws for regulation of systems and rejection of matched perturbations. The advantage of the Filippov method consists in a minimal (from set-theoretic point if view) extension of an ill-posed equation to a well-posed equation. This implies a better matching of the dynamics of the control system by the regularized model. In the paper, the equivalent control method is developed for affine-in-control systems modeled by integro-differential equations. The well-posedness of regularized systems is studied. Filippov method and equivalent control approach are the main tools for analysis of control systems with sliding modes.
To illustrate theoretical results, sliding mode control is designed for a class of systems with distributed input delay. The dynamics of the closed-loop system in the sliding mode is obtained in the form of an algebro-integro-differential equation. The equivalent control is defined as a solution of the Volterra integral equation of the second kind.
A SMC for a plant modeled by infinite-dimensional differential equation with finite-dimensional inputs and outputs can also be designed using the developed theory of discontinuous IDEs. Such a discontinuous infinite-dimensional control system can be regularized based on a finite-dimensional IDE representing its input-output dynamics. This scheme is demonstrated on a heat control system modeled by PDE.
The obtained theoretical results provide all necessary tools for extension of other methods of SMC theory (e.g., high order SMC [Levant2003:IJC] or implicit SMC discretization [AcaryBrogliato2010:SCL]) to systems modeled by integro-differential equations and by differential equations in Banach spaces (e.g., PDEs).

### v3

@@
 The paper deals with analysis and design of sliding mode control systems modeled by finite-dimensional integro-differential equations.
-Filippov method and equivalent control approach are extended to a class of nonlinear discontinuous integro-differential equations.
-Sliding mode control algorithms are designed for distributed input delay systems and for systems modeled by differential equations in Banach spaces.
-The obtained results are illustrated by numerical examples.
+Filippov method and equivalent control approach are extended to a class of nonlinear discontinuous integro-differential equations and to a class of control systems modeled by infinite-dimensional differential equations in Banach spaces.
+Sliding mode control algorithms are designed for distributed input delay systems and for a heat control system.
 1 Introduction
@@
 Sliding Mode Control (SMC) is a robust control methodology, which uses discontinuous feedback laws for rejection of matched disturbances.
-The design and analysis of SMC is well-developed for finite-dimensional systems [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book].
+The design and analysis of SMC is well-developed for finite-dimensional systems [49], [14], [46].
 Mathematical models of finite-dimensional SMC systems are ill-posed discontinuous ODEs (Ordinary Differential Equations).
-Filippov method [Filippov1988:Book], [Cortes2008:CSM] is aimed at regularization of discontinuous ODEs (Ordinary Differential Equations).
+Filippov method [17], [11] is aimed at regularization of discontinuous ODEs.
 The method consists in a minimal (in a set-theoretic sense) extension of an ill-posed discontinuous ODE to a well-posed ODI (Ordinary Differential Inclusion).
-The Filippov regularization has a rather simple mathematical formulation [Filippov1988:Book] based on a notion of the set of Lebesgue measure zero.
-Filippov method is the main tool for regularization of a SMC system [Filippov1988:Book], [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book], [PolyakovFridman2014:JFI], [Utkin_etal2020:Book] while an analysis of the closed-loop dynamics of the SMC system is usually based on the so-called equivalent control method [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book].
-Formally, the finite-dimensional Filippov method is inapplicable to discontinuous infinite-dimensional systems in abstract spaces [DaleckyKrein1974:Book] (e.g., due to the absence of a notion of the set of Lebesgue measure zero in an abstract space).
-Traditionally (see, e.g., [OrlovUtkin1985:DAN], [Orlov_elal2004:IJC], [Orlov2008:Book] and references therein), SMC design for infinite-dimensional systems is based on an equivalent control method.
-Well-posedness analysis of infinite-dimensional SMC systems requires rather complicated constructions [Levaggi2002:DIE], [Pisano_etal2011:SIAM_JCO], [Levaggi2013:DCDS], [Orlov2020:Book].
-Infinite-dimensional differential inclusion obtained from discontinuous infinite-dimensional differential equation in an abstract space (like in [Mohet_etal2023:SCL]) is a non-trivial object for research [Tolstonogov2000:Book].
-Some generalized solutions to discontinuous functional differential equations of retarded type are introduced on [KolmanovskiiMyshkis1992:Book, page 58].
+The Filippov regularization has a rather simple mathematical formulation [17] based on the notion of a set of Lebesgue measure zero.
+Filippov method is the main tool for regularization of a SMC system [17], [49], [14], [46], [43], [50] while an analysis of the closed-loop dynamics of the SMC system is usually based on the so-called equivalent control method [49], [14], [46].
+Formally, the finite-dimensional Filippov method is inapplicable to discontinuous infinite-dimensional systems in abstract spaces [12] (e.g., due to the absence of a notion of the set of Lebesgue measure zero in an abstract space).
+Traditionally (see, e.g., [36], [35], [37] and references therein), SMC design for infinite-dimensional systems is based on an equivalent control method.
+Well-posedness analysis of infinite-dimensional SMC systems requires rather complicated constructions [27], [41], [28], [38].
+Infinite-dimensional differential inclusion obtained from discontinuous infinite-dimensional differential equation in an abstract space (like in [31]) is a non-trivial object for research [47].
+Some generalized solutions to discontinuous functional differential equations of retarded type are introduced on [24, page 58].
 The corresponding regularization procedure consists in a construction of a convex hull of limits of all possible convergent sequences in a function space.
 Applicability of such non-constructive procedure to analysis and design of SMC systems seems questionable too.
-Integro-Differential Equations (IDEs) [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book] model various finite and infinite dimensional systems, such as epidemics [KermackMcKendrick1991:BMB], dispersive waves [Debnath2005:Book], systems with delays [Hale1977:Book], [Fridman2014:Book], [Orlov2020:Book], and others.
+Integro-Differential Equations (IDEs) [26], [34] model various finite and infinite dimensional systems, such as epidemics [23], dispersive waves [13], systems with delays [21], [18], [38], and others.
 Input-output dynamics of some systems modeled by PDEs (Partial Differential Equations) can also be described by IDEs (see Section 1.2).
-Being a model of (possibly) infinite-dimensional system, an IDE is usually defined by functions in finite-dimensional spaces (see, e.g., [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book]).
+Being a model of (possibly) infinite-dimensional system, an IDE is usually defined by functions in finite-dimensional spaces (see, e.g., [26], [34]).
 To the best of author’s knowledge, finite-dimensional Filippov method and SMC theory has never been developed for IDEs.
 1.2 Motivation example: SMC design for a heat system
-Let us consider a control system modeled by the PDE
+Let us consider a control system modeled by the heat PDE
 $\tfrac{\partial x}{\partial t}\!=\!\nu\tfrac{\partial^{2}x}{\partial z^{2}}+\!\beta(z)(u(t)\!+\!\gamma(t)),\quad t\geq t_{0}\vskip-2.84526pt$ (1)
@@
 $x^{0}\in\tilde{\mathcal{D}}=\{x\in C^{2}([0,1];\mathbb{R}):x(0)=x(1)=0\}.$ (4)
-A control system similar to (1) also appears in the case of the boundary control design using the Dirichlet lifting [PrieurTrelat2019:TAC], [KatzFridman2021:EJC].
+A control system similar to (1) also appears in the case of the boundary control design using the Dirichlet lifting [45], [22].
 Let a measured output of the system be modeled as
@@
 $u(t)\!=\!q(t)\mathrm{sign}(y(t)),\quad q\in L^{\infty}([t_{0},+\infty);\mathbb{R}).\vskip-2.84526pt$ (6)
-This is a SMC provided that $q$ is properly selected [Orlov_elal2004:IJC], [Orlov2020:Book].
+This is a SMC provided that $q$ is properly selected [35], [38].
 The infinite-dimensional closed-loop control system (1) – (6) is ill-posed in the SMC case.
-A regularization of this discontinuous PDEs with SMC is a nontrivial problem (see, e.g., [Orlov2020:Book, pages 24-28], [Mohet_etal2023:SCL]).
+A regularization of this discontinuous PDEs with SMC is a nontrivial problem (see, e.g., [38, pages 24-28], [31]).
 However, the considered infinite-dimensional control system has the scalar input $u$ and the scalar output $y$.
@@
 $\dot{y}(t)\!=\!p(t-t^{0})+b(u(t)+\gamma(t))+\!\!\int^{t}_{t_{0}}\!\!\!\!\Phi(t-\tau)(u(\tau)+\gamma(\tau))d\tau,\!\!\vskip-5.69054pt$ (7)
-where $y(0)=\int^{1}_{0}\xi(z)x^{0}(z)dz$, $b=\int^{1}_{0}\xi(z)\beta(z)dz$, and $p,\Phi$ are some decreasing functions (see Section 4).
+where $y(0)=\int^{1}_{0}\xi(z)x^{0}(z)dz$, $b=\int^{1}_{0}\xi(z)\beta(z)dz$, and $p,\Phi$ are some continuous functions (see Example 8 in Section 4).
 If the control law $u$ is defined by the formula (6), then the input-output dynamics of the system (1) - (6) should satisfy the discontinuous IDE (7), (6).
-Due to discontinuity of $u$ on $y$, this IDE is ill-posed [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book].
-However, the IDE is defined in terms of finite-dimensional functions, which admit Filippov regularization [Filippov1988:Book].
+Due to discontinuity of $u$ on $y$, this IDE is ill-posed [26], [34].
+However, the IDE is defined in terms of finite-dimensional functions, which admit Filippov regularization [17].
 This paper extends the Filippov method to a class of finite-dimensional IDEs, which covers the IDE (7), (6) as a particular case.
-A regularization of the IDE (7), (6) allows the original discontinuous PDE (1) – (6) to be easily regularized (see, e.g., [Polyakov2025:CDC]).
-Such an IDE-based approach to SMC design and analysis is applicable to a rather large class of systems modeled by differential equations in Banach spaces
+A regularization of the IDE (7), (6) allows the original discontinuous PDE (1) – (6) to be easily regularized (see, e.g., [44]).
+Such an IDE-based approach to SMC design and analysis is applicable to a rather large class of systems modeled by differential equations in Banach spaces.
 1.3 Contributions of the paper
@@
   • The equivalent control method is developed for affine-in-control IDEs (i.e., for $f=a+bu$ and $\tilde{f}=\tilde{a}+\tilde{b}u$).
-  • A SMC design for some control systems with distributed input delay is presented.
-  • An IDE-based method of SMC design and analysis for systems modeled by differential equations in Banach spaces is developed and illustrated on the system (1)-(6).
+  • A SMC design for control systems modeled by linear IDEs is presented.
+  • An IDE-based SMC design and analysis for some control systems modeled by differential equations in Banach spaces is developed and illustrated on the system (1)-(6).
 1.4 Organization of the paper
@@
 The surveyed results are utilized in the next section, where an extension of Filippov method to IDEs is proposed.
-After that, the developed theoretical results are applied for SMC design and analysis of finite-dimensional systems with distributed input delay and of infinite-dimensional systems modeled by differential equations in Banach spaces.
+After that, the obtained theoretical results are applied to SMC design and analysis for finite-dimensional systems with distributed input delay and for infinite-dimensional systems modeled by differential equations in Banach spaces.
 Finally, concluding remarks are given.
@@
 2.1 Some classes of functions
-For $1\!\leq\!p\!\leq\!\infty$, $L^{p},$ denotes the well-known Lebesgue space. A function $(t,x)\in\mathbb{R}\times\mathbb{R}^{n}\mapsto f(t,x)\in\mathbb{R}^{n}$ is said to satisfy
-  • Carathéodory Condition if it is measurable in $t$ for any $x$, continuous in $x$ for almost any $t$ and for any $\alpha,r>0$ there exists $m\in L^{1}((-\alpha,\alpha);\mathbb{R})$ such that $\left\|f(t,x)\right\|\leq m(t)$ for almost all $t\!\in\!(-\alpha,\alpha)$, for all $x\!\in\!\mathcal{B}(r)$, where $\mathcal{B}(r)$ is an open ball of the radius $r>0$.
+For $1\!\leq\!p\!\leq\!\infty$, $L^{p}$ denotes the well-known Lebesgue space. A function $(t,x)\in\mathbb{R}\times\mathbb{R}^{n}\mapsto f(t,x)\in\mathbb{R}^{n}$ is said to satisfy
+  • Carathéodory Condition if $f$ is measurable in $t$ for all $x$, continuous in $x$ for almost all $t$ and for any $\alpha,r>0$ there exists $m\in L^{1}((-\alpha,\alpha);\mathbb{R})$ such that $\left\|f(t,x)\right\|\leq m(t)$ for almost all $t\!\in\!(-\alpha,\alpha)$, for all $x\!\in\!\mathcal{B}(r)$, where $\mathcal{B}(r)$ is an open ball of the radius $r>0$, $\|\cdot\|$ - Euclidean norm.
   • One-Sided Lipschitz Condition on $\mathcal{X}$ if $\exists\ell\in L^{\infty}_{\rm loc}(\mathbb{R};\mathbb{R})$:
@@
 defines a rule of an evolution of the state $x(t)\in\mathbb{R}^{n}$ of the system in time $t\in\mathbb{R}$.
-Let $\mathcal{I}$ be one of the following intervals: $[a,b]$, $(a,b)$, $[a,b)$ or $(a,b]$, where $a,b\in[-\infty,+\infty]$.
+Let $\mathcal{I}$ be one of the following intervals: $[a,b]$, $(a,b)$, $[a,b)$ or $(a,b]$, where $a,b\in[-\infty,+\infty],a<b$.
 Definition 1 (Classical solution of ODE).
-A function $x\in C^{1}(\mathcal{I};\mathbb{R}^{n})$ is said to be a classical solution of ODE (11) if it satisfies the equation (11) everywhere on $\mathcal{I}$.
+A function $x\in C^{1}(\mathcal{I};\mathbb{R}^{n})$ is said to be a classical solution of ODE ($1$) if it satisfies the equation (11) everywhere on $\mathcal{I}$.
 The Initial Value Problem (IVP) is to find a solution (11) for a given initial state $x_{0}\in\mathbb{R}^{n}$ at a given initial time $t_{0}\in\mathbb{R}$:
 $x(t^{0})=x^{0}\in\mathbb{R}^{n}.\vskip-2.84526pt$ (12)
-For the proofs of the classical theorem surveyed in this subsection we refer the reader, for example, to [CoddingtonLevinson1955:Book], [Filippov1988:Book].
+For the proofs of the classical theorems surveyed in this subsection we refer the reader, for example, to [10], [17].
 Theorem 1 (Peano Existence Theorem, 1890).
@@
 Definition 2 (Carathéodory solution of ODE).
-An absolutely continuous function $x:\mathcal{I}\subset\mathbb{R}^{n}\to\mathbb{R}^{n}$ is said to be a Carathéodory (or strong) solution of ODE (11) if it satisfies the equation (11) almost everywhere on the time interval $\mathcal{I}$.
+An absolutely continuous function $x:\mathcal{I}\subset\mathbb{R}^{n}\to\mathbb{R}^{n}$ is said to be a Carathéodory (or strong) solution of ODE ($1$) if it satisfies the equation (11) almost everywhere on the time interval $\mathcal{I}$.
 The existence of Carathéodory solutions can be proven for ODEs with discontinuous in $t$ right-hand sides.
 Theorem 2 (Carathéodory Existence Theorem, 1918).
-f $f$ satisfies Carathéodory Condition then the IVP (11), (12) has a Carathéodory solution defined, at least, locally in time.
+If $f$ satisfies Carathéodory Condition then the IVP (11), (12) has a Carathéodory solution defined, at least, locally in time.
 The strong solutions allow some systems with deterministic noise to be modeled by ODEs, e.g., the system $\dot{x}(t)=-x(t)+\gamma(t)$ does not have a classical solution if $\gamma\in L^{\infty}(\mathbb{R};\mathbb{R})$ is a discontinuous signal, but it has the strong solution.
 Let $x(t,t^{0},x^{0})$ denote a solution of the IVP (11), (12).
-Theorem 3 (Continuity of solutions on initial data [CoddingtonLevinson1955:Book], page 58).
+Theorem 3 (Continuity of solutions on initial data [10], page 58).
 If $f$ satisfies Carathéodory Condition, then $x(t,t^{1},x^{1})\to x(t,t^{0},x^{0})$ as $(t^{1},x^{1})\to(t^{0},x^{0})$ uniformly on $\mathcal{I}$ provided that the solution $x(t,t^{0},x^{0})$ is unique on $\mathcal{I}$.
@@
 2.3 Regularization of ill-posed ODE by Filippov method
-Systems with discontinuous in $x$ right-hand sides may have Carathéodory solutions in some cases [Liberzon2003:Book].
-However, SMC theory deals with control systems modeled by ill-posed ODEs with discontinuous right-hand sides [Utkin_etal2009:Book, Liberzon2003:Book, GeligLeonovYakubovich2004:Book].
+Systems with discontinuous in $x$ right-hand sides may have Carathéodory solutions in some cases [30].
+However, SMC theory deals with control systems modeled by ill-posed ODEs with discontinuous right-hand sides [48, 30, 19].
 Example 1.
@@
 In addition, formally, if $x(t)=0$ for all $t>t^{0}+T(t^{0},x^{0})$ then $\dot{x}(t)=0=u(t)+0.5\sin(3t)$ for $t>t^{0}+T(t^{0},x^{0})$, but, by definition, $\mathrm{sign}(x(t))=0$ for $x(t)=0$.
-To define a solution of the system (11) with discontinuous (on $x$) right-hand side, a special regularization procedure known as the Filippov method [Filippov1988:Book] can be utilized.
+To define a solution of the system (11) with discontinuous (on $x$) right-hand side, a special regularization procedure known as the Filippov method [17] can be utilized.
 It extends an ill-posed ODE to a well-posed ODI
 $\dot{x}\in K[f](t,x),\vskip-2.84526pt$ (14)
-where $K[f]:\mathbb{R}\times\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$ is defined as [Filippov1988:Book, page 85]:
+where $K[f]:\mathbb{R}\times\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$ is defined as [17, page 85]:
 […]
-where $\overline{\mathrm{co}}$ denotes the convex closure, and the intersections are taken over all sets $M\subset\mathbb{R}^{n}$ of measure zero and all $\delta>0$.
-The set-valued mapping $(t,x)\to K[f](t,x)$ is nonempty-valued, compact-valued, convex-valued and upper semi-continuous in $x$ for almost all $t$ provided that $f$ is a measurable function (see, e.g., [Filippov1988:Book]).
-Filippov regularization (15) does not modify continuous in $x$ function $f$.
+where $\overline{\mathrm{co}}$ denotes the convex closure, and the intersections are taken over all sets $M\subset\mathbb{R}^{n}$ of measure zero and all $\delta>0$ and $2^{\mathbb{R}^{n}}$ denotes the power set of $\mathbb{R}^{n}$.
+The set-valued mapping $(t,x)\to K[f](t,x)$ is nonempty-valued, compact-valued, convex-valued and upper semi-continuous in $x$ for almost all $t$ provided that $f$ is a measurable function (see, e.g., [17]).
+Notice that Filippov regularization (15) does not modify continuous in $x$ function $f$.
 Definition 3 (Filippov solution).
@@
 If $f$ satisfies Filippov Condition, then the IVP (11), (12) has a Filippov solution defined, at least, locally in time.
-Continuity of Filippov solutions on initial data [Filippov1988:Book, page 98] repeats the classical result if solutions of (14) are unique.
+Continuity of Filippov solutions on initial data [17, page 98] repeats the classical result if solutions of (14) are unique.
 Theorem 6 (Continuity of Filippov solutions on initial data).
@@
 Filippov solution of a discontinuous ODE is understood as a solution of an ODI.
-The uniqueness analysis for solutions of the ODI is a non-trivial problem in the general case [Filippov1988:Book], [Tolstonogov2000:Book].
+The uniqueness analysis for solutions of the ODI is a non-trivial problem in the general case [17], [47].
 However, One-Sided Lipschitz Condition is still applicable for uniqueness analysis of Filippov solutions.
-Theorem 7 (Uniqueness of Filippov solutions, [Filippov1988:Book], §10).
+Theorem 7 (Uniqueness of Filippov solutions, [17], §10).
 If a locally one-sided Lipschitz function $f$ satisfies Filippov Condition then the IVP (11), (12) has a unique Filippov solution defined, at least, locally in the forward time.
@@
 […]
-where the set-valued function $\mathcal{N}:\mathcal{S}\to 2^{\{1,2,...,N\}}$ indicates the domains, which have a common boundary point $x\in\mathcal{S}$, i.e., $\mathcal{N}(x)=\left\{j\in\{1,2,...,N\}:x\in\partial G^{j}\right\}.$
-Let $\mathcal{C}_{\Omega}(x)\subset\mathbb{R}^{n}$ denote the (Bouligand) tangent cone to the set $\Omega\subset\mathbb{R}^{n}$ at $x\in\mathbb{R}^{n}$ defined as (see., e.g. [Nagumno1942:PPMSJ])
+where, $\mathcal{S}=\bigcup_{i=1}^{N}\partial G^{j}$ is the discontinuity set of $f$, but the set-valued function $\mathcal{N}:\mathcal{S}\to 2^{\{1,2,...,N\}}$ indicates the continuity domains $G^{j}$, which have a common boundary point $x\in\mathcal{S}$, i.e., $\mathcal{N}(x)=\left\{j\in\{1,2,...,N\}:x\in\partial G^{j}\right\}.$
+Let $\mathcal{C}_{\Omega}(x)\subset\mathbb{R}^{n}$ denote the (Bouligand) tangent cone to the set $\Omega\subset\mathbb{R}^{n}$ at $x\in\mathbb{R}^{n}$ defined as (see., e.g. [32])
 $\mathcal{C}_{\Omega}(x)=\left\{z\in\mathbb{R}^{n}:\liminf_{h\to 0^{+}}\tfrac{\inf_{y\in\Omega}\|x+hz-y\|}{h}=0\right\}.\vskip-2.84526pt$ (18)
-Tangent cones characterize positively invariant sets [^2] A set $M\!\subset\!\mathbb{R}^{n}$ is positively invariant if $x(t^{0})\!\in\!M\Rightarrow x(t)\!\in\!M,t\!\geq\!t_{0}.$ of dynamical systems (see, e.g., [Clarke_etal1995:JDCS], [Blanchini1999:Aut], [AubinEkeland1984:Book] and references therein).
+Tangent cones characterize positively invariant sets [^2] A set $M\!\subset\!\mathbb{R}^{n}$ is positively invariant if $x(t^{0})\!\in\!M\Rightarrow x(t)\!\in\!M,t\!\geq\!t_{0}.$ of dynamical systems (see, e.g., [9], [7], [2] and references therein).
 Assumption 1.
 Let a piece-wise one-sided Lipschitz continuous function $f$ be such that $f^{j}(t,x)\!\notin\!\mathcal{C}_{\bar{G}^{j}}(x)$ for all $x\!\in\!\partial G^{j}$, for almost all $t\!\in\!\mathbb{R}$ and for all j=1,2,,Nj\!=\!1,2,...,N.
-Let us denote $F_{\mathcal{S}}(t,x)=K[f](t,x)\cap\mathcal{C}_{\mathcal{S}}(x),t\in\mathbb{R},x\in\mathcal{S}$. The following result is partially inspired by [Filippov1988:Book, §10].
+The following result is partially inspired by [17, §10].
 Theorem 8 (Uniqueness of solutions in sliding mode case).
-If the set $F_{\mathcal{S}}$ is a singleton $F_{\mathcal{S}}(t,x)=\{f^{0}(t,x)\}$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$ and for some locally one-sided Lipschitz continuous function $f^{0}$, then, under Assumption 1, the IVP (11), (12) has a Filippov solution being unique, at least, locally in the forward time.
+Let
+$F_{\mathcal{S}}(t,x)=K[f](t,x)\cap\mathcal{C}_{\mathcal{S}}(x),\quad t\in\mathbb{R},\quad x\in\mathcal{S}.\vskip-2.84526pt$ (19)
+If the set $F_{\mathcal{S}}$ is a singleton, i.e., $F_{\mathcal{S}}(t,x)=\{f^{0}(t,x)\}$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$, with for some locally one-sided Lipschitz continuous function $f^{0}$, then, under Assumption 1, the IVP (11), (12) has a Filippov solution being unique, at least, locally in the forward time.
 Proof.
@@
 $\dot{x}(t)\in F_{\mathcal{S}}(t,x(t)).\vskip-2.84526pt$ (20)
-The system (16) satisfies the above theorem with $F_{\mathcal{S}}=\{0\}$.
-The relay SMC system is the conventional example of discontinuous ODE satisfying the above theorem [Utkin1992:Book], [Filippov1988:Book].
+The system (16) satisfies the above theorem with $F_{\mathcal{S}}=\{0\}$. The relay SMC system is the conventional example of discontinuous ODE satisfying the above theorem [49], [17].
 Example 2.
-The simplest example of discontinuous ODE with unique solutions can be discovered in [Utkin1992:Book] for the system
+A simple example of discontinuous ODE with unique solutions can be discovered in [49] for the SMC system
 $\dot{x}(t)=Ax(t)+Bu(t),\vskip-2.84526pt$ (21)
@@
 Assumption 2.
-Let $f$ be piece-wise one-sided Lipschitz continuous and $F_{\mathcal{S}}(t,x)$$\emptyset$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$.
+Let $f$ be piece-wise one-sided Lipschitz continuous and $F_{\mathcal{S}}(t,x)=\emptyset$ for all $x\in\mathcal{S}$, for almost all $t\in\mathbb{R}$.
 Theorem 9 (Uniqueness of solutions in switching case).
@@
 Let $f$ be piece-wise one-sided Lipschitz continuous function and there exists a closed set $\mathcal{M}\subset\mathcal{S}\subset\mathbb{R}^{n}$:
-  • $F_{\mathcal{S}}(t,x)$$\emptyset$, $\forall x\!\in\!\mathcal{S}\backslash\mathcal{M}$, for almost all $t\!\in\!\mathbb{R}$;
+  • $F_{\mathcal{S}}(t,x)=\emptyset$, $\forall x\!\in\!\mathcal{S}\backslash\mathcal{M}$, for almost all $t\!\in\!\mathbb{R}$;
   • $f^{j}(t,x)\!\notin\!\mathcal{C}_{\bar{G}^{j}}(x)$ for all $x\!\in\!\partial G^{j}\cap\mathcal{M}$, for almost all $t\!\in\!\mathbb{R}$ and for all $j\!=\!1,2,...,N:\partial G^{j}\cap\mathcal{M}\neq\emptyset$.
@@
 Under Assumption 3, if $F_{\mathcal{S}}(t,x)\!=\!\{f^{0}(t,x)\}$ is a singleton for all $x\!\in\!\mathcal{M}$ and almost all $t\!\in\!\mathbb{R}$ and $f^{0}$ is locally one-sided Lipschitz continuous, then the IVP (11), (12) has a Filippov solution being unique, at least, locally in the forward time.
-Assumption 3 ensures that $\mathcal{M}$ is the sliding set and the (switching) set $\mathcal{S}\backslash\mathcal{M}$ is crossed by trajectories without sliding.
+Assumption 3 ensures that $\mathcal{M}$ is the sliding set and the switching set $\mathcal{S}\backslash\mathcal{M}$ is crossed by trajectories without sliding.
 Example 3.
@@
 where $t$ are $x$ are as before, $f\in\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ and $u:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a measurable function (e.g., SMC).
-In this case, some alternative methods of regularization of discontinuous ODE can be proposed (e.g., [Utkin1992:Book], [Filippov1988:Book], [PolyakovFridman2014:JFI]).
+In this case, some alternative methods of regularization of discontinuous ODE can be proposed (e.g., [49], [17], [43]).
 Definition 4 (Utkin solution).
@@
 Example 4 (Neimark 1960).
-A simple analysis (see, e.g., [Neimark1961:IFAC_Congress] or [PolyakovFridman2014:JFI, Example 2]) shows that the set of Utkin solutions of the first kind do not coincide with the set of Utkin solutions of the second kinds for the linear time-invariant control system (20) with $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, $A=\mathbf{0}$, $B=I_{2}$ and the static relay feedback law $u=(u_{1},u_{2})^{\top},u_{1}=u_{2}=\mathrm{sign}(x_{1})$.
-The function $u_{eq}$ is the so-called equivalent control [Utkin1992:Book], [Filippov1988:Book], which, in SMC theory, is defined as a solution of a special algebraic equation (see below).
-The equivalent control method is the main tool for analysis and design of SMC systems [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book], [Utkin_etal2020:Book], [Orlov2020:Book].
-The existence of $u_{\rm eq}$ follows from [Filippov1962:SIAM].
+A simple analysis (see, e.g., [33] or [43, Example 2]) shows that the set of Utkin solutions of the first kind do not coincide with the set of Utkin solutions of the second kinds for the linear time-invariant control system (21) with $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, $A=\mathbf{0}$, $B=I_{2}$ and the static relay feedback law $u=(u_{1},u_{2})^{\top},u_{1}=u_{2}=\mathrm{sign}(x_{1})$.
+The function $u_{eq}$ is the so-called equivalent control [49], [17], which, in SMC theory, is defined as a solution of a special algebraic equation (see below).
+The equivalent control method is the main tool for analysis and design of SMC systems [49], [14], [46], [50], [38].
+The existence of $u_{\rm eq}$ follows from [16].
 Lemma 1 (Filippov 1959).
@@
 $\dot{x}(t)\in f(t,x(t),U(t,x(t))),\quad t\in\mathbb{R},\vskip-2.84526pt$ (24)
-has a solution (see, e.g., [Filippov1988:Book] or [PolyakovFridman2014:JFI]).
-The mapping $(t,x)\mapsto f(t,x,U(t,x))$ may be nonconvex-valued, so the conventional existence theorems for ODIs [Filippov1988:Book] are not applicable [BartoliniZolezzi1985:IEEE_TAC].
-The equivalence of Filippov and Utkin solutions for affine-in-control systems was studied in the literature [Utkin1992:Book].
+has a solution (see, e.g., [17] or [43]). The mapping $(t,x)\mapsto f(t,x,U(t,x))$ may be nonconvex-valued, so the conventional existence theorems for ODIs [17] are not applicable [6].
+The equivalence of Filippov and Utkin solutions for affine-in-control systems was studied in the literature (see, e.g., [49]).
 Theorem 10.
@@
 being fulfilled for almost all $t$, due to continuity of $a$, $b$ in $x$.
-A sufficient condition of equivalence of Utkin solutions of the first and of the second kinds is presented, for example, in [Zolezzi2002:inBook, Theorem 14, page 44] for affine-in-control systems.
-If $\mathcal{S}=\{x\in\mathbb{R}^{n}:s(x)=\mathbf{0}\}$, $s\in C^{1}(\mathbb{R}^{n};\mathbb{R}^{k})$ is a sliding set of the system (21) then the equivalent control on $\mathcal{S}$ is defined as a solution of the algebraic equation [Utkin1992:Book]:
+A sufficient condition of equivalence of Utkin solutions of the first and of the second kinds is presented, for example, in [52, Theorem 14, page 44] for affine-in-control systems.
+If $\mathcal{S}=\{x\in\mathbb{R}^{n}:s(x)=\mathbf{0}\}$, $s\in C^{1}(\mathbb{R}^{n};\mathbb{R}^{k})$ is a sliding set of the system (22) then the equivalent control on $\mathcal{S}$ is defined as a solution of the algebraic equation [49]:
 $\tfrac{\partial s(x)}{\partial x}f(t,x,u_{\rm eq})=\mathbf{0}.\vskip-2.84526pt$ (25)
@@
 $\det\left(\tfrac{\partial s(x)}{\partial x}b(t,x)\right)\neq 0,\quad\forall x\in\mathcal{S},\quad\forall t\in\mathbb{R},\vskip-2.84526pt$ (26)
-for the affine-in-control system from Theorem 10, then [Utkin1992:Book]:
+for the affine-in-control system from Theorem 10, then [49]:
 $u_{\rm eq}(t,x)=-\left(\tfrac{\partial s(x)}{\partial x}b(t,x)\right)^{-1}\tfrac{\partial s(x)}{\partial x}a(t,x),\quad x\in\mathcal{S}.\vskip-5.69054pt$ (27)
 If $(t,x)\!\mapsto\!\!f^{0}(t,x)\!=\!a(t,x)\!-\!b(t,x)\!\left(\!\frac{\partial s(x)}{\partial x}b(t,x)\!\right)^{\!\!-\!1}\!\!\frac{\partial s(x)}{\partial x}a(t,x)$ is locally one-sided Lipschitz on $\mathcal{S}$ continuous function then any Filippov solution belonging to the sliding set is unique.
-This does not mean all solutions with $x_{0}\in\mathcal{S}$ are unique, since $\mathcal{S}$ may not be a strongly positively invariant set of the system.
+This does not mean that all solutions with $x_{0}\in\mathcal{S}$ are unique, since the sliding set $\mathcal{S}$ may not be strongly positively invariant.
 3 Discontinuous Integro-differential equations
@@
 3.1 Well-posed IDEs
-Continuous integro-differential equations of the form (27) are well studied in the literature (see, e.g., [Lakshmikantham_etal1995:Book]).
+Continuous integro-differential equations of the form (28) are well studied in the literature (see, e.g., [26]).
 Below we use the notation $\mathcal{I}=[t^{0},t^{0}+\alpha)$ with $\alpha>0$ or $\mathcal{I}=[t^{0},+\infty)$.
@@
 The existence theorem of classical solutions of the integro-differential equation (28) repeats the Peano existence theorem.
-Theorem 11 (see, e.g., [Lakshmikantham_etal1995:Book]).
+Theorem 11 (see, e.g., [26]).
 If $f$, $\tilde{f}$, $\Phi$ are continuous functions then the IVP (28), (12) has a classical solution defined, at least, locally in the forward time.
-Carathéodory solutions of integro-differential equations are defined similarly to solutions of ODEs [OReganMeehan1998:Book].
+Carathéodory solutions of integro-differential equations are defined similarly to solutions of ODEs [34].
 Definition 6 (Carathéodory solution of IDE).
 An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ with $x(t^{0})=x^{0}$ is said to be a Carathéodory (or strong) solution of the IVP (28), (12) if it satisfies (28) almost everywhere on $\mathcal{I}$.
-Carathéodory solutions of IDEs are well-studied too [OReganMeehan1998:Book].
+Carathéodory solutions of IDEs are well-studied too [34].
 Theorem 12.
 Let Assumption 4 hold. If $f~\tilde{f}$ satisfy Carathéodory Condition then the IVP (28), (12) has a Carathéodory solution defined, at least, locally in the forward time.
-Similarly to ODEs, a Lipschitz condition is sufficient for uniqueness of solutions of IDEs (see, e.g., [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book]).
+Similarly to ODEs, a Lipschitz condition is sufficient for uniqueness of solutions of IDEs (see, e.g., [26], [34]).
 Theorem 13.
@@
 Definition 7 (Filippov solution of IDE).
-An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ is said to be a Filippov solution of the IVP (27), (12) if $x(t^{0})=x^{0}$ and, almost everywhere on $\mathcal{I}$, it satisfies the integral differential inclusion [^4] The integral of the set-valued function is understood in the sense of Aumann [Aumann1965:JMAA].
+An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ is said to be a Filippov solution of the IVP (28), (12) if $x(t^{0})=x^{0}$ and, almost everywhere on $\mathcal{I}$, it satisfies the integral differential inclusion [^4] The integral of the set-valued function is understood in the sense of Aumann [3].
 $\begin{split}\dot{x}(t)\in&K[f](t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)K[\tilde{f}](\tau,x(\tau))d\tau.\vskip-2.84526pt\end{split}$ (29)
@@
 Proof.
-1) The set-valued mappings $K[f]$ and $K[\tilde{f}$] are nonempty-valued, convex-valued, compact-valued and upper semi-continuous in $x$ for almost all $t$ (see, e.g., [Filippov1988:Book, page 85]).
+1) The set-valued mappings $K[f]$ and $K[\tilde{f}$] are nonempty-valued, convex-valued, compact-valued and upper semi-continuous in $x$ for almost all $t$ (see, e.g., [17, page 85]).
 Let us define the functions $f^{\eta},\tilde{f}^{\eta}:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}$ as follows
 $f^{\eta}(t,x)\!=\!\tfrac{1}{\eta}\!\!\int\limits_{\mathcal{B}(\eta)}\!\!\!f(t,x+y)dy,\;\;\;\tilde{f}^{\eta}(t,x)\!=\!\tfrac{1}{\eta}\!\!\int\limits_{\mathcal{B}(\eta)}\!\!\tilde{f}(t,x+y)dy,\vskip-2.84526pt$
-where $\eta>0$. In [Filippov1988:Book, Theroem 8, page 85] it is shown that these functions satisfy Carathéodory Condition and
+where $\eta>0$. In [17, Theroem 8, page 85] it is shown that these functions satisfy Carathéodory Condition and
 […]
@@
 $\dot{x}^{\eta}(t)=f^{\eta}(t,x)+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{f}^{\eta}(\tau,x^{\eta}(\tau))d\tau\vskip-2.84526pt$ (30)
-has a Carathéodory solution. By definition of Aumann’s Integral [Aumann1965:JMAA], for almost all $t\in\mathcal{I}$, this solution satisfies
+has a Carathéodory solution. By definition of Aumann’s Integral [3], for almost all $t\in\mathcal{I}$, this solution satisfies
 $\dot{x}^{\eta}(t)\in F^{\eta}(t,x^{\eta}(t))+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{F}^{\eta}(\tau,x^{\eta}(\tau))d\tau.\vskip-2.84526pt$ (31)
@@
 Let $\bar{\eta}=\min\{1/4,1/(4\overline{\Phi}\alpha)\}$.
-Since, by Filippov Condition, for any $\alpha>0$ and any $r>0$ there exist a Lebesgue integrable function $m:[t^{0},t^{0}+\alpha]\to\mathbb{R}_{+}$ such that $\|f(t,x)\|\leq m(t)$ and $\|{\tilde{f}}(t,x)\|\leq m(t)$ for almost all $t\in[t^{0},t^{0}+\alpha]$ and almost all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$ then (see, e.g., [Filippov1988:Book, page 76])
+Since, by Filippov Condition, for any $\alpha>0$ and any $r>0$ there exist a Lebesgue integrable function $m:[t^{0},t^{0}+\alpha]\to\mathbb{R}_{+}$ such that $\|f(t,x)\|\leq m(t)$ and $\|\tilde{f}(t,x)\|\leq m(t)$ for almost all $t\in[t^{0},t^{0}+\alpha]$ and almost all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$ then (see, e.g., [17, page 76])
 $K[f](t,x)\subset\mathcal{B}(m(t)),\quad K[\tilde{f}](t,x)\subset\mathcal{B}(m(t))\vskip-2.84526pt$
-for almost all $t\in\mathcal{I}=[t^{0},t^{0}+\alpha]$ and for all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$, and (see, e.g., [Filippov1988:Book, page 82])
+for almost all $t\in\mathcal{I}=[t^{0},t^{0}+\alpha]$ and for all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$, and (see, e.g., [17, page 82])
 $F^{\eta}(t,x)\subset\mathcal{B}(m(t)+\eta),\quad\tilde{F}^{\eta}(t,x)\subset\mathcal{B}(m(t)+\eta)\vskip-2.84526pt$
@@
 for almost all $t\in\mathcal{I}$ and all $\eta\in(0,\bar{\eta}]$.
-Since $m$ is Lebesgue integrable on $\mathcal{I}$ then $M:=\int^{t^{0}+\alpha}_{t^{0}}m(\tau)d\tau<+\infty$ and for any $\epsilon>0$ there exist $\delta\in(0,\epsilon\min\{1,1/(4\overline{\Phi}M)\})$ such that for all $a,b]\subset\mathcal{I}:|b-a|\leq\delta$ and all $\eta\in\left(0,\bar{\eta}\right)$ it holds
+Since $m$ is Lebesgue integrable on $\mathcal{I}$ then $M:=\int^{t^{0}+\alpha}_{t^{0}}m(\tau)d\tau<+\infty$ and for any $\epsilon>0$ there exist
+$\delta\in(0,\epsilon\min\{1,1/(4\overline{\Phi}M)\})$
+such that for all $a,b]\subset\mathcal{I}:|b-a|\leq\delta$ and all $\eta\in\left(0,\bar{\eta}\right)$ it holds
 $\begin{split}\|x^{\eta}(b)-x^{\eta}(a)\|\leq&\left\|\int^{b}_{a}{\dot{x}}^{\eta}(s)ds\right\|\leq\int^{b}_{a}\|{\dot{x}}^{\eta}(s)\|ds\\ \leq&\int^{b}_{a}m(s)ds+(b-a)(\eta+\overline{\Phi}M+\overline{\Phi}\alpha\eta)\\ \leq&\frac{\epsilon}{4}+\delta\eta+\delta\overline{\Phi}M+\delta\overline{\Phi}\alpha\eta\leq\epsilon.\vskip-2.84526pt\end{split}$
 This means that the family of function $\{x^{\eta}\},\eta\in(0,\bar{\eta}]$ is equicontinuous and uniformly bounded for all $t\in\mathcal{I}$,
-$\|x^{\eta}(t)\|\leq\|x_{0}\|+\int^{t}_{t_{0}}\|\dot{x}^{\eta}(\tau)\|d\tau\leq M+\alpha\eta+\overline{\Phi}(M\alpha+\alpha^{2}\eta),\vskip-2.84526pt$
+$\|x^{\eta}(t)\|-\|x_{0}\|\leq\int^{t}_{t_{0}}\|\dot{x}^{\eta}(\tau)\|d\tau\leq M+\alpha\eta+\overline{\Phi}(M\alpha+\alpha^{2}\eta),\vskip-2.84526pt$
 so, by Arzelà–Ascoli theorem, it contains a sequence $x^{\eta_{k}}$ uniformly convergent to a continuous function $x$ as $\eta_{k}\to\mathbf{0}$.
 The above inequalities imply that $x$ is absolutely continuous.
-3) The set-valued mappings $F^{\eta}$ and $\tilde{F}^{\eta}$ are nonempty-valued, compact-valued, convex-valued and upper semi-continuous continuous in $x$ for almost all $t$ (see, e.g., [Filippov1988:Book, page 86]).
+3) The set-valued mappings $F^{\eta}$ and $\tilde{F}^{\eta}$ are nonempty-valued, compact-valued, convex-valued and upper semi-continuous continuous in $x$ for almost all $t$ (see, e.g., [17, page 86]).
 For any solution $x^{\eta}$ of the IDE (30), the set-valued mappings $t\to F^{\eta}(t,x^{\eta}(t))$ and $\tau\to\Phi(t,\tau)\tilde{F}^{\eta}(t,x^{\eta}(t))$ are compact-valued, convex-valued, measurable and bounded by the integrable function $t\mapsto\bar{\Phi}(m(t)+\eta)$.
-By [Aumann1965:JMAA, Theorems 1, 2 and 4] [^5] The requirement of the paper [Aumann1965:JMAA] about Borel graph of the integrated function has been relaxed (see, e.g., [CastaingValadier1977:Book, Chapter 3]) using Kuratowski–Ryll-Nardzewski measurable selection theorem published in the same year [KuratowskiRyll-Nardzewski1965:BAPC]., the right-hand side of (30)
+By [3, Theorems 1, 2 and 4] [^5] The requirement of the paper [3] about Borel graph of the integrated function has been relaxed (see, e.g., [8, Chapter 3]) using Kuratowski–Ryll-Nardzewski measurable selection theorem published in the same year [25]., the right-hand side of (31)
 $Q^{\eta}(t):=F^{\eta}(t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{F}^{\eta}(\tau,x(\tau))d\tau\subset\mathbb{R}^{n}\vskip-2.84526pt$
 is convex-valued and compact-valued for almost all $t$.
-Moreover, by construction, for almost all $t$ and for all $x$, we have $K[f](t,x)\subset F^{\eta}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{\eta}(t,x)$, $\forall\eta\in[0,\overline{\eta}]$, and $F^{\eta}(t,x)\to K[f](t,x)$, $\tilde{F}^{\eta}(t,x)\to K[\tilde{f}](t,x)$ as $\eta\to 0$, where the limit is understood in the Hausdorff topology (see, e.g., [Filippov1988:Book, §5]).
-By [Aumann1965:JMAA, Theprem 5], for almost all $t\in\mathcal{I}$, we conclude that $Q^{\eta_{k}}(t)\to Q(t)$ as $\eta_{k}\to 0$, where
+Moreover, by construction, for almost all $t$ and for all $x$, we have $K[f](t,x)\subset F^{\eta}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{\eta}(t,x)$, $\forall\eta\in[0,\overline{\eta}]$, and $F^{\eta}(t,x)\to K[f](t,x)$, $\tilde{F}^{\eta}(t,x)\to K[\tilde{f}](t,x)$ as $\eta\to 0$, where the limits are understood in the Hausdorff topology (see, e.g., [17, §5]).
+By [3, Theprem 5], for almost all $t\in\mathcal{I}$, we conclude that $Q^{\eta_{k}}(t)\to Q(t)$ as $\eta_{k}\to 0$, where
 $Q(t)=K[f](t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)K[\tilde{f}](\tau,x(\tau))d\tau\subset\mathbb{R}^{n}\vskip-2.84526pt$
@@
 almost everywhere on $t^{0},t^{0}+\alpha]$. This means $\dot{x}(t)\in Q(t)$ for almost all $t\in\mathcal{I}$. The proof is complete. ∎
-Under some assumptions, the integro-differential equation (27) can be interpreted as a functional differential equation of retarded type [Hale1977:Book], [Fridman2014:Book].
+Under some assumptions, the integro-differential equation (28) can be interpreted as a functional differential equation of retarded type [21], [18].
 Indeed, if, there exist $d>0$ such that $\Phi(t,\tau)=0$ for $t-\tau>d$, then $\dot{x}(t)$ depends only on $x(\tau)$ with $\tau\in[\max\{t^{0},t-d\},t]$, $t\geq t^{0}$.
-Generalized solutions of discontinuous retarded functional differential equations are introduced in [KolmanovskiiMyshkis1992:Book, page 58] by means of a limiting procedure in a function space, which does not have such a simple formalization by the Filippov regularization (15).
+Generalized solutions of discontinuous retarded functional differential equations are introduced in [24, page 58] by means of a limiting procedure in a function space, which does not have such a simple formalization by the Filippov regularization (15).
 A study of links between the generalized solutions and the Filippov solutions is an interesting problem for the future research.
@@
 $\tilde{F}^{*}_{\mu}(t,x)=\overline{\mathrm{co}}\bigcup_{y\in\mathcal{B}(m_{\mu}(t)),z\in\mathcal{B}(m_{\mu}(t))}K[\tilde{f}](t,x+y)+z.\vskip-2.84526pt$
-For almost all $t\in(t^{0},t^{0}+T)$ and almost all $x\in x^{0}+\mathcal{B}(\bar{r})$, by construction $K[f](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$, and $K[f_{\mu}](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$ due to positive measure of $E(t,x)$ (see, [Filippov1988:Book, page 98] for more details).
+For almost all $t\in(t^{0},t^{0}+T)$ and almost all $x\in x^{0}+\mathcal{B}(\bar{r})$, by construction $K[f](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$, and $K[f_{\mu}](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$ due to positive measure of $E(t,x)$ (see, [17, page 98] for more details).
 Since $r_{\mathcal{M}}=\sup_{\mu\in\mathcal{M}}\|m_{\mu}\|_{L^{\infty}(t^{0},t^{0}+T)}<+\infty$ then, due to Filippov Condition, there exists $m\in L^{1}((t^{0},t^{0}+T),\mathbb{R}_{+})$ such that $\sup_{y\in F^{*}_{\mu}(t,x)}\|y\|\leq m(t)+m_{\mu}(t)$ for almost all $t\in(t^{0},t^{0}+T)$, for all $x\in x^{0}+\mathcal{B}(\bar{r}+r_{\mathcal{M}})$ and for all $\mu\in\mathcal{M}$.
@@
 and $\|\dot{x}_{\mu}(t)\|\in m(t)+m_{\mu}(t)+\overline{\Phi}\int^{t}_{t^{0}}m(\tau)+m_{\mu}(\tau)\,d\tau$ almost everywhere, where $\overline{\Phi}\geq 0$ is defined in the proof of Theorem 14.
-Repeating the step 2) of the proof of Theorem 14, we show that the family of solutions of (33) with $\mu\in\mathcal{M}$ is uniformly bounded and equicontinuous.
-Since, by assumption, $\|m_{\mu}\|_{L^{1}}\to 0$ as $\mu\to\mu_{0}$, then, repeating the step 3) of the proof of Theorem 14, we conclude that, for any uniformly convergent sequence of solutions $x_{\mu_{i}}\to x_{\mu_{0}}$ as $\mu_{i}\to\mu_{0}$, the limit $x_{\mu_{0}}$ is a Filippov solution of the IVP (27), (12).
-Applying [Filippov1988:Book, Lemma 6, page 10] we complete the proof.
+Hence, for any $a,b]\subset(t^{0},t^{0}+T)$, we have
+$\vskip-2.84526pt\begin{split}\|x_{\mu}(b)\!-\!x_{\mu}(a)\|\!\leq\!&\left\|\int^{b}_{a}{\dot{x}}^{\eta}(s)ds\right\|\leq\int^{b}_{a}\|{\dot{x}}^{\eta}(s)\|ds\\ \leq&\int^{b}_{a}\!\!m(s)ds\!+\!(b-a)(r_{\mathcal{M}}\!+\!\overline{\Phi}M\!+\!T\overline{\Phi}r_{\mathcal{M}}),\end{split}$
+where $M=\int^{t^{0}+T}_{t^{0}}m(\tau)d\tau<+\infty$.
+This implies the equicontinuity of the family $\{x_{\mu}\}_{\mu\in\mathcal{M}}$ of absolutely continuous functions (see, the step 2 of the proof of Theorem 14 for more details).
+Similarly, for all $t\in(t^{0},t^{0}+T)$, we derive
+$\vskip-2.84526pt\begin{split}\|x_{\mu}(t)\|\leq&\|x_{0}\|+\delta+\int^{t}_{t^{0}}\|\dot{x}_{\mu}(\tau)\|d\tau\\ \leq&\|x_{0}\|+\delta+M+T(r_{\mathcal{M}}+\overline{\Phi}M\!+\!T\overline{\Phi}r_{\mathcal{M}}),\end{split}$
+i.e., the family $\{x_{\mu}\}_{\mu\in\mathcal{M}}$ is uniformly bounded. Since, by assumption, $\|m_{\mu}\|_{L^{1}}\to 0$ as $\mu\to\mu_{0}$, then,
+for almost all $t$ and for all $x$, we have $F^{*}_{\mu}(t,x)\to K[f](t,x)$, $\tilde{F}^{*}_{\mu}(t,x)\to K[\tilde{f}](t,x)$ as $\mu\to\mu_{0}$, where the limits are understood in the Hausdorff topology.
+By [3, Theprem 5], for almost all $t\in\mathcal{I}$, we conclude that $Q^{*}_{\mu}\to Q(t)$ as $\mu\to\mu_{0}$, where
+$\vskip-2.84526pt\begin{split}Q^{*}_{\mu}(t)=&F^{*}_{\mu}(t,x_{\mu}(t))+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{F}^{*}_{\mu}(\tau,x_{\mu}(\tau))d\tau\subset\mathbb{R}^{n},\\ Q(t)=&K[f](t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)K[\tilde{f}](\tau,x(\tau))d\tau\subset\mathbb{R}^{n}\end{split}$
+are convex and compact sets.
+Therefore, for any uniformly convergent sequence of solutions $x_{\mu_{i}}\to x_{\mu_{0}}$ as $\mu_{i}\to\mu_{0}$, the limit $x_{\mu_{0}}$ is a Filippov solution of the IVP (28), (12) (see, the step 3 of the proof of Theorem 14 for more details).
+Applying [17, Lemma 6, page 10] we complete the proof.
 ∎
@@
 If $x_{1},x_{2}$ are two solutions of the IVP (28), (12) then, for $e(t)=x_{1}(t)-x_{2}(t)$ and for almost all $t$, it holds
-$\begin{split}\tfrac{d\|e(t)\|^{2}}{dt}=&2e^{\top}(t)(f_{1}(t)-f_{2}(t))\\ &+\!2\!\!\int^{t}_{t^{0}}\!\!\!e^{\top}\!(t)\Phi(t,\tau)\tilde{f}(\tau,x_{1}(\tau)\!-\!\tilde{f}(\tau,x_{2}(\tau))d\tau\\ \leq&2\ell(t)\|e(t)\|^{2}+2\bar{\Phi}\|e(t)\|\int^{t}_{t^{0}}\tilde{\ell}(\tau)\|e(\tau)\|d\tau\\ \leq&(2\ell(t)+\bar{\Phi}^{2})\|e(t)\|^{2}+\|\tilde{\ell}\|_{L^{2}}^{2}\int^{t}_{t^{0}}\|e(\tau)\|^{2}d\tau,\end{split}\vskip-2.84526pt$ (35)
-where $f_{1}(t)\in K[f](t,x_{1}(t)),f_{2}(t)\in K[f](t,x_{2}(t))$.
-One-Sided Lipschitz Condition for $K[f]$ [Filippov1988:Book, page 106], Lipschitz Condition for $\tilde{f}$ and Cauchy–Schwartz inequality in $L^{2}$ have been utilized on the last steps.
-Since, $e(t^{0})=0$ then the obtained inequality implies e(t)=0e(t)\!=\!0 for $t\!>\!t^{0}$ (see [Pachpatte1975:JMAA]).
+$\small\begin{split}\!\!\!\!\!\tfrac{d\|e(t)\|^{2}}{dt}\!=&2e^{\!\top}\!(t)\!\left(\!f_{1}(t)\!-\!f_{2}(t)\!+\!\!\!\int^{t}_{t^{0}}\!\!\Phi(t,\tau)\left(\tilde{f}_{1}(\tau)\!-\!\tilde{f}_{2}(\tau)\right)d\tau\!\right)\\[-5.69054pt] \leq&2\ell(t)\|e(t)\|^{2}+2\bar{\Phi}\|e(t)\|\int^{t}_{t^{0}}\tilde{\ell}(\tau)\|e(\tau)\|d\tau\\[-5.69054pt] \leq&2\ell(t)\|e(t)\|^{2}\!+\!(\bar{\Phi}\|e(t)\|)^{2}\!+\!\left(\int^{t}_{t^{0}}\!\!\tilde{\ell}(\tau)\|e(\tau)\|d\tau\right)^{2}\\[-5.69054pt] \leq&2\ell(t)\|e(t)\|^{2}+\bar{\Phi}^{2}\|e(t)\|^{2}+\|\tilde{\ell}\|_{L^{2}}^{2}\!\!\int^{t}_{t^{0}}\!\!\|e(\tau)\|^{2}d\tau,\end{split}\!\!\!\!\vskip-2.84526pt$ (36)
+where $f_{1}(t)\in K[f](t,x_{1}(t)),f_{2}(t)\in K[f](t,x_{2}(t))$, $\tilde{f}_{1}(\tau)=\tilde{f}(\tau,x_{1}(\tau))$, $\tilde{f}_{2}(\tau)=\tilde{f}(\tau,x_{2}(\tau))$, $\overline{\Phi}\geq 0$ is defined in the proof of Theorem 14.
+One-Sided Lipschitz Condition for $f$ (and, consequently, for $K[f]$, see [17, page 106]) with $\ell\in L^{\infty}_{\rm loc}$, Lipschitz Condition for $\tilde{f}$ with $\tilde{\ell}\in L^{\infty}_{\rm loc}$ and Cauchy–Schwartz inequality in $L^{2}((0,t);\mathbb{R})$ have been utilized to derive (36).
+Since, $e(t^{0})=0$ then the obtained inequality implies e(t)=0e(t)\!=\!0 for $t\!>\!t^{0}$ (see [39]).
 ∎
-For discontinuous $\tilde{f}$, the uniqueness analysis is challenging.
-Let us denote $F_{\mathcal{S}}(t,x,\gamma)\!=\!\left(K[f](t,x)+\gamma\right)\,\cap\,\mathcal{C}_{\mathcal{S}}(x)$, where $\gamma\in\mathbb{R}^{n}$ and $\mathcal{S}\subset\mathbb{R}^{n}$ is the discontinuity set of $f$.
+Let $F_{\mathcal{S}}(t,x,\gamma)\!=\!\left(K[f](t,x)+\gamma\right)\,\cap\,\mathcal{C}_{\mathcal{S}}(x)$, where $\gamma\in\mathbb{R}^{n}$, $\mathcal{S}\!\subset\!\mathbb{R}^{n}$ is discontinuity set of $f$ and $\mathcal{C}_{\mathcal{S}}$ is tangent cone for $\mathcal{S}$.
 Theorem 17 (Uniqueness of solution in sliding mode case).
@@
     $K[f](t,x)\cap\mathcal{C}_{\tilde{\mathcal{S}}}(x)\!=\!\emptyset$ for all $x\!\in\!\tilde{\mathcal{S}}$ and almost all $t\!\in\!\mathbb{R}$;
-then then the IVP (27), (12) has a Filippov solution being unique, at least, locally in the forward time.
+then the IVP (28), (12) has a Filippov solution being unique, at least, locally in the forward time.
 Proof.
 Let us study the case $x^{0}\in\mathcal{S}$.
-This means that $x^{0}\in\partial G^{j}$ for all $j\in\mathcal{N}(x^{0})$, where, as before, $\mathcal{N}:\mathbb{R}^{n}\to 2$ is the indicator function (i.e., $j\in\mathcal{N}(x_{0})$ if $x_{0}\in\overline{G}^{j}$).
+This means that $x^{0}\in\partial G^{j}$ for all $j\in\mathcal{N}(x^{0})$, where, as before, $\mathcal{N}:\mathcal{S}\to 2^{\{1,2,...,N\}}$ is the indicator function, i.e., $j\in\mathcal{N}(x_{0})$ if $x_{0}\in\partial G^{j}$.
 Since the function $f$ is piece-wise continuous then there exists $\bar{\epsilon}>0$ such that $\mathcal{N}(x^{0}+\delta)\subset\mathcal{N}(x^{0})$ for $x_{0}+\delta\in S:\|\delta\|\leq\bar{\epsilon}$.
@@
 Hence, taking into account, $f^{j}\in C(\mathbb{R}\times\mathbb{R}^{n};\mathbb{R}^{n})$ we conclude that there exists $\epsilon^{j}>0$ such that
-$f^{j}(t^{0}+s,x^{0}+B(\epsilon^{j}))+B(\epsilon^{j})\subset\left(\mathbb{R}^{n}\backslash{\mathcal{C}}_{\bar{G}^{j}}(x^{0})\right)\vskip-2.84526pt$
-for all $s\in(0,\epsilon^{j})$.
-Let $\epsilon=\min\left\{\bar{\epsilon},{\hat{\epsilon}},\min_{j\in\mathcal{N}(x^{0})}\{\epsilon^{j}\}\right\}$.
-Repeating considerations of the proof of Theorem 8, we conclude that any solution $x^{\epsilon}$ of the IVP
-$\dot{x}^{\epsilon}(t)\in K[f](t,x^{\epsilon}(t))+B(\epsilon),\quad x^{\epsilon}(t^{0})=x^{0}\in\mathcal{S}\vskip-2.84526pt$ (36)
-belongs to $\mathcal{S}$ as long as $x^{\epsilon}(t)\in x^{0}+B(\epsilon)$. For any continuous function $t\in[t^{0},t^{0}+\epsilon]\to x(t)\in\mathbb{R}^{n}$, the set-valued function
+$f^{j}(t^{0}+s,x^{0}+\mathcal{B}(\epsilon^{j}))+\mathcal{B}(\epsilon^{j})\subset\left(\mathbb{R}^{n}\backslash{\mathcal{C}}_{\bar{G}^{j}}(x^{0})\right),\;\forall s\in(0,\epsilon^{j}),\vskip-2.84526pt$
+where $\mathcal{B}(\epsilon^{j})\subset\mathbb{R}^{n}$ is an open ball of the radius $\epsilon^{j}>0$.
+Let $\epsilon=\min\left\{\bar{\epsilon},\hat{\epsilon},\min_{j\in\mathcal{N}(x^{0})}\{\epsilon^{j}\}\right\}$. Repeating the proof of Theorem 8, we conclude that any solution $x^{\epsilon}$ of the IVP
+$\dot{x}^{\epsilon}(t)\in K[f](t,x^{\epsilon}(t))+\mathcal{B}(\epsilon),\quad x^{\epsilon}(t^{0})=x^{0}\in\mathcal{S}\vskip-2.84526pt$ (37)
+belongs to $\mathcal{S}$ as long as $x^{\epsilon}(t)\in x^{0}+\mathcal{B}(\epsilon)$. For any continuous function $t\in[t^{0},t^{0}+\epsilon]\to x(t)\in\mathbb{R}^{n}$, the set-valued function
 $t\in[t^{0},t^{0}+\epsilon]\mapsto\tilde{Q}(t):=\int^{t}_{t^{0}}\Phi(t,\tau)K[\tilde{f}](\tau,x(\tau))d\tau\subset\mathbb{R}^{n}\vskip-2.84526pt$
@@
 A Filippov solution of the IVP
-$\left\{\begin{split}\dot{x}_{1}(t)=&\;u(x(t)),\quad\quad\quad\quad\quad\quad\quad\quad\quad x_{1}(0)=0,\\ \dot{x}_{2}(t)=&\int^{t}_{0}{\sin(t+\tau)}u(x(\tau))d\tau,\quad\quad x_{2}(0)=0,\\ u(x)=&-\mathrm{sign}(x_{1}),\quad\quad\quad\;\;x=(x_{1},x_{2})^{\top}\in\mathbb{R}^{2}\end{split}\right$.
+$\vskip-2.84526pt\left\{\begin{split}\dot{x}_{1}(t)=&\;u(x(t)),\quad\quad\quad\quad\quad\quad\quad\quad\quad x_{1}(0)=0,\\[-2.84526pt] \dot{x}_{2}(t)=&\int^{t}_{0}\sin(t+\tau)u(x(\tau))d\tau,\quad\quad x_{2}(0)=0,\\[-2.84526pt] u(x)=&-\mathrm{sign}(x_{1}),\quad\quad\quad\;\;x=(x_{1},x_{2})^{\top}\in\mathbb{R}^{2}\end{split}\right$.
 (38)
@@
 ∎
-As a straightforward consequence of the two above theorems we deduce the following corollary.
+The corollary given below follows from above theorems.
 Corollary 2.
@@
 then the IVP (28), (12) has a Filippov solution being unique, at least, locally in the forward time.
-Remark 1 (On explicit Euler method for IDEs).
-If $f,\tilde{f}$ are piece-wise locally Lipschitz continuous functions in $x$ then approximate (numerical solution) of the IDE (27) can be found using the explicit Euler method
-$\begin{split}x_{k+1}=&x_{k}+hf(t^{0}+kh,x_{k})+hI_{k},\\ I_{k}=&h\sum^{k}_{i=0}\Phi(t^{0}+kh,t^{0}+ih)\tilde{f}(t^{0}+ih,x_{i}),\vskip-2.84526pt\end{split}$ (38)
-where $h>0$ is the sampling period, $k=0,1,2,$..., $x_{k}$ approximates $x(t^{0}+kh)$ and $I_{k}$ approximates the integral $\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{f}(\tau,x(\tau))d\tau$ by means of the method of rectangles.
-Indeed, since, on continuity domains, both functions $f,\tilde{f}$ are locally Lipschitz continuous in $x$, any solution $x:\mathcal{I}\to\mathbb{R}^{n}$ of IDE (27) belonging to this domain is unique and the sequence of the Euler approximates
-$x_{h}(t)\!=\!\tfrac{t^{0}\!+(k+1)h-t}{h}x_{k}+\tfrac{t-t^{0}\!-kh}{h}x_{k+1},t\!\in\![t^{0}\!+kh,t^{0}\!+(k+1)h],$
-(being uniformly bounded and equicontinuous on $\mathcal{I}$) converges to $x$ uniformly on $\mathcal{I}$ as $h\to 0$.
-This conclusion remains valid if the discontinuous IDE (27) has no sliding mode.
-In the sliding mode case, similarly to ODEs, the explicit Euler method correctly approximates the Filippov solution of the IDE away an $\epsilon$-neighborhood of a sliding set $\mathcal{S}$, where $\epsilon\to 0$ as $h\to 0$.
-A consistent discrete-time approximation of the motion on the sliding surface requires development of advanced numerical schemes [AcaryBrogliato2010:SCL], [Polyakov_etal2019:SIAM_JCO].
-This goes out of the scope of the paper.
+Theorems 17, 18 and Corollary 2 prove only a local-in-time uniqueness of solutions.
+For ODEs, due to the semigroup property of solutions (see, e.g., [10]), the local-in-time uniqueness implies the uniqueness on the whole interval of the existence of solution.
+For IDEs, the semigroup property is not valid, so, a Filippov solution of IDE may be unique on a short interval of time and non-unique on a long interval of time.
+Utkin solutions introduced below solve this problem in some cases.
 3.4 Equivalent control method for systems modeled by IDEs
@@
 almost everywhere on $\mathcal{I}$, where $U=K[u]$ is defined by (15) (resp., $U=(K[u_{1}],...,K[u_{m}])^{\top}$ for $u=(u_{1},u_{2},...,u_{m})^{\top}$).
-Similarly to ODEs, the set of Utkin solutions of the first kind is a subset of the set of the Utkin solutions of the second kind, since $K[u](t,x(t))\subset(K[u_{1}](t,x(t)),...,K[u_{m}(t,x(t))])^{\top}$.
+Similarly to ODEs, the set of Utkin solutions of the first kind is a subset of the set of the Utkin solutions of the second kind, since $K[u](t,x(t))\subset(K[u_{1}](t,x(t)),...,K[u_{m}](t,x(t)))^{\top}$.
 They always coincide for single input systems (i.e., if $m=1$).
@@
 If control inputs $u_{1}$ and $u_{2}$ are assumed to be independent (e.g., realized by different actuators), then, in practice, the input signals cannot be identical.
-In this particular case, the Utkin solution of the second kind (40) provides a more realistic description of the system motion than the Utkin solution of the first kind.
+In this particular case, the Utkin solution of the second kind provides a more realistic description of the system motion than the Utkin solution of the first kind.
+Utkin solutions of the first kind for an affine-in-control IDE with a smooth sliding surface may be unique globally in time.
 Corollary 3 (On equivalent control on smooth surface).
@@
     where $\mathcal{K}:L^{\infty}((t^{*},t^{1});\mathbb{R}^{m})\to L^{\infty}((t^{*},t^{1});\mathbb{R}^{m})$ is a linear bounded operator
-    $(\mathcal{K}\tilde{u})(t)\!=\!\!\int^{t}_{t^{*}}\!\!\!G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\tilde{u}(\tau)d\tau,\vskip-2.84526pt$ (42)
+    $\small(\mathcal{K}\tilde{u})(t)\!=\!\!\int^{t}_{t^{*}}\!\!\!G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\tilde{u}(\tau)d\tau,\vskip-2.84526pt$ (42)
     and $g\!\in\!L^{\infty}((t^{*},t^{1});\mathbb{R}^{n})$ is as follows
-    $g(t)\!=\!-G^{-1}(t,x(t))\left.\tfrac{\partial s(y)}{\partial y}\right|_{y=x(t)}\!\left(a(t,x(t))\!+\!\tilde{g}(t)\right),\vskip-5.69054pt$
-    $\tilde{g}(t)\!=\!\int^{t^{*}}_{t^{0}}\!\!\!\!\!\Phi(t,\tau)\left(a(\tau,x(\tau))\!+\!b(\tau,x(\tau))u(t,x(\tau)\right)d\tau.\vskip-2.84526pt$
+    $\small\!\!g(t)\!=\!-G^{-1}(t,x(t))\left.\tfrac{\partial s(y)}{\partial y}\right|_{y=x(t)}\!\left(a(t,x(t))\!+\!\tilde{g}(t)\right),\vskip-5.69054pt$
+    $\small\tilde{g}(t)\!=\!\int^{t^{*}}_{t^{0}}\!\!\!\!\!\Phi(t,\tau)\left(a(\tau,x(\tau))\!+\!b(\tau,x(\tau))u(t,x(\tau)\right)d\tau.\vskip-2.84526pt$
   • ii)
-    If $a,b,\tilde{a},\tilde{b},\frac{\partial s}{\partial x}$ are locally Lipschitz continuous and $G(t,y_{1})=G(t,y_{2}),\tilde{b}(t,y_{1})=\tilde{b}(t,y_{2})$, $\forall y_{1},y_{2}\in\mathcal{S}$ then any Utkin solution of (8), (12) belonging to $\mathcal{S}$ is unique.
+    If $a,b,\tilde{a},\tilde{b},\frac{\partial s}{\partial x}$ are locally Lipschitz continuous and
+    $G(t,y_{1})=G(t,y_{2}),\tilde{b}(t,y_{1})=\tilde{b}(t,y_{2}),\quad\forall y_{1},y_{2}\in\mathcal{S}$
+    then any Utkin solution of (8), (12) belonging to $\mathcal{S}$ is unique.
 Proof.
 i) If $x(t)\!\in\!\mathcal{S}$ then $\frac{ds(x(t))}{dt}\!=\!\left.\frac{\partial s(y)}{\partial y}\right|_{y=x(t)}\dot{x}(t)\!=\!0$ and, taking into account,
-$\begin{split}\dot{x}(t)\!=&a(t,x(t))+b(t,x(t))u_{\rm eq}(t)\\ &+\!\int^{t}_{t^{0}}\!\!\!\Phi(t,\tau)\left(\tilde{a}(\tau,x(\tau))\!+\!\tilde{b}(\tau,x(\tau))u_{\rm eq}(\tau)\right)d\tau,\vskip-5.69054pt\end{split}$ (43)
+$\small\begin{split}\dot{x}(t)\!=&a(t,x(t))+b(t,x(t))u_{\rm eq}(t)\\[-2.84526pt] &+\!\int^{t}_{t^{0}}\!\!\!\Phi(t,\tau)\left(\tilde{a}(\tau,x(\tau))\!+\!\tilde{b}(\tau,x(\tau))u_{\rm eq}(\tau)\right)d\tau,\vskip-5.69054pt\end{split}$ (43)
 we conclude that, in the sliding mode, $u_{\rm eq}$ is the unique solution of the integral Volterra equation of the second kind
-$u_{\rm eq}(t)\!=\!g(t)\!+\!(\mathcal{K}u_{\rm eq})(t),\quad t\in[t^{*},t^{1}].\vskip-2.84526pt$ (44)
+$\small u_{\rm eq}(t)\!=\!g(t)\!+\!(\mathcal{K}u_{\rm eq})(t),\quad t\in[t^{*},t^{1}].\vskip-2.84526pt$ (44)
 This integral equation has a unique solution since
-$M\!=\!\!\operatorname*{ess\,sup}_{t^{*}\leq\tau\leq t\leq t^{1}}\|G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\|d\tau\!<\!+\infty.\vskip-5.69054pt$
+$\small M\!=\!\!\operatorname*{ess\,sup}_{t^{*}\leq\tau\leq t\leq t^{1}}\|G^{-1}(t,x(t))\Phi(t,\tau)\tilde{b}(\tau,x(\tau))\|d\tau\!<\!+\infty.\vskip-5.69054pt$
 Indeed, since $\|\mathcal{K}^{n}\|_{\mathcal{L}(L^{\infty};L^{\infty})}\leq\frac{M^{n}(t^{1}-t^{*})^{n}}{n!}$, then the Neumann series $\sum^{\infty}_{i=0}\mathcal{K}^{i}$ is convergent in the Banach space $\mathcal{L}(L^{\infty};L^{\infty})$ of bounded linear operators $L^{\infty}\to L^{\infty}$, and $u_{\rm eq}=\sum^{\infty}_{i=0}\mathcal{K}^{i}g$ is the solution of the integral equation (44).
@@
 ∎
-The table 1 summarizes classes of solutions of the IDE (8) and some conditions to its right-hand side, which are required to ensure existence/uniqueness.
 Table 1: Types of solutions of the IDE (8) and conditions to $f,\tilde{f},u$
@@
 Classical continuous continuous Lipschitz
-[Lakshmikantham_etal1995:Book] in $(t,x,u)$ in $(t,x)$
+[26] in $(t,x,u)$ in $(t,x)$
 Carathéodory discontin. in $t$; discontin. in $t$; Lipschitz
-[OReganMeehan1998:Book] contin. in $(x,u)$ continuous in $x$
+[34] contin. in $(x,u)$ continuous in $x$
 Filippov discontinuous discontinuous piece-wise Lipschitz;
@@
 continuous in $(t,x)$ affine-in-control system;
-in $(x,u)$ smooth sliding surface
+in $(x,u)$ smooth and invariant
+sliding surface
+The table 1 summarizes classes of solutions of the IDE (8) and some conditions to its right-hand side, which are required to ensure existence/uniqueness.
+Remark 1 (On explicit Euler method for IDEs).
+If $f,\tilde{f}$ are piece-wise locally Lipschitz continuous functions in $x$ then approximate (numerical solution) of the IDE (28) can be found using the explicit Euler method
+$\vskip-2.84526pt\begin{split}x_{k+1}=&x_{k}+hf(t^{0}+kh,x_{k})+hI_{k},\\ I_{k}=&h\sum^{k}_{i=0}\Phi(t^{0}+kh,t^{0}+ih)\tilde{f}(t^{0}+ih,x_{i}),\end{split}$ (45)
+where $h>0$ is the sampling period, $k=0,1,2,$..., $x_{k}$ approximates $x(t^{0}+kh)$ and $I_{k}$ approximates the integral $\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{f}(\tau,x(\tau))d\tau$ by means of the method of rectangles.
+Indeed, since, on continuity domains, both functions $f,\tilde{f}$ are locally Lipschitz continuous in $x$, any solution $x:\mathcal{I}\to\mathbb{R}^{n}$ of IDE (28) belonging to this domain is unique and the sequence of the Euler approximates
+$x_{h}(t)\!=\!\tfrac{t^{0}\!+(k+1)h-t}{h}x_{k}+\tfrac{t-t^{0}\!-kh}{h}x_{k+1},t\!\in\![t^{0}\!+kh,t^{0}\!+(k+1)h],$
+(being uniformly bounded and equicontinuous on $\mathcal{I}$) converges to $x$ uniformly on $\mathcal{I}$ as $h\to 0$. This conclusion remains valid if the discontinuous IDE (28) has no sliding mode.
+In the sliding mode case, similarly to ODEs, the explicit Euler method correctly approximates the Filippov/Utkin solution of the IDE away an $\epsilon$-neighborhood of a sliding set $\mathcal{S}$, where $\epsilon\to 0$ as $h\to 0$.
+A consistent discrete-time approximation of the motion on the sliding surface requires development of advanced numerical schemes [1], [42].
+This goes out of the scope of the paper.
 4 Application of discontinuous IDEs for SMC design and analysis
@@
 This IDE can be considered as a linear system with a distributed input delay.
-SMC design for a linear system with a distributed state delay can be found in [Gouaisbaut_etal2001:TDS].
+SMC design for a linear system with a distributed state delay can be found in [20].
 The aim is to stabilize the output
 $y(t)=Cx(t),\quad C\in\mathbb{R}^{m\times n},\quad\det(CB)\neq 0\vskip-2.84526pt$ (47)
-to zero in a finite time. For ODEs (i.e. for $\Phi=\mathbf{0}$), the SMC theory (see, [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book]) suggests the feedback
-$u(t)=-\rho(t)(CB)^{-1}\frac{y(t)}{\|y(t)\|_{P}},\quad\rho(t)>0,\vskip-2.84526pt$ (47)
+to zero in a finite time. For ODEs (i.e. for $\Phi=\mathbf{0}$), the SMC theory (see, [49], [14], [46]) suggests the feedback
+$u(t)=-\rho(CB)^{-1}\frac{y(t)}{\|y(t)\|_{P}},\quad\rho>0,\vskip-2.84526pt$ (48)
 where $\|y\|_{P}=\sqrt{y^{\top}Py}$ is the weighted Euclidean norm with a positive definite matrix $P\succ 0$.
@@
 The condition $\det(CB)\neq 0$ means that the output has the relative degree 1 with respect to the input.
-This is the standard assumption for the first order sliding mode control design [Utkin1992:Book], [Utkin_etal2020:Book].
-If the output has a higher relative degree with respect to the input, then, to solve the considered regulation problem, the high order sliding mode control design algorithms [Levant2003:IJC], [Shtessel_etal2014:Book] must be extended to the IDE (45).
+This is the standard assumption for the first order sliding mode control design [49], [50].
+If the output has a higher relative degree with respect to the input, then, to solve the considered regulation problem, the high order sliding mode control design algorithms [29], [46] must be extended to the IDE (46).
 This extension goes out of the scope of the paper.
@@
 If $CA\!=\!\Lambda C$ with $\Lambda\!\in\!\mathbb{R}^{m\times m}$ such that
-[…]
-$M=\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|C\Phi(t,\tau)\tilde{B}(CB)^{-1}\|_{P}d\tau<1,\vskip-2.84526pt$ (49)
-$\rho(t)\!\geq\!\tfrac{\left\|C(B\gamma(t)+p(t))\right\|_{P}+\int^{t}_{t^{0}}\!\|C\Phi(t,\tau)\tilde{B}\gamma(\tau)\|_{P}d\tau+\delta}{1-M},\;\delta\!>\!0,\vskip-2.84526pt$ (50)
+$\Lambda^{\top}P+P\Lambda\preceq 0,\quad 0\prec P=P^{\top}\in\mathbb{R}^{m\times m},\vskip-2.84526pt$ (49)
+$\rho\!\geq\!\tfrac{\left\|C(B\gamma(t)+p(t))\right\|_{P}+\int^{t}_{t^{0}}\!\|C\Phi(t,\tau)\tilde{B}\gamma(\tau)\|_{P}d\tau+\delta}{1-M},\vskip-2.84526pt$ (50)
+$\small\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|C\Phi(t,\tau)\tilde{B}(CB)^{-1}\|_{P}d\tau\leq M<1,\;\;\delta\!>\!0,\vskip-2.84526pt$ (51)
 then, any Filippov solution of the closed-loop system (46), (47), (48) exists and defined for all $t\geq t^{0}$.
@@
 For any Filippov solution of the closed-loop system (46), (47), (48) it holds
-$\begin{split}\tfrac{d\|x(t)\|}{dt}\!\leq&\|Ax(t)\|\!+\!\rho\|B(CB)^{-1}y\|/\|y\|_{P}\!+\!\|B\gamma\|_{L^{\infty}}\!+\!\|p\|_{L^{\infty}}\!\\ &\!\!+\!\left(\!\rho\|\tilde{B}(CB)^{-1}y\|/\|y\|_{P}\!+\!\|\tilde{B}\gamma\|_{L^{\infty}}\!\right)\!\!\int^{t}_{t^{0}}\!\!\|\Phi(t,\tau)\|d\tau.\end{split}\vskip-2.84526pt$
-Since $\Phi$ is measurable and locally bounded on $\Delta$ then the function $t\mapsto\int^{t}_{t^{0}}\|\Phi(t,\tau)\|d\tau$ is measurable and locally bounded on $t_{0},+\infty)$.
-So, using Grönwall–Bellman inequality, we conclude that $\|x(t)\|<+\infty$ for any finite $t\geq t^{0}$.
+$\small\begin{split}\tfrac{d\|x(t)\|}{dt}\!\leq&\|Ax(t)\|\!+\!\rho\|B(CB)^{-1}y\|/\|y\|_{P}\!+\!\|B\gamma\|_{L^{\infty}}\!+\!\|p\|_{L^{\infty}}\!\\ &\!\!+\!\left(\!\rho\|\tilde{B}(CB)^{-1}y\|/\|y\|_{P}\!+\!\|\tilde{B}\gamma\|_{L^{\infty}}\!\right)\!\!\int^{t}_{t^{0}}\!\!\|\Phi(t,\tau)\|d\tau.\end{split}\vskip-2.84526pt$
+Since $\Phi$ is measurable and locally bounded on $\Delta$ then the function
+$t\mapsto\int^{t}_{t^{0}}\|\Phi(t,\tau)\|d\tau$
+is measurable and locally bounded on $t_{0},+\infty)$. So, using Grönwall–Bellman inequality, we conclude that $\|x(t)\|<+\infty$ for any finite $t\geq t^{0}$.
 Assumption $CA=\Lambda C$ implies that the output $y$ satisfies
-$\begin{split}\dot{y}(t)\in&\,{\Lambda y(t)+}CB\gamma(t)-\tfrac{\rho y(t)}{\|y(t)\|_{P}}+Cp(t)\\ &+\!\int^{t}_{t^{0}}C\Phi(t,\tau)\tilde{B}\left(\gamma(\tau)-\tfrac{\rho(CB)^{-1}y(\tau)}{\|y(\tau\|)_{P}}\right)d\tau\end{split}\vskip-2.84526pt$ (52)
+$\small\begin{split}\dot{y}(t)\in&\,\Lambda y(t)+CB\gamma(t)-\tfrac{\rho y(t)}{\|y(t)\|_{P}}+Cp(t)\\ &+\!\int^{t}_{t^{0}}C\Phi(t,\tau)\tilde{B}\left(\gamma(\tau)-\tfrac{\rho(CB)^{-1}y(\tau)}{\|y(\tau\|)_{P}}\right)d\tau\end{split}\vskip-2.84526pt$ (53)
 for almost all $t\geq t^{0}$. Using (51), (50) and (49), we derive
@@
 as long as $|y(t)|\neq 0$, so $y(t)=0$ for all $t\geq T$. ∎
+If $\|\gamma\|_{L^{\infty}}\leq\bar{\gamma}$ and $\|p\|_{L^{\infty}}\leq\bar{p}$ then the inequality (51) holds for $\rho=(\sqrt{\|P\|}(\|CB\|\bar{\gamma}+\|C\|\bar{p}+M\|CB\|\bar{\gamma})+\delta)/(1-M)$.
+Such a design based on upper estimates on uncertain parameters of the system perfectly follows SMC theory [49], [14] and appreciated by SMC practice [48].
 The above theorem presents a sufficient condition under which $Cx=0$ is a sliding surface of the closed-loop system (46), (47), (48).
-The SMC theory [Utkin1992:Book] also studies a motion of the system on this surface.
+The SMC theory [49] also studies a motion of the system on this surface.
 Similarly to ODEs, the sliding motion can be studied using equivalent control method.
@@
 $\begin{split}\!\!\!\!\!\dot{x}(t)\!=&\,Ax(t)\!+\Pi p(t)\!+\!\Pi\!\int^{t}_{t-d}\!\!\!\!\Phi(t,\tau)\tilde{B}\tilde{u}_{\rm eq}(\tau)d\tau,\\ 0=&\,CB\tilde{u}_{\rm eq}(t)+Cp(t)+\int^{t}_{t-d}C\Phi(t,\tau)\tilde{B}\tilde{u}_{\rm eq}(\tau)d\tau,\vskip-2.84526pt\end{split}$ (62)
+Hence, the condition (61) implies that $\operatorname*{ess\,sup}_{s\geq t}\|\tilde{u}_{\rm eq}(t)\|\to 0$ as $t\to+\infty$ and
+$\operatorname*{ess\,sup}_{s\geq t}\left\|\int^{t}_{t-d}\Phi(t,\tau)\tilde{B}\tilde{u}_{\rm eq}(\tau)d\tau\right\|\to 0\text{ as }t\to+\infty.$
 […]
@@
 Therefore, by Theorem 20, the SMC (48) with $\rho>6\|\gamma\|_{L^{\infty}}$ stabilizes the output $y=Cx$ of the system (46) to zero in a finite time.
-The results of numerical simulations for $\rho=4$, $\gamma(t)=0.5\cos(2t)$ and $t^{0}=0$ are depicted on Figures 1, 2.
+The results of numerical simulations for $\rho=4$, $\gamma(t)=0.5\cos(2t)$ and $t^{0}=0$ are depicted on Figures 1, 2 and 3.
 The simulations have been done by means of explicit Euler method (45) with the sampling period $h=0.001$.
 The integral is computed using the method of rectangles using the same sampling period.
-The control signal depicted on Figure 3 is the typical signal generated by SMC, which has a bounded chattering during the sliding mode (after the time instant $t=T\approx 0.55$).
-Recall that (see, e.g., [Utkin1992:Book]), in practice, the equivalent control can be estimated by filtering the discontinuous input signal $u(t)$:
+The applicability of this method is discussed in Remark 1.
+To compare the SMC for IDE with the SMC for ODE , we also present the simulation results of the system (46)-(48) with $\Phi=\mathbf{0}$ (please see Figures 4, 5 and 6).
+The simulations show a faster convergence of the ODE ($\Phi=\mathbf{0}$) system to zero.
+The control signal depicted on Figure 3 is the typical signal generated by SMC, which has a chattering during the sliding mode (after the time instant $t=T\approx 0.55$).
+The chattering is caused by the use of the explicit method, which correctly approximates the dynamics of the SMC system only away from the sliding surface (see Remark 1 or [1]).
+Recall that (see, e.g., [49]), in practice, the equivalent control can be estimated by filtering the discontinuous input signal $u(t)$:
 $\epsilon\dot{u}_{\epsilon}(t)=-u_{\epsilon}(t)+u(t),\quad t>t^{0},\quad u_{\epsilon}(t^{0})=0,\vskip-2.84526pt$ (63)
@@
 $\delta(t)=CB(u_{\epsilon}(t)+\gamma(t))+\int^{t}_{t^{0}}C\Phi(t,\tau)\tilde{B}(u(\tau)+\gamma(\tau))d\tau\vskip-2.84526pt$
-can be considered as a sliding mode indicator.
-It must be close to zero for $t\geq T$, provided that Utkin solution is an adequate description of a motion of the considered discontinuous control system.
+can be considered as a practical sliding mode indicator.
+Indeed, if the Utkin solution is an adequate description of a motion of the considered discontinuous control system and the explicit Euler method gives a precise enough approximate of Utkin solution of IDE then the approximate identity $\sigma(t)\approx 0$ may be treated as an indication of a practical sliding mode for $t\geq T$.
 The evolution of the signal $\delta$ computed for $\epsilon=0.01$ is depicted on Figure 7.
-The numerical computation of $\delta$ has been done using the sampling period $h=0.0001$.
-The obtained $\delta$ is, indeed, close to zero for $t\geq 0.6$.
-This indirectly confirms a validity of the Utkin solutions for description of a motion of discontinuous SMC control systems modeled by IDEs.
-A small delay in detection of SMC is caused by the inertia of the linear filter (62) and by numerical errors.
+The numerical computation of $\delta$ has been done with the sampling period $h=0.0001$.
+The obtained $\delta(t)$ is, indeed, close to zero for $t\geq 0.6$.
+This may be interpreted as some heuristic confirmation of the validity of Utkin solutions for description of a motion of discontinuous SMC control systems modeled by IDEs.
+A small delay in detection of SMC is caused by an inertia of the linear filter (63) and by numerical errors of the explicit Euler method.
+Development of a rigorous numerical analysis for IDEs with SMC is the important problem for the future research.
 Figure 1: The states of the closed-loop system (46)-(48)
@@
 Figure 3: The control input of the closed-loop system (46)-(48)
-Figure 4: The sliding mode indicator $\delta$ of the closed-loop system (45)- (47)
-To compare the SMC for IDE with the SMC for ODE , we also present the simulation results of the system (45)-(47) with $\Phi=\mathbf{0}$ (please see Figures 5, 6 and 7).
-The simulations show a faster convergence of the delay-free ($\Phi=\mathbf{0}$) system to zero.
-Figure 5: The states of the closed-loop system (45)-(47) with $\Phi=\mathbf{0}$
-Figure 6: The output of the closed-loop system (45)-(47) with $\Phi=\mathbf{0}$
-Figure 7: The control input of the closed-loop system (45)-(47) with $\Phi=\mathbf{0}$
 4.2 SMC for dynamical systems in Banach spaces
-Let us consider a control system modeled by a differential equation in a Banach space [DaleckyKrein1974:Book], [Yosida1980:Book], [Pazy1983:Book]
+Let us consider a control system modeled by a differential equation in a Banach space [12], [51], [40]
 $\dot{x}(t)\!=\!\!Ax(t)\!+\!B(u(t)+\gamma(t)),\,t\!>\!t_{0},\,x(t_{0})\!=\!x^{0}\!\!\in\!\mathcal{D}(A),\vskip-2.84526pt$ (64)
@@
 where $C:\mathbb{B}\to\mathbb{R}^{k}$ is a linear bounded operator.
-For any $u\in L^{1}_{\rm loc}([t^{0},+\infty);\mathbb{R}^{m})$, the open-loop systems (63) is well-posed and has the so-called mild solution [Pazy1983:Book, page 106]
+For any $u\in L^{1}_{\rm loc}([t^{0},+\infty);\mathbb{R}^{m})$, the open-loop systems (64) is well-posed and has the so-called mild solution [40, page 106]
 $x(t)=\Psi(t-t^{0})x^{0}+\int^{t}_{t^{0}}\Psi(t-\tau)B(u(\tau)+\gamma(\tau))d\tau,\vskip-2.84526pt$ (66)
-where the integral is understood in the sense of Bochner [Yosida1980:Book].
-The mild solution $x:[t^{0},t^{0}+T)\to\mathbb{B}$ with $0<T\leq+\infty$ is the classical (resp., strong [Pazy1983:Book, page 109]) solution of (63) if $x(t)\!\in\!\mathcal{D}(A)$ satisfies (63) for (almost) all $t\in[t^{0},t^{0}+T)$.
+where the integral is understood in the sense of Bochner [51].
+The mild solution $x:[t^{0},t^{0}+T)\to\mathbb{B}$ with $0<T\leq+\infty$ is the classical (resp., strong [40, page 109]) solution of (64) if $x(t)\!\in\!\mathcal{D}(A)$ satisfies (64) for (almost) all $t\in[t^{0},t^{0}+T)$.
+Figure 4: The states of the closed-loop system (46)-(48) with $\Phi=\mathbf{0}$
+Figure 5: The output of the closed-loop system (46)-(48) with $\Phi=\mathbf{0}$
+Figure 6: The control input of the closed-loop system (46)-(48) with $\Phi=\mathbf{0}$
 Let a feedback law be defined as follows
@@
 where $\tilde{u}$ is a measurable function.
-Even if a function $\tilde{u}$ is continuous in $y$, the non-linearity in the right-hand side of the closed-loop system may not be locally Lipschitz continuous, so, the conventional theory of nonlinear abstract differential equations (see, e.g., [Pazy1983:Book], [EngelNagel2000:Book]) is not applicable for its analysis.
-In the case of SMC design [Orlov_elal2004:IJC], the closed-loop system (63), (64), (66) is ill-posed.
-Let us regularize the system (63), (64) following the concept of Utkin solutions [Orlov2008:Book], [PolyakovFridman2014:JFI].
+Even if a function $\tilde{u}$ is continuous in $y$, the non-linearity in the right-hand side of the closed-loop system may not be locally Lipschitz continuous, so, the conventional theory of nonlinear abstract differential equations (see, e.g., [40], [15]) is not applicable for its analysis.
+In the case of SMC design [35], the closed-loop system (64), (65), (67) is ill-posed.
+Let us regularize the system (64), (65) following the concept of Utkin solutions [37], [43].
 Definition 9.
@@
 In other words, an Utkin solution $x$ of (64), (65), (67) is a mild solution with a properly selected $u(t)\!\in\!U(t,Cx(t))$.
-If $B$ maps $\mathbb{R}^{m}$ into the domain of $A$ (i.e., $\mathrm{range}(B)\!\subset\!\mathcal{D}(A)$), then, by [Pazy1983:Book, Theorem 2.9, Chapter 4], any mild solution is a strong solution.
-By [Pazy1983:Book, Theorem 2.4, Chapter 1], the function
+If $\mathrm{range}(B)\!\subset\!\mathcal{D}(A)$, then, by [40, Theorem 2.9, Chapter 4], any mild solution is a strong solution. By [40, Theorem 2.4, Chapter 1], the function
 $s\in[0,+\infty)\mapsto\Phi(s):=C\Psi(s)AB\in\mathbb{R}^{k\times m}\vskip-2.84526pt$
 is continuous.
-Moreover, taking into account $\frac{d\Psi(s)B}{ds}=\Psi(s)AB$ and $\frac{d\Psi(s)x^{0}}{ds}=\Psi(s)Ax^{0},\forall s>0,\forall x^{0}\in\mathcal{D}(A)$ (see, [Pazy1983:Book, Theorem 2.4, Chapter 1]), we derive
-$\begin{split}\tfrac{d}{dt}y(t)=\tfrac{d}{dt}Cx(t)=&\frac{d}{dt}C\Psi(t-t)x^{0}\\ &+\frac{d}{dt}\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)B}_{\in\mathbb{R}^{k\times m}}(u(\tau)+\gamma(\tau))d\tau\\ =&C\Psi(t-t)Ax^{0}+CB(u(t)+\gamma(t))\\ &+\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)AB}_{\Phi(t-\tau)}(u(\tau)+\gamma(\tau))d\tau\end{split}$
+Moreover, taking into account $\frac{d\Psi(s)B}{ds}=\Psi(s)AB$ and $\frac{d\Psi(s)x^{0}}{ds}=\Psi(s)Ax^{0},\forall s>0,\forall x^{0}\in\mathcal{D}(A)$ (see, [40, Theorem 2.4, Chapter 1]), we derive
+$\small\begin{split}\tfrac{d}{dt}y(t)=\tfrac{d}{dt}Cx(t)=&\frac{d}{dt}C\Psi(t-t)x^{0}\\ &+\frac{d}{dt}\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)B}_{\in\mathbb{R}^{k\times m}}(u(\tau)+\gamma(\tau))d\tau\\ =&C\Psi(t-t)Ax^{0}+CB(u(t)+\gamma(t))\\ &+\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)AB}_{\Phi(t-\tau)}(u(\tau)+\gamma(\tau))d\tau\end{split}$
 almost everywhere. Therefore, the input-output dynamics of the system (64), (65) is given by the IDE
-$\dot{y}(t)=p(t-t^{0})+CB(u(t)+\gamma(t))+\int^{t}_{\sigma}\!\!\Phi(t-\tau)(u(\tau)+\gamma(\tau))d\tau,\vskip-2.84526pt$ (67)
+$\dot{y}(t)=p(t-t^{0})+CB(u(t)+\gamma(t))+\int^{t}_{\sigma}\!\!\Phi(t,\tau)(u(\tau)+\gamma(\tau))d\tau,\vskip-2.84526pt$ (68)
 where $t>t^{0}$, $CB\in\mathbb{R}^{k\times m}$ and
@@
 Notice that the function $s\mapsto p(s)$ is continuous if $x_{0}\in\mathcal{D}(A)$.
+Figure 7: The sliding mode indicator $\delta$ of the closed-loop system (46)- (48)
 Let $y_{\rm eq}$ be an Utkin solution of the IDE (68), (69) with the feedback law (67), and let $u_{\rm eq}$ be the corresponding equivalent control.
 By Theorem 19, the pair $(y,u_{\rm eq})$ always exists if $\tilde{u}$ is a locally essentially bounded measurable function.
-If $x_{\rm eq}$ is a mild solution of (63) given by (65) with $u\!=\!u^{\rm eq}$, then
-$\begin{split}\tfrac{d}{dt}Cx_{\rm eq}(t)=&C\Psi(t-t)Ax^{0}+CB(u_{\rm eq}(t)+\gamma(t))\\ &+\int^{t}_{\sigma}\!\!\Phi(t-\tau)(u_{\rm eq}(\tau)+\gamma(\tau))d\tau,\end{split}\vskip-2.84526pt$
+If $x_{\rm eq}$ is a mild solution of (64) given by (66) with $u\!=\!u_{\rm eq}$, then
+$\small\begin{split}\tfrac{d}{dt}Cx_{\rm eq}(t)=&C\Psi(t-t)Ax^{0}+CB(u_{\rm eq}(t)+\gamma(t))\\ &+\int^{t}_{\sigma}\!\!\Phi(t-\tau)(u_{\rm eq}(\tau)+\gamma(\tau))d\tau,\end{split}\vskip-2.84526pt$
 i.e., $y_{\rm eq}(t)=Cx_{\rm eq}(t)$ and $u_{\rm eq}(t)\in U(t,Cx(t))$ almost everywhere.
@@
 Since $\mathrm{range}(B)\subset\mathcal{D}(A^{2})$ then $\frac{d\Psi(s)AB}{ds}=\Psi(s)A^{2}B,\forall s>0$ and
-$\begin{split}-\xi(t,t^{0},x^{0})=&\,\Phi(t-t^{0})\xi(t^{0},t^{0},x^{0})-\Phi(0)\xi(t,t^{0},x^{0})\\ =&\int^{t-t^{0}}_{0}\frac{d}{d\sigma}\left(\Phi(\sigma)\xi(t-\sigma,t^{0},x^{0})\right)d\sigma\\ =&\int^{t-t^{0}}_{0}C\Psi(\sigma)A^{2}B\xi(t-\sigma,t^{0},x^{0})d\sigma\\ &-\int^{t-t^{0}}_{0}\Phi(\sigma)\dot{\xi}(t-\sigma,t^{0},x^{0})d\sigma\end{split}\vskip-2.84526pt$
+$\small\begin{split}-\xi(t,t^{0},x^{0})=&\,\Phi(t-t^{0})\xi(t^{0},t^{0},x^{0})-\Phi(0)\xi(t,t^{0},x^{0})\\ =&\int^{t-t^{0}}_{0}\frac{d}{d\sigma}\left(\Phi(\sigma)\xi(t-\sigma,t^{0},x^{0})\right)d\sigma\\ =&\int^{t-t^{0}}_{0}C\Psi(\sigma)A^{2}B\xi(t-\sigma,t^{0},x^{0})d\sigma\\ &-\int^{t-t^{0}}_{0}\Phi(\sigma)\dot{\xi}(t-\sigma,t^{0},x^{0})d\sigma\end{split}\vskip-2.84526pt$
 Hence, $\int^{t}_{t^{0}}\Phi(t-\tau)\dot{\xi}(\tau,t^{0},x^{0})d\tau=\xi(t,t^{0},x^{0})+\int^{t}_{t^{0}}C\Psi(t-\tau)A^{2}B\xi(\tau,t^{0},x^{0})d\tau$ and the pair $y(t,t^{0},x^{0})=Cx(t,t^{0},x^{0})$, $\xi(t,t^{0},x^{0})$ is an Utkin solution of the IDE
-$\begin{split}\dot{y}(t,t^{0},x^{0})\in&\,p(t,t^{0},x^{0})+CB(U(t,y(t,t^{0},x^{0}))+\gamma(t))\\ &+\int^{t}_{t^{0}}\Phi(t-\tau)\gamma(\tau)d\tau+\xi(t,t^{0},x^{0})\\ &+\int^{t}_{t^{0}}C\Psi(t-\tau)A^{2}B\xi(\tau,t^{0},x^{0})d\tau,\\ \dot{\xi}(t,t^{0},x^{0})\in&\,U(t,y(t,t^{0},x^{0})),\quad\xi(t^{0})=0,\quad y(t^{0})=Cx^{0},\end{split}\vskip-2.84526pt$
+$\small\begin{split}\dot{y}(t,t^{0},x^{0})\in&\,p(t,t^{0},x^{0})+CB(U(t,y(t,t^{0},x^{0}))+\gamma(t))\\ &+\int^{t}_{t^{0}}\Phi(t-\tau)\gamma(\tau)d\tau+\xi(t,t^{0},x^{0})\\ &+\int^{t}_{t^{0}}C\Psi(t-\tau)A^{2}B\xi(\tau,t^{0},x^{0})d\tau,\\ \dot{\xi}(t,t^{0},x^{0})\in&\,U(t,y(t,t^{0},x^{0})),\qquad\xi(t^{0})=0,\qquad y(t^{0})=Cx^{0},\end{split}\vskip-2.84526pt$
 where $p(t,t^{0},x^{0})=C\Psi(t-t^{0})Ax^{0}$.
@@
 The example (1)–(6) admits the representation (64) with $A\!=\!\nu\frac{\partial^{2}}{\partial z^{2}}$, $D(A)\!=\!H^{2}((0,1);\mathbb{R})\cap H^{1}_{0}((0,1);\mathbb{R})$, $(Bx)(z)\!=\!\beta(z)x(z)$, where $H^{2}$, $H^{1}_{0}$ are Sobolev spaces.
-The closed densely defined operator $A:\mathcal{D}(A)\subset L^{2}((0,1);\mathbb{R})\to L^{2}((0,1);\mathbb{R})$ is a generator of the strongly continuous semigroup of linear bounded operators on $L^{2}((0,1);\mathbb{R})$ (see, e.g., [Pazy1983:Book, page 212]).
+The closed densely defined operator $A:\mathcal{D}(A)\subset L^{2}((0,1);\mathbb{R})\to L^{2}((0,1);\mathbb{R})$ is a generator of the strongly continuous semigroup of linear bounded operators on $L^{2}((0,1);\mathbb{R})$ (see, e.g., [40, page 212]).
 The output (65) is given by $y(t)=Cx(t)=\int^{1}_{0}\xi(z)x(t,z)dz$.
@@
 Since $\beta\in\mathcal{D}(A)$, then, by Corollary 6, the system (1)–(6) has an Utkin solution $x(t)\!\in\!\mathcal{D}(A)$, $\forall x^{0}\in\mathcal{D}(A)$, $\forall q\!\in\!L^{\infty}((t^{0},+\infty);\mathbb{R})$ in (6).
-2) SMC design using PDE-based approach.
+2) SMC design using PDE.
 The closed-loop heat control system (1)-(6) is forward complete, i.e., all Utkin solutions exist globally in time.
 Indeed, considering $V(t)\!=\!\int^{1}_{0}\!x^{2}(t,z)dz$ and using integration by parts, we derive
-$\dot{V}(t)=2\int^{1}_{0}\nu x(t,z)\tfrac{\partial^{2}x(t,z)}{\partial z^{2}}+x(t,z)\beta(z)u_{\rm eq}(t)dz\vskip-5.69054pt$
-$2\int^{1}_{0}\nu\left(\tfrac{\partial x(t,z)}{\partial z}\right)^{2}+x(t,z)\beta(z)u_{\rm eq}(t)dz,\vskip-2.84526pt$
+$\vskip-2.84526pt\small\begin{split}\dot{V}(t)=&\,2\int^{1}_{0}\nu x(t,z)\tfrac{\partial^{2}x(t,z)}{\partial z^{2}}+x(t,z)\beta(z)(u_{\rm eq}(t)+\gamma(t))dz\\[-5.69054pt] =&-2\int^{1}_{0}\nu\left(\tfrac{\partial x(t,z)}{\partial z}\right)^{2}+x(t,z)\beta(z)(u_{\rm eq}(t)+\gamma(t))dz,\end{split}$
 where $u_{\rm eq}(t)\in q(t)\,\overline{\mathrm{sign}}(Cx(t))$.
@@
 Using Wirtinger and Cauchy-Schwartz inequalities, we come up with the estimate
-$\dot{V}(t)\leq-\nu 2\pi^{2}V(t)+2\|q\|_{L^{\infty}}\|\beta\|_{L^{2}}\sqrt{V(t)},$
+$\dot{V}(t)\leq-\nu 2\pi^{2}V(t)+2(\|q\|_{L^{\infty}}+\|\gamma\|_{L^{\infty}})\|\beta\|_{L^{2}}\sqrt{V(t)},$
 that is fulfilled almost everywhere. Therefore, any Utkin solution of the closed-loop system is bounded and
-$\limsup_{t\to+\infty}\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{\|q\|_{L^{\infty}}\|\beta\|_{L^{2}}}{\nu\pi^{2}}.\vskip-2.84526pt$ (69)
+$\limsup_{t\to+\infty}\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{(\|q\|_{L^{\infty}}+\|\gamma\|_{L^{\infty}})\|\beta\|_{L^{2}}}{\nu\pi^{2}}.\vskip-2.84526pt$ (70)
 Since $CB=\int^{1}_{0}\xi(z)\beta(z)dz$ and $Cx(t)=\int^{1}_{0}\xi(z)x(t,z)dz$ then, due to $\xi(0)=\xi(1)=x(t,0)=x(t,1)=0$ and $\xi\in C^{2}([0,1];\mathbb{R})$, the integration by parts yields
 $\begin{split}\tfrac{dCx(t)}{dt}=&\int^{1}_{0}\nu\xi(z)\tfrac{\partial^{2}x(t,z)}{\partial z^{2}}dz+CB(u_{\rm eq}(t)+\gamma(t))\\ =&\int^{1}_{0}\nu\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz+CB(u_{\rm eq}(t)+\gamma(t))\\ =&-\nu\lambda Cx(t)+\nu\int^{1}_{0}\!\!\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}\!+\!\lambda\xi(z)\right)x(t,z)dz\\ &+CB(u_{\rm eq}(t)+\gamma(t)),\quad\quad\forall\lambda\geq 0.\end{split}\!\vskip-2.84526pt$ (71)
-Taking into account that $u_{\rm eq}(t)\in-q(t)\overline{\mathrm{sign}}(Cx(t))$, we conclude that, for $Cx(t)\!\neq\!0$, it holds $u_{\rm eq}(t)\!=\!-q(t)\mathrm{sign}(Cx(t))$,
+Taking into account that $u_{\rm eq}(t)\in q(t)\overline{\mathrm{sign}}(Cx(t))$, we conclude that, for $Cx(t)\!\neq\!0$, it holds $u_{\rm eq}(t)\!=\!q(t)\mathrm{sign}(Cx(t))$,
 $\begin{split}\tfrac{d|Cx(t)|}{dt}\leq&\,-\nu\lambda|Cx(t)|+CBq(t)\\ &+\nu\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\!\|x(t,\cdot)\|_{L^{2}}+\|CB\gamma\|_{L^{\infty}}\end{split}\vskip-2.84526pt$
-almost everywhere. If $CB=\int^{1}_{0}\xi(z)\beta(z)dz\neq 0$ then, taking $\delta>0$ and $q(t)=-\frac{\rho}{CB}$ with $\rho>0$, we have due to (69)):
-$\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{\rho\|\beta\|_{L^{2}}}{\pi^{2}|CB|}+\delta\vskip-2.84526pt$
+almost everywhere. If $CB=\int^{1}_{0}\xi(z)\beta(z)dz\neq 0$ then, taking $\delta>0$ and $q(t)=-\frac{\rho}{CB}$ with $\rho>0$, we have due to (70):
+$\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{(\rho+\|\gamma\|_{L^{\infty}}|CB|)\|\beta\|_{L^{2}}}{\nu\pi^{2}|CB|}+\delta\vskip-2.84526pt$
 for almost all $t\geq t^{0}+\tilde{T}(x_{0})$ provided that $\tilde{T}(x_{0})\geq 0$ is large enough. If $\exists\lambda\geq 0$ such that
@@
 then, for
-$\rho>\tfrac{\pi^{2}|CB|\left(\bar{\gamma}|CB|+\nu\delta\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\right)}{\pi^{2}\left|CB\right|-\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}},\quad\bar{\gamma}\geq\|\gamma\|_{L^{\infty}},\vskip-2.84526pt$
+$\rho\!>\!\tfrac{|CB|\left(\pi^{2}\bar{\gamma}|CB|+\left(\nu\delta\pi^{2}+\bar{\gamma}\|\beta\|_{L^{2}}\right)\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\right)}{\pi^{2}\left|CB\right|-\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}},\quad\bar{\gamma}\!\geq\!\|\gamma\|_{L^{\infty}},\vskip-2.84526pt$
 we derive the following estimate
-$\tfrac{d|Cx(t)|}{dt}\!\leq\!\nu\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}\!+\!\lambda\xi\right\|_{L^{2}}\!\left(\tfrac{\rho\|\beta\|_{L^{2}}}{\nu\pi^{2}|CB|}\!+\!\delta\right)\!+|CB|\cdot\|\gamma\|_{\infty}-\rho\!<\!0\;$
+$\small\vskip-2.84526pt\begin{split}\tfrac{d|Cx(t)|}{dt}\!\leq&\nu\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}\!+\!\lambda\xi\right\|_{L^{2}}\!\left(\tfrac{(\rho+\|\gamma\|_{L^{\infty}}|CB|)\|\beta\|_{L^{2}}}{\nu\pi^{2}|CB|}\!+\!\delta\right)\\ &+|CB|\cdot\|\gamma\|_{\infty}-\rho<0\end{split}$
 for almost all $t\geq t^{0}+\tilde{T}(x^{0})$ such that $Cx(t)\neq 0$.
@@
 for almost all $t\geq T$.
-3) SMC design using IDE-based approach. The same result can be obtained considering the input-output dynamics of the heat control system. Using the modal decomposition
+3) SMC design using IDE. The same result can be obtained considering the input-output dynamics of the heat control system. Using the modal decomposition
 $x(t,z)=\sum_{i=1}^{\infty}x_{i}(t)\phi_{i}(z),\quad x_{i}(t)\in\mathbb{R},\vskip-2.84526pt$ (75)
-where $\phi_{i}(z)=\sqrt{2}\sin(\sqrt{\lambda_{i}}z),\lambda_{i}=\pi^{2}i^{2}$, the system (1) - (3) can be rewritten (see, e.g.
-[Balas1988:JMAA], [PrieurTrelat2019:TAC], [KatzFridman2021:EJC]) in the form
+where $\phi_{i}(z)=\sqrt{2}\sin(\sqrt{\lambda_{i}}z),\lambda_{i}=\pi^{2}i^{2}$, the system (1) - (3) can be rewritten (see, e.g. [5], [45], [22]) in the form
 $\dot{x}_{i}(t)=-\nu\lambda_{i}x_{i}(t)+b_{i}(u(t)+\gamma(t)),\quad t>t_{0},\vskip-2.84526pt$ (76)
@@
 $\begin{split}\dot{y}(t)=&-\nu\lambda y(t)+p(t,t^{0})+CB(u(t)+\gamma(t))\\ &+\int^{t}_{t^{0}}\Phi(t,\tau)(u(\tau)+\gamma(\tau))d\tau,\end{split}\vskip-5.69054pt$ (77)
-$p(t,t^{0})=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-t^{0})}x_{i}^{0}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz,\vskip-5.69054pt$
-$\Phi(t,\tau)=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-\tau)}b_{i}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz.\vskip-2.84526pt$
-Since $|\Phi(t,\tau)|\leq\nu e^{-\nu\pi^{2}(t-\tau)}\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}$ then the inequality (71) yields $M=\frac{\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|\Phi(t,\tau)|d\tau}{|CB|}\!<\!1$.
-Let $q(t)=-\frac{|CB|\cdot\|\gamma\|_{L_{\infty}}+|p(t,t^{0})|+M\|\gamma\|_{L^{\infty}}+\delta}{(1-M)CB}$ with $\delta>0$.
+$p(t{\color[rgb]{0,0,1}-}t^{0})=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-t^{0})}x_{i}^{0}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz,\vskip-5.69054pt$
+$\Phi(t{\color[rgb]{0,0,1}-}\tau)=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-\tau)}b_{i}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz.\vskip-2.84526pt$
+Since $|\Phi(t{\color[rgb]{0,0,1}-}\tau)|\leq\nu e^{-\nu\pi^{2}(t-\tau)}\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}$ then the inequality (72) yields $M=\frac{\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|\Phi(t{\color[rgb]{0,0,1}-}\tau)|d\tau}{|CB|}\!<\!1$.
+Let $q(t)=-\frac{|CB|\bar{\gamma}+\bar{p}+M\bar{\gamma}+\delta}{(1-M)CB}$ with $\delta>0$ and $\bar{p}\geq|p(t-t^{0})|$.
 By Theorem 20 and Corollary 4, the system (77), (6) has the unique Utkin solution and $y(t)=0$ for all $t\geq T$, where $T\geq t^{0}$ is some instant of time.
 This means that the Utkin solution of the system (1)–(6) satisfies (74) for all $t\geq T$.
-4) Stability of the heat system in the sliding mode.
+4) Stability analysis of the heat system in the sliding mode.
 Let us show that the system (73), (74), (3) is globally exponentially stable.
@@
 $\dot{V}(t)\leq-2\nu\pi^{2}V(t)-2\nu\tfrac{\int^{1}_{0}\frac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz\int^{1}_{0}\beta(z)x(t,z)dz}{\int^{1}_{0}\xi(z)\beta(z)dz}.\vskip-2.84526pt$
-The identity (73) gives
-$\begin{split}\left|\int^{1}_{0}\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz\right|=&\left|\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)x(t,z)dz\right|\\ \leq&\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\sqrt{V(t)}.\vskip-2.84526pt\end{split}$
+The identity (74) and Cauchy-Schwartz inequality give
+$\small\vskip-2.84526pt\begin{split}\left|\int^{1}_{0}\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz\right|=&\left|\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)x(t,z)dz\right|\\ \leq&\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\sqrt{V(t)}.\vskip-2.84526pt\end{split}$
 Using the inequality $\left|\int^{1}_{0}\beta(z)x(t,z)dz\right|\leq\|\beta\|_{L^{2}}\sqrt{V(t)}$ and the condition (72) we come up with the estimate
-$\dot{V}(t)\leq 2V(t)\left(\tfrac{\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\|\beta\|_{L^{2}}}{|CB|}-\pi^{2}\right)<0\text{ for }V(t)\neq 0,\vskip-2.84526ptx$
-i.e., the system (72), (73), (3) is globally exponentially stable.
+$\dot{V}(t)\leq 2V(t)\left(\tfrac{\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\|\beta\|_{L^{2}}}{|CB|}-\pi^{2}\right)<0\,\text{ for }\,V(t)\neq 0.\vskip-2.84526pt$
+So, the system (73), (74), (3) is globally exponentially stable.
 The condition (72) is fulfilled, for example, if
@@
 The initial state $x^{0}\in\mathcal{D}(A)$ is defined as follows $x_{0}(z)=10z(1-z),z\in[0,1]$.
-The method of modal decomposition [Balas1988:JMAA], [PrieurTrelat2019:TAC], [KatzFridman2021:EJC], [Ayamou_etal2025:IMA] (with 60 modes) is utilized for numerical simulation of the closed-loop SMC system.
+The method of modal decomposition [5], [45], [22], [4] with 60 modes is utilized for numerical simulation of the closed-loop SMC system.
 The time sampling is $0.001$.
@@
 Figure 10: The control input of the heat system (1)-(5) with the SMC (6)
-The finite-dimensional IDE-based procedure of regularization of infinite-dimensional systems with sliding mode control can be extended to other systems using, for example, the methodology suggested in [Polyakov2025:CDC].
-Further study of the SMC design for such systems goes out of the scope of this paper.
+The finite-dimensional IDE-based procedure of regularization of infinite-dimensional systems with sliding mode control can be extended to other systems using, for example, the methodology suggested in [44].
+Being interesting future research topic, a further study of the SMC design for such systems goes out of the scope of this paper.
 5 Conclusions
@@
 Such ill-posed discontinuous systems appear in sliding mode control theory, which uses discontinuous feedback laws for regulation of systems and rejection of matched perturbations.
-The advantage of the Filippov method consists in a minimal (from set-theoretic point if view) extension of an ill-posed equation to a well-posed equation.
+The advantage of the Filippov method consists in a minimal (from set-theoretic point of view) extension of an ill-posed equation to a well-posed inclusion.
 This implies a better matching of the dynamics of the control system by the regularized model.
@@
 The well-posedness of regularized systems is studied.
-Filippov method and equivalent control approach are the main tools for analysis of control systems with sliding modes.
+Filippov method and equivalent control approach are key tools for analysis of control systems with sliding modes.
 To illustrate theoretical results, sliding mode control is designed for a class of systems with distributed input delay.
@@
 The equivalent control is defined as a solution of the Volterra integral equation of the second kind.
-A SMC for a plant modeled by infinite-dimensional differential equation with finite-dimensional inputs and outputs can also be designed using the developed theory of discontinuous IDEs.
-Such a discontinuous infinite-dimensional control system can be regularized based on a finite-dimensional IDE representing its input-output dynamics.
-This scheme is demonstrated on a heat control system modeled by PDE.
-The obtained theoretical results provide all necessary tools for extension of other methods of SMC theory (e.g., high order SMC [Levant2003:IJC] or implicit SMC discretization [AcaryBrogliato2010:SCL]) to systems modeled by integro-differential equations and by differential equations in Banach spaces (e.g., PDEs).
+For a class of infinite-dimensional control systems (64) with finite-dimensional inputs and outputs, a sliding mode control can be designed and analyzed using the developed theory.
+Indeed, the input-output dynamics of such an infinite-dimensional system can always be represented as a finite-dimensional integro-differential equation.
+If the input-output dynamics admits the representation (46), (47), then the SMC for the original system can be designed in the form (48) provided that all conditions of Theorem 20 are fulfilled.
+In Example 8, this approach is demonstrated for a heat control system modeled by PDE.
+The obtained theoretical results provide all necessary tools for extension of other methods of SMC theory (e.g., high order SMC [29] or implicit SMC discretization [1]) to systems modeled by integro-differential equations and by differential equations in Banach spaces (e.g., PDEs).